### v1

# In-Context Learning as Nonparametric Conditional Probability Estimation: Risk Bounds and Optimality

Abstract
This paper investigates the expected excess risk of In-Context Learning (ICL) for multi-class classification. We model each task as a sequence of labeled prompt samples and a query input, where a pre-trained model estimates the conditional class probabilities of the query. The expected excess risk is defined as the average truncated Kullback–Leibler (KL) divergence between the predicted and ground-truth conditional class distributions, averaged over a specified family of tasks.
We establish a new oracle inequality for the expected excess risk based on KL divergence in multiclass classification. This allows us to derive tight upper and lower bounds for the expected excess risk in transformer-based models, demonstrating that the ICL estimator achieves the minimax optimal rate—up to a logarithmic factor—for conditional probability estimation.
From a technical standpoint, our results introduce a novel method for controlling generalization error using the uniform empirical covering entropy of the log-likelihood function class. Furthermore, we show that multilayer perceptrons (MLPs) can also perform ICL and achieve this optimal rate under specific assumptions, suggesting that transformers may not be the exclusive architecture capable of effective ICL.
Keywords: expected excess risk, in-context learning, multi-class classification, multilayer perceptrons, transformers
1 Introduction
The transformer architecture (Vaswani et al., 2017) represents a pivotal milestone in modern machine learning. It has been extensively applied across domains including natural language processing (Devlin et al., 2019), computer vision (Dosovitskiy et al., 2021), and reinforcement learning (Chen et al., 2021), catalyzing significant research advancements in the field of large language models (LLMs). Brown et al. (2020) observed that when a model reaches a certain scale, it can emerge the remarkable capability of in-context learning (ICL). Specifically, pre-trained LLMs can perform novel tasks with just a few labeled examples in the prompt, without any parameter updates or fine-tuning. This capability facilitates exceptional few-shot generalization in LLMs while establishing a novel paradigm for few-shot learning (Brown et al., 2020; Garcia et al., 2023). Consequently, it accelerates the deployment of deep learning models, enabling a foundational shift toward LLMs as instruction-following systems.
A recent line of research has quantified ICL by studying transformer models pretrained on linear regression tasks with Gaussian priors (Garg et al., 2022; Akyürek et al., 2023; Li et al., 2023; Raventós et al., 2023; Wu et al., 2024). Concurrently, a growing body of research has revealed a close connection between the forward pass of multi-layer transformer models and multi-step gradient descent algorithms (Akyürek et al., 2023; Von Oswald et al., 2023; Bai et al., 2023; Ahn et al., 2023; Zhang et al., 2024), suggesting that transformers may implement ICL by implicitly simulating gradient descent. Recent empirical findings show that transformers fail to exhibit ICL unless they are pretrained on a sufficiently diverse set of tasks (Raventós et al., 2023), highlighting the crucial role of task diversity in pretraining. The pioneering experimental study by Garg et al. (2022) demonstrated that transformers can in-context learn a variety of fundamental function classes, thereby sparking theoretical analyses of ICL in statistics. Wu et al. (2024) further quantified how the number of pretraining tasks influences the prediction error of ICL in linear regression. Kim et al. (2024) considered deep neural network with one-layer linear transformer structure and established the minimax optimality of the in-context learner from the view of nonparametric regression.
Although substantial research has analyzed the ICL capabilities of large language models, achieving significant progress in understanding this mechanism, there are still notable limitations and challenges in its theoretical analysis.
First, a significant gap exists between the limit assumptions underlying theoretical analyses and the scenarios typically encountered in practical applications. Most existing studies focus on the combined regime of task scaling and context scaling (Abedsoltan et al., 2024), where both the number of pretraining tasks and the context length are assumed to be large. Some analyses further consider the asymptotic limit as both the task scaling and the context scaling approach infinity. However, the practical importance of few-shot scenarios is neglected, as inference is typically constrained by limited context length. For example, GPT-3 was evaluated under few-shot settings  (Brown et al., 2020), where the context length only allowed one or two labeled examples per class in the prompt. To our best knowledge, Wu et al. (2024) rigorously analyzed the impact of task scaling on the prediction risk of ICL at a fixed context length.
Second, there are substantial limitations in both task settings and model architecture, as current theoretical results do not cover multi-classification tasks or complete transformer architecture. Most theoretical studies on ICL concentrate on simplified transformer architectures, such as linear transformers or single-head attention, to facilitate analytical tractability (Zhang et al., 2024; Huang et al., 2024; Kim and Suzuki, 2024; Li et al., 2024a). However, empirical research demonstrates that in a wide range of ICL tasks, multi-head dot-product attention mechanisms offers substantial performance gains over its single-head counterpart (Cui et al., 2024; Xing et al., 2024). For regression problems, some progresses have been made in understanding ICL based on multi-head dot-product attention (Chen et al., 2024a, b; Deora et al., 2024). Li et al. (2024b) and Li et al. (2024c) considered the binary classification and analyzed the training dynamics of single-layer transformer models. Frei and Vardi (2024) investigated in-context classifcation of Gaussian mixtures using a linear transformer model. Meanwhile, multiclass classification, which is fundamental and widely used in deep learning, has not received adequate theoretical attention. Shen et al. (2024) studied the training dynamics of a single-layer linear transformer model for multi-classification of Gaussian mixtures in the context of ICL.
To address these limitations, we focus on multi-classification tasks and develop a theoretical framework for bounding the expected excess risk of ICL under the task-scaling regime. Specifically, we derive non-asymptotic error bounds under the assumption that the number of pretraining tasks is much larger than the context length. We further analyze asymptotic behavior when the number of pretraining tasks tends to infinite while the context length remains fixed. From the perspective of modeling, we frame ICL as a nonparametric conditional probability estimation problem, where the model estimates conditional class probabilities of a query input given labeled prompt examples. We use a deep transformer model with multi-head dot-product attention and analyze its generalization performance in ICL.
Our contributions are summarized as follows:
  1. 1.
    Our work extends existing theoretical frameworks by relaxing main assumptions in experimental and practical settings. Specifically, we relax constraints commonly adopted in previous studies, such as focusing solely on regression task, using simplified transformer architecture, or assuming asymptotic regimes for context scaling. This enables closer alignment with empirical deployments while preserving analytical rigor. Our work bridges critical gaps between theoretical analysis and real-world implementation.
  2. 2.
    We establish a new oracle inequality for the expected excess risk based on Kullback–Leibler (KL) divergence in multiclass classification. To the best of our knowledge, this is the first result that decomposes the expected excess risk under KL divergence into approximation, optimization, and generalization errors, where the generalization error is controlled by the uniform empirical covering entropy of the log-likelihood function class.
  3. 3.
    Our analysis finds that transformers are not necessarily the uniquely optimal architecture for ICL. By comparing the convergence rates of transformer and MLP architectures, we demonstrate that both can achieve minimax-optimal rates under certain conditions. This reveals that transformers are not the only viable architecture foundation for ICL, which is also corroborated by recent empirical studies (Tong and Pehlevan, 2025; Kratsios and Furuya, 2025).
2 Related Work
We briefly review related work on ICL and theoretical analyses of transformer architectures across various settings.
2.1 Theoretical Analyses of Transformers in ICL
Recent theoretical studies have rigorously analyzed various aspects of transformer-based models in ICL. For linear function classes, the statistical complexity and representation power of transformers have been precisely characterized (Zhang et al., 2024; Wu et al., 2024; Cheng et al., 2024), with further extensions to mixture models and broader function classes explored in (Jin et al., 2024; Guo et al., 2024; Collins et al., 2024). In terms of optimization, convergence properties and landscape analyses of transformer-based ICL objectives have been investigated primarily under simplified settings, focusing on linear attention and regression tasks (Zhang et al., 2024; Huang et al., 2024; Kim et al., 2024; Nichani et al., 2024; Kim and Suzuki, 2024). More recently, Chen et al. (2024a) have studied single-layer transformers with multi-head attention in nonlinear regimes, although their analysis remains restricted to shallow architectures and regression. In contrast, we establish expected excess risk bounds for deep transformers with multi-head attention in the multiclass classification setting, thereby substantially extending existing theoretical results.
2.2 Mechanistic and Algorithmic Perspectives of Transformers in ICL
Recent work has revealed that transformers can emulate classical learning algorithms such as gradient descent and its variants (Akyürek et al., 2023; Von Oswald et al., 2023; Bai et al., 2023; Fu et al., 2024). Studies have also examined global minimizers and critical points that emerge from adaptive gradient descent methods within transformer architectures (Ahn et al., 2023; Mahankali et al., 2024). Additionally, Bayesian analyses suggest that transformers implicitly approximate Bayesian inference, further enhancing the interpretability of their ICL mechanisms (Xie et al., 2022; Müller et al., 2022; Zhang et al., 2022; Panwar et al., 2024; Jeon et al., 2024). Transformers’ ability to perform ICL on data generated by Markov chains has also been investigated (Collins et al., 2024; Edelman et al., 2024; Makkuva et al., 2024), as has their potential for in-context decision-making within reinforcement learning frameworks (Lin et al., 2024; Sinii et al., 2024). Broader theoretical perspectives concerning the generalization behavior, algorithmic interpretations, and meta-learning properties of transformer-based models have been extensively studied by Li et al. (2023), Dai et al. (2023), and Raventós et al. (2023). Analyses of multi-head attention mechanisms, from both representational and Bayesian viewpoints, have further illuminated the learning dynamics in these models (An et al., 2020; Mahdavi et al., 2024). A detailed discussion of these results, however, lies beyond the scope of this paper.
2.3 ICL Beyond Transformers
Recent literature has begun to investigate whether models beyond transformers can exhibit in-context learning behavior. Tong and Pehlevan (2025) empirically demonstrated that multi-layer perceptrons (MLPs) can achieve ICL performance comparable to transformers on synthetic tasks. Complementing this, Kratsios and Furuya (2025) proved that MLPs possess universal approximation capabilities in an ICL setting. Our theoretical results further support these empirical and theoretical findings by showing that, under comparable conditions, MLPs attain minimax-optimal rates for multiclass classification. This provides evidence that attention-based architectures may not be essential for ICL.
3 Preliminaries
Notation
We write $n]:=\{1,2,\ldots,n\}$. The set of natural numbers including zero is denoted by $\mathbb{N}:=\mathbb{Z}_{\geq 0}$. For a vector $\mathbf{x}=(x_{1},\cdots,x_{p})^{\top}\in\mathbb{R}^{p}$ and a matrix $\mathbf{X}=(X_{ij})\in\mathbb{R}^{p\times N}$, we define the entrywise max-norms $\|\mathbf{x}\|_{\infty}:=\max_{i}|x_{i}|$ and $\|\mathbf{X}\|_{\infty}:=\max_{i,j}|X_{ij}|$, the Euclidean norm $\|\mathbf{x}\|:=(\sum_{i}x_{i}^{2})^{1/2}$, and the Frobenius norm $\|\mathbf{X}\|_{F}:=(\sum_{i,j}X_{ij}^{2})^{1/2}$. The counting norm $\|\mathbf{x}\|_{0}$ denotes the number of nonzero entries in $\mathbf{x}$, and analogously for matrices. For vector-valued functions $\mathbf{f},\mathbf{g}:\Omega\subset\mathbb{R}^{d}\to\mathbb{R}^{K}$, we write $\|\mathbf{f}-\mathbf{g}\|_{\infty}:=\sup_{\mathbf{z}\in\Omega}\max_{j\in[K]}|f_{j}(\mathbf{z})-g_{j}(\mathbf{z})|$, omitting $\Omega$ when clear. For vectors $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{K}$, define $\log(\mathbf{x}/\mathbf{x}^{\prime}):=\log(\mathbf{x})-\log(\mathbf{x}^{\prime})$, applied coordinatewise. For any $\in\mathbb{R}, x\lfloor x\rfloor$ and $\lceil x\rceil$ denote the floor and ceiling, respectively. A standard basis vector in $\mathbb{R}^{K}$ has the form $(0,\ldots,0,1,0,\ldots,0)^{\top}$. The (K1)(K{-}1)-simplex is $\mathcal{S}^{K}:=\{\mathbf{p}\in\mathbb{R}^{K}:\sum_{k}p_{k}=1,\;p_{k}\geq 0\}$. For probability measures $P$ and $Q$, the Kullback–Leibler divergence is $\mathrm{KL}(P\|Q):=\int\log(dP/dQ)\,dP$ if $P\ll Q$, and $\infty$ otherwise. We use the notation $f(x)\lesssim g(x)$ to signify that $f(x)\leq Cg(x)$ for some constant $C>0$. Similarly, $f(x)\gtrsim g(x)$ signifies $f(x)\geq Cg(x)$. The hardmax operator is $\sigma_{H}(\mathbf{x}):=\lim_{c\to\infty}\exp(c\mathbf{x})/\|\exp(c\mathbf{x})\|_{1}$, applied columnwise for matrices. The Hadamard product $\odot$ denotes entrywise multiplication. The metric entropy of the function class $\mathcal{F}$ with respect to the metric $\rho$ is defined as $\mathcal{V}(\delta,\mathcal{F},\rho):=\log\mathcal{N}(\delta,\mathcal{F},\rho)$, where $\mathcal{N}$ is the $\delta$-covering number of $\mathcal{F}$ under the metric $\rho$. Throughout this paper, we use $C$ to denote a generic constant, which may differ from line to line.
3.1 In-Context Learning
In-context learning (ICL) refers to the phenomenon where a model performs prediction by conditioning on a sequence of input-output examples, known as a prompt, without updating its parameters (Brown et al., 2020; Garg et al., 2022). A prompt consists of a sequence $(\mathbf{x}_{1},\mathbf{y}_{1},\ldots,\mathbf{x}_{N},\mathbf{y}_{N},\mathbf{x}_{N+1})$, where each $\mathbf{y}_{i}=f(\mathbf{x}_{i})$ for some unknown function $f$, and $\mathbf{x}_{N+1}$ is a query input. The goal is to output a prediction $\hat{\mathbf{y}}(\mathbf{x}_{N+1})$ such that $\hat{\mathbf{y}}(\mathbf{x}_{N+1})\approx f(\mathbf{x}_{N+1})$.
In ICL, each task is specified by a joint distribution $P$ over $(\mathbf{x},\mathbf{y})\in[0,1]^{p}\times[K]$, where $P$ is drawn from a family $\mathcal{P}$. For a given task, the model observes $N$ i.i.d. examples from $P$ as a prompt and is then required to predict the label $\mathbf{y}_{N+1}$ for a query input $\mathbf{x}_{N+1}$, where $(\mathbf{x}_{N+1},\mathbf{y}_{N+1})\sim P$ independently of the prompt.
For a given task, we draw $(\mathbf{x}_{1},\mathbf{y}_{1}),\ldots,(\mathbf{x}_{N+1},\mathbf{y}_{N+1})\overset{\text{i.i.d.}}{\sim}P$, where the first $N$ pairs form the prompt and $\mathbf{x}_{N+1}$ is the query. Following Bai et al. (2023), we encode the prompt and the query into a matrix $\mathbf{Z}\in\mathbb{R}^{(p+K)\times(N+1)}$ as
$\mathbf{Z}=\begin{bmatrix}\mathbf{X}&\mathbf{x}_{N+1}\\ \mathbf{Y}&\mathbf{o}_{K}\end{bmatrix}$,
where $\mathbf{X}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})\in\mathbb{R}^{p\times N}$, $\mathbf{Y}=(\mathbf{y}_{1},\ldots,\mathbf{y}_{N})\in\mathbb{R}^{K\times N}$, and $\mathbf{o}_{K}\in\mathbb{R}^{K}$ is a zero vector placeholder for the label of the query input $\mathbf{x}_{N+1}$, which the model is expected to predict based on the prompt.
We consider a dataset of $T$ tasks, denoted by $\mathcal{D}_{T}=\{(\mathbf{Z}^{(t)},\mathbf{y}^{(t)}_{N+1})\}_{t=1}^{T}$, where each prompt matrix $\mathbf{Z}^{(t)}\in[0,1]^{(p+K)\times(N+1)}$ is flattened into a vector in $0,1]^{d}$ with $d:=(p+K)(N+1)$ for notational simplicity.
3.2 The Multiclass Classification Model
The core objective of in-context classification is to predict the label of a query input based on a finite set of observed input-output pairs, collectively referred to as the prompt. In the multiclass setting with $K$ possible classes, this task can be formalized as estimating the conditional probability distribution over labels for the query, given the prompt. Estimating the full conditional distribution is essential for both uncertainty quantification and reliable prediction.
Consider task $t$ with prompt $\mathbf{Z}^{(t)}=(\mathbf{X}^{(t)},\mathbf{Y}^{(t)},\mathbf{x}_{N+1}^{(t)})$, where $(\mathbf{X}^{(t)},\mathbf{Y}^{(t)})$ denotes the set of $N$ in-context examples and $\mathbf{x}_{N+1}^{(t)}$ is the query input. Let $\mathbf{y}_{N+1}^{(t)}\in\{0,1\}^{K}$ be the one-hot encoded label of the query, with $y_{N+1,k}^{(t)}$ denoting its $k$-th component. The true conditional class probabilities are defined as
$p_{k}^{0}(\mathbf{Z}^{(t)}):=\mathbb{P}\left(y_{N+1,k}^{(t)}=1\mid\mathbf{Z}^{(t)}\right)=\mathbb{P}\left(y_{N+1,k}^{(t)}=1\mid\mathbf{X}^{(t)},\mathbf{Y}^{(t)},\mathbf{x}_{N+1}^{(t)}\right),\quad k\in[K],$
and the full vector of probabilities is given by
$\mathbf{p}_{0}(\mathbf{Z}^{(t)}):=\left(p_{1}^{0}(\mathbf{Z}^{(t)}),\dots,p_{K}^{0}(\mathbf{Z}^{(t)})\right)\in\mathcal{S}^{K},$
where $\mathcal{S}^{K}$ denotes the $K$-dimensional probability simplex.
To estimate the conditional class probabilities, a standard approach in deep learning is to minimize the empirical negative log-likelihood via (stochastic) gradient descent. Let $\mathcal{D}_{T}=\{(\mathbf{Z}^{(t)},\mathbf{y}_{N+1}^{(t)})\}_{t=1}^{T}$ be a dataset of $T$ tasks, where each $\mathbf{Z}^{(t)}$ denotes the prompt input and $\mathbf{y}_{N+1}^{(t)}\in\{0,1\}^{K}$ is a one-hot encoded label vector. Given a conditional class probability vector $\mathbf{p}(\mathbf{Z})=(p_{1}(\mathbf{Z}),\dots,p_{K}(\mathbf{Z}))$, the likelihood over $\mathcal{D}_{T}$ is defined as $\mathcal{L}(\mathbf{p}\mid\mathcal{D}_{T})=\prod_{t=1}^{T}\prod_{k=1}^{K}\left(p_{k}(\mathbf{Z}^{(t)})\right)^{y_{N+1,k}^{(t)}}$.
Taking the negative logarithm yields the empirical negative log-likelihood (or cross-entropy loss):
$\ell(\mathbf{p},\mathcal{D}_{T})=-\frac{1}{T}\sum_{t=1}^{T}\sum_{k=1}^{K}y_{N+1,k}^{(t)}\log p_{k}(\mathbf{Z}^{(t)})=-\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}_{N+1}^{(t)\,\top}\log\mathbf{p}(\mathbf{Z}^{(t)}),$
where the logarithm is applied element-wise and denotes the transpose. This loss measures the average discrepancy between the predicted and true class distributions across tasks. Minimizing it corresponds to empirical risk minimization under the cross-entropy loss. The resulting estimator $\hat{\mathbf{p}}\in\mathcal{F}$ is obtained by optimizing over a function class $\mathcal{F}$, typically parameterized by a deep neural network and constrained to produce valid probability vectors.
We aim to investigate a non-asymptotic bound of the expected excess risk of the estimator $\hat{\mathbf{p}}$, as measured by the expected cross-entropy loss:
$\mathbb{E}_{\mathcal{D}_{T},(\mathbf{Z},\mathbf{y}_{N+1})}\left[\mathbf{y}_{N+1}^{\top}\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\hat{\mathbf{p}}(\mathbf{Z})}\right]=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[\mathbf{p}_{0}(\mathbf{Z})^{\top}\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\hat{\mathbf{p}}(\mathbf{Z})}\right]=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[{\rm KL}(\mathbf{p}_{0}(\mathbf{Z})\|\hat{\mathbf{p}}(\mathbf{Z}))\right],$
where $(\mathbf{Z},\mathbf{y}_{N+1})$ is independent of the data set $\mathcal{D}_{T}$, $\mathbf{p}_{0}(\mathbf{Z})$ is the true conditional distribution, and $\hat{\mathbf{p}}(\mathbf{Z})$ is the predicted distribution, and ${\rm KL}(\mathbf{p}_{0}(\mathbf{Z})\|\hat{\mathbf{p}}(\mathbf{Z}))$ denotes the Kullback–Leibler (KL) divergence between the two distributions, conditional on $\mathbf{Z}$ and $(\mathcal{D}_{T},\mathbf{Z})$, respectively.
However, the expected KL divergence may diverge when the likelihood ratio becomes unbounded (see Lemma 2.1 in Bos and Schmidt-Hieber (2022)). To address this limitation, following the suggestion of Bos and Schmidt-Hieber (2022), we consider a truncated version of the KL divergence at level $B>0$:
$\mathrm{KL}_{B}(\mathbf{p}_{0}(\mathbf{Z})\,\|\,\hat{\mathbf{p}}(\mathbf{Z}))=\mathbf{p}_{0}(\mathbf{Z})^{\top}\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\hat{\mathbf{p}}(\mathbf{Z})}\right),$
where $a\wedge b=\min\{a,b\}$. Therefore, the expected excess risk of the estimator $\hat{\mathbf{p}}$, with respect to the expected cross-entropy loss, is defined as the expected truncated Kullback-Leibler divergence:
$R_{B}(\mathbf{p}_{0},\hat{\mathbf{p}}):=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[\mathrm{KL}_{B}(\mathbf{p}_{0}(\mathbf{Z})\,\|\,\hat{\mathbf{p}}(\mathbf{Z}))\right]=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[\mathbf{p}_{0}(\mathbf{Z})^{\top}\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\hat{\mathbf{p}}(\mathbf{Z})}\right)\right].$
[…]
$R(\mathbf{p}_{0},\hat{\mathbf{p}})=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[KL(\mathbf{p}_{0}(\mathbf{Z}),\hat{\mathbf{p}}(\mathbf{Z}))\right].$
Due to the non-convexity of the function class $\mathcal{F}$, we introduce a parameter to quantify the gap between the empirical risk achieved by an estimator $\hat{\mathbf{p}}$ and the global minimum within $\mathcal{F}$. Formally, this empirical risk gap is defined as:
$\Delta_{T}(\mathbf{p}_{0},\hat{\mathbf{p}}):=\mathbb{E}_{\mathcal{D}_{T}}\left[-\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}_{N+1}^{(t)\,\top}\log(\hat{\mathbf{p}}(\mathbf{Z}^{(t)}))-\min_{\mathbf{p}\in\mathcal{F}}\left(-\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}_{N+1}^{(t)\,\top}\log(\mathbf{p}(\mathbf{Z}^{(t)}))\right)\right].$ (1)
This quantity measures the expected difference between the empirical risk of $\hat{\mathbf{p}}$ and the optimal empirical risk achievable within the class $\mathcal{F}$.
3.3 Model Classes: Transformers and MLPs
We begin by defining the model classes for our estimator, employing a transformer encoder for prompt representation and a subsequent MLP for processing the encoded output. This hybrid architecture—utilizing a transformer as a feature extractor with an MLP classifier atop—constitutes a prevalent design paradigm in practice. Transformers excel at capturing complex contextual dependencies through self-attention mechanisms, generating enriched latent representations of sequential data. The subsequent MLP leverages these high-level features to perform efficient classification or regression, benefiting from the transformer’s representational capacity while preserving computational tractability. Empirical studies across NLP benchmarks (e.g., BERT+MLP for GLUE tasks) demonstrate that such configurations offer an optimal balance: transformers extract transferable linguistic features, while lightweight MLP heads enable rapid task-specific adaptation without exhaustive retraining. The decoupled design further facilitates modular optimization, allowing domain-specific fine-tuning of the transformer alongside dedicated calibration of the classifier.
Transformer architecture.
Adapting the architecture proposed in previous theoretical work on transformer-based models Gurevych et al. (2022); Bai et al. (2023); Jiao et al. (2024), we adopt a transformer encoder to process the input prompt.
Let $\mathbf{Z}\in\mathbb{R}^{D\times(N+1)}$ denote the input prompt, where $D=p+K$ is the feature dimension and $N+1$ is the in-context length. A transformer network $\boldsymbol{\phi}$ maps $\mathbf{Z}$ to an output of the same dimension through a sequence of $M$ transformer blocks, each comprising a self-attention (SA) and a feed-forward (FF) sublayer. The architecture is defined as:
$\boldsymbol{\phi}=E_{\mathrm{out}}\circ F_{M}^{(\mathrm{FF})}\circ F_{M}^{(\mathrm{SA})}\circ\dots\circ F_{1}^{(\mathrm{FF})}\circ F_{1}^{(\mathrm{SA})}(\mathbf{Z}),$ (2)
where $E_{\mathrm{out}}$ denotes the final output projection.
Each self-attention (SA) sublayer applies a multi-head attention mechanism with a residual connection:
$F^{(\mathrm{SA})}(\mathbf{Z})=\mathbf{Z}+\sum_{s=1}^{H}\mathbf{W}_{O,s}\left(\mathbf{W}_{V,s}\mathbf{Z}\right)\left[\left(\mathbf{W}_{K,s}\mathbf{Z}\right)^{\top}\left(\mathbf{W}_{Q,s}\mathbf{Z}\right)\right]\odot\sigma_{H}\left(\left(\mathbf{W}_{K,s}\mathbf{Z}\right)^{\top}\left(\mathbf{W}_{Q,s}\mathbf{Z}\right)\right),$
where $\sigma_{H}$ denotes the hardmax activation, and each attention head is parameterized by learnable projection matrices $\mathbf{W}_{Q,s},\mathbf{W}_{K,s},\mathbf{W}_{V,s},\mathbf{W}_{O,s}\in\mathbb{R}^{D\times D}$. Each feed-forward (FF) sublayer is a position-wise two-layer network with ReLU activation and residual connection:
$F^{(\mathrm{FF})}(\mathbf{Z})=\mathbf{Z}+\mathbf{W}_{F_{2}}\,\sigma\left(\mathbf{W}_{F_{1}}\mathbf{Z}\right),$
where $\mathbf{W}_{F_{1}}\in\mathbb{R}^{D^{\prime}\times D}$, $\mathbf{W}_{F_{2}}\in\mathbb{R}^{D\times D^{\prime}}$, and $\sigma(\cdot)$ is the ReLU activation function.
To generate a prediction for the query label $\mathbf{y}_{N+1}$, we apply an output projection to the final representation of the (N+1)(N{+}1)-th token:
$E_{\mathrm{out}}(\mathbf{Z})=\mathbf{A}_{\mathrm{out}}\,\mathbf{z}_{N+1}+\mathbf{b}_{\mathrm{out}},$
where $\mathbf{z}_{N+1}\in\mathbb{R}^{D}$ is the last column of $\mathbf{Z}$, and $\mathbf{A}_{\mathrm{out}}\in\mathbb{R}^{K\times D}$, $\mathbf{b}_{\mathrm{out}}\in\mathbb{R}^{K}$ are learnable parameters.
Based on the above structure, we define a class of transformer networks with bounded output, Lipschitz continuity, and sparsity constraints:
[…]
$\ \sup_{\mathbf{Z}}\|\boldsymbol{\phi}(\mathbf{Z})\|\leq B_{1},$
$\ \|\boldsymbol{\phi}(\mathbf{Z}_{1})-\boldsymbol{\phi}(\mathbf{Z}_{2})\|\leq\gamma\|\mathbf{Z}_{1}-\mathbf{Z}_{2}\|_{F},\;\forall\ \mathbf{Z}_{1},\mathbf{Z}_{2}\in[0,1]^{d},$
$\ \sum_{r=1}^{M}\sum_{s=1}^{H}\left(\|\mathbf{W}_{Q,r,s}\|_{0}+\|\mathbf{W}_{K,r,s}\|_{0}+\|\mathbf{W}_{V,r,s}\|_{0}+\|\mathbf{W}_{O,r,s}\|_{0}\right)$
$\ +\sum_{r=1}^{M}\left(\|\mathbf{W}_{r,F_{1}}\|_{0}+\|\mathbf{W}_{r,F_{2}}\|_{0}\right)+\|\mathbf{A}_{\mathrm{out}}\|_{0}+\|\mathbf{b}_{\mathrm{out}}\|_{0}\leq S_{1},$
$\ \max(D,D^{\prime}_{r})\leq W_{1},\ \forall\ r=1,\dots,M\big{\}},$ (3)
where $d=(p+K)(N+1)$ and $\|\cdot\|_{0}$ counts the number of nonzero elements.
MLP architecture.
The class $\mathcal{F}(L,W)$ consists of depth-$L$ ReLU networks with maximum width $W$:
$\mathcal{F}(L,W):=\left\{\mathbf{f}:\mathbb{R}^{K}\to\mathbb{R}^{K}\;\middle|\;\mathbf{f}(\mathbf{x})=\mathbf{W}_{L}\sigma_{\boldsymbol{\nu}_{L}}\cdots\mathbf{W}_{1}\sigma_{\boldsymbol{\nu}_{1}}\mathbf{W}_{0}\mathbf{x}\right\},$ (4)
where $W=\max_{j}K_{j}$, $\mathbf{W}_{j}\in\mathbb{R}^{K_{j+1}\times K_{j}}$, $\boldsymbol{\nu}_{j}\in\mathbb{R}^{K_{j}}$, and $\sigma_{\boldsymbol{\nu}}(\mathbf{x}):=\max(\mathbf{x}-\boldsymbol{\nu},0)$ denotes the biased ReLU activation.
To control the capacity of the MLP class, we define a constrained subclass $\mathcal{F}_{\mathrm{id}}(L,W,S)$ with bounded parameters and sparsity:
$\mathcal{F}_{\mathrm{id}}(L,W,S):=\bigg{\{}\mathbf{f}\in\mathcal{F}(L,W)\Big{|}\max_{j\in[L]}\left(\|\mathbf{W}_{j}\|_{\infty}\vee\|\boldsymbol{\nu}_{j}\|_{\infty}\right)\leq 1,\sum_{j=0}^{L}\left(\|\mathbf{W}_{j}\|_{0}+\|\boldsymbol{\nu}_{j}\|_{0}\right)\leq S\bigg{\}},$ (5)
where $\|\cdot\|_{\infty}$ is the element-wise maximum norm.
Assumption 1
Assume that the MLP class $\mathcal{F}_{\mathrm{id}}(L,W,S)$ is $L_{1}$-Lipschitz with respect to the $\ell_{\infty}$ norm, i.e.,
$\|\mathbf{f}(\mathbf{y}_{1})-\mathbf{f}(\mathbf{y}_{2})\|_{\infty}\leq L_{1}\|\mathbf{y}_{1}-\mathbf{y}_{2}\|_{\infty},\quad\forall\ \mathbf{f}\in\mathcal{F}_{\mathrm{id}}(L,W,S),\ \forall\ \mathbf{y}_{1},\mathbf{y}_{2}\in\mathbb{R}^{K}.$
This assumption simplifies our analysis. Although the Lipschitz constant $L_{1}$ of an MLP can be bounded in terms of its depth, width, and weight norms, computing its exact value is NP-hard (Virmaux and Scaman, 2018). Moreover, obtaining tight upper bounds remains challenging (Fazlyab et al., 2019; Latorre et al., 2020; Kim et al., 2021). Therefore, we do not attempt to estimate it explicitly and instead assume a uniform bound $L_{1}$.
3.4 Function Class for Conditional Probability Estimation
We now define the full function class used to estimate conditional class probabilities from a given prompt. Each function integrates a transformer encoder for contextual representation learning, followed by an MLP that maps the latent representation to class probabilities via a final softmax layer:
$\mathcal{F}(M,H,D,W_{1},B_{1},S_{1},\gamma,L,W,S)$
$:=\left\{\mathbf{p}(\mathbf{Z})=\boldsymbol{\Phi}\circ\mathbf{f}\circ\boldsymbol{\phi}(\mathbf{Z})\;\middle|\;\mathbf{f}\in\mathcal{F}_{\mathrm{id}}(L,W,S),\ \boldsymbol{\phi}\in\mathcal{T}(M,H,D,W_{1},B_{1},S_{1},\gamma)\right\},$ (6)
where $\boldsymbol{\Phi}(\cdot)$ is the softmax function.
4 The Risk Bounds for In-Context Learning
In this section, we establish risk bounds for ICL in the multiclass classification setting. We first introduce some key definitions and assumptions used in our analysis.
Definition 1 (Small Value Bound)
Let $\alpha\geq 0$ and $\mathcal{H}$ be a function class. A function $\mathbf{p}=(p_{1},\cdots,p_{K})\in\mathcal{H}$ is said to be $\alpha$-small value bounded (or $\alpha$-SVB) if there exists some constant $C>0$ such that for any $k\in[K]$ and $t\in(0,1]$, the following holds:
$\mathbb{P}\left\{p_{k}(\mathbf{Z})\leq t\right\}\leq Ct^{\alpha}.$
Moreover, if all $\mathbf{p}\in\mathcal{H}$ are $\alpha$-SVB, we say that $\mathcal{H}$ is $\alpha$-SVB.
The concept of $\alpha$-SVB generalizes the standard assumption in nonparametric classification that true conditional class probabilities are bounded away from zero, allowing them to approach zero. This extension, which follows the conditions outlined in Bos and Schmidt-Hieber (2022), tightens error bounds and accelerates convergence. The index $\alpha$ will appear in the convergence rate analysis.
Definition 2 (Hölder Smoothness)
Let $\beta>0$, $Q>0$, and $\Omega\subset\mathbb{R}^{d}$. The ball of $\beta$-Hölder continuous functions with radius $Q$, denoted by $C^{\beta}(\Omega,Q)$, is defined as
$C^{\beta}(\Omega,Q):=\left\{f:\Omega\to\mathbb{R}:\sum_{\|\gamma\|_{1}<\beta}\|\partial^{\gamma}f\|_{\infty}+\sum_{\|\gamma\|_{1}=\lfloor\beta\rfloor}\sup_{\begin{subarray}{c}\mathbf{x},\mathbf{y}\in\Omega\\ \mathbf{x}\neq\mathbf{y}\end{subarray}}\frac{|\partial^{\gamma}f(\mathbf{x})-\partial^{\gamma}f(\mathbf{y})|}{\|\mathbf{x}-\mathbf{y}\|_{\infty}^{\beta-\lfloor\beta\rfloor}}\leq Q\right\},$
where $\gamma=(\gamma_{1},\dots,\gamma_{d})\in\mathbb{N}^{d}$ is a multi-index and $\partial^{\gamma}:=\partial^{\gamma_{1}}_{x_{1}}\cdots\partial^{\gamma_{d}}_{x_{d}}$.
Definition 3 (Hölder-Smooth Class of Probability)
For any $\beta>0$ and $Q>0$, the $\beta$-Hölder-smooth class of the conditional probability functions is defined as
$\mathcal{G}(\beta,Q):=\left\{\mathbf{p}_{0}=(p_{1}^{0},\dots,p_{K}^{0}):[0,1]^{d}\to\mathcal{S}^{K},\;p_{k}^{0}\in C^{\beta}([0,1]^{d},Q)\text{ for all }k\in[K]\right\},$ (7)
where $\mathcal{S}^{K}\subset\mathbb{R}^{K}$ denotes the probability simplex. Furthermore, we write $\mathcal{G}_{\alpha}(\beta,Q)$ for all functions in $\mathcal{G}(\beta,Q)$ that are $\alpha$-SVB.
For a class of conditional class probabilities $\mathcal{F}$, we define its coordinate-wise logarithmic transform as
$\log(\mathcal{F}):=\left\{\log\mathbf{p}=(\log p_{1},\dots,\log p_{K}):\mathbf{p}\in\mathcal{F}\right\}.$
Definition 4 ($\delta$-Covering Number and $\delta$-Entropy)
Let $(\mathcal{V},d)$ be a pseudometric space with metric $d(\cdot,\cdot)$ and $\mathcal{F}$ be a subset of $\mathcal{V}$. A finite set $\mathcal{F}(\delta)$ is called a $\delta$-cover (or $\delta$-net) of $\mathcal{F}$ if, for all $\mathbf{f}\in\mathcal{F}$, there exists $\mathbf{g}\in\mathcal{F}(\delta)$ such that $d(\mathbf{f},\mathbf{g})\leq\delta$. Moreover, the $\delta$-covering number of $\mathcal{F}$ with respect to the metric $d$ is the smallest cardinality of such a $\delta$-cover $\mathcal{F}(\delta)$, denoted by $\mathcal{N}(\delta,\mathcal{F},d)$. The quantity $\log\mathcal{N}(\delta,\mathcal{F},d)$ is referred to as the $\delta$-entropy of $\mathcal{F}$ with respect to the metric $d$.
Definition 5 (Empirical and Uniform Supremum Covering Number)
Given $\mathcal{Z}_{T}=\{\mathbf{Z}^{(1)},\ldots,\mathbf{Z}^{(T)}\}\subset\mathbb{R}^{d}$, we define the pseudometric $d=\|\cdot\|_{\infty,\mathcal{Z}_{T}}$ as
$d(\mathbf{f},\mathbf{g}):=\max_{1\leq t\leq T}\max_{1\leq k\leq K}\left|\mathbf{f}_{k}(\mathbf{Z}^{(t)})-\mathbf{g}_{k}(\mathbf{Z}^{(t)})\right|.$
The corresponding metric covering number $\mathcal{N}(\delta,\mathcal{F},\|\cdot\|_{\infty,\mathcal{Z}_{T}})$ is referred as the empirical sup-norm covering number. For a fixed $T$, the largest sup-norm covering number over all $\mathcal{Z}_{T}\subset\mathbb{R}^{d}$ is referred to as the uniform empirical sup-norm covering number:
$\mathcal{N}_{\infty}(\delta,\mathcal{F},T):=\sup_{\mathcal{Z}_{T}}\mathcal{N}(\delta,\mathcal{F},\|\cdot\|_{\infty,\mathcal{Z}_{T}}).$
Assumption 2
Let $\mathcal{F}$ be a class of conditional class probability functions $\mathbf{p}=(p_{1},\dots,p_{K})$ used to estimate conditional class probabilities. Suppose that for any $\mathbf{p}\in\mathcal{F}$, there exist two constants $C_{1}>0$ and $0<C_{2}<1$ such that $p_{k}(\mathbf{Z})\geq C_{1}/T^{C_{2}}$ for all $k\in[K]$ and $\mathbf{Z}\in[0,1]^{d}$.
Assumption 2 requires all estimated probabilities to be bounded away from zero. This avoids the singularity of the KL divergence and enables the use of empirical covering numbers to control the complexity of $\log\mathcal{F}$ in the oracle inequality.
Remark 6
[…]
Theorem 7 (Oracle Inequality)
Let $\mathcal{F}$ be any class of conditional class probability functions that satisfies Assumption 2, and let $\widehat{\mathbf{p}}\in\mathcal{F}$ be any estimator of $\mathbf{p}_{0}\in\mathcal{S}^{K}$. For any truncation level $B\geq 2$ and task number $T$, the expected excess risk satisfies
$R_{B}(\mathbf{p}_{0},\widehat{\mathbf{p}})\leq C\left(\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})+\Delta_{T}(\mathbf{p}_{0},\widehat{\mathbf{p}})+\frac{B^{2}\log(\mathcal{N}_{T})\,\log^{2}T}{T}\right),$ (8)
[…]
Theorem 7 provides a general oracle inequality for any function class $\mathcal{F}$ satisfying Assumption 2. This result is independent of the specific architecture used to model the conditional class probabilities. In the subsequent analysis, we instantiate $\mathcal{F}$ as a class of neural networks that estimate conditional class probabilities from a given prompt, as defined in (3.4).
Error Decomposition.
[…]
To derive the ICL risk, we focus on two key components: the approximation error $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$ and the uniform empirical covering entropy $\mathcal{N}_{T}$ in the oracle inequality (8). We first establish the approximation capacity of transformer networks for functions in the Hölder class $\mathcal{G}(\beta,Q)$ in the following lemma.
Lemma 8
Let $0<\varepsilon<1$, $\beta>0$, and $Q>0$. For every function $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$, there exists a transformer network $\boldsymbol{\phi}\in\mathcal{T}(M,H,D,W_{1},B_{1},S_{1},\gamma)$ with
$M\lesssim\log\left(\frac{Q}{\varepsilon}\right),\quad H\lesssim\left(\frac{Q}{\varepsilon}\right)^{d/\beta},\quad D\lesssim 1,$
$W_{1}\lesssim\left(\frac{Q}{\varepsilon}\right)^{d/\beta},\quad B_{1}\lesssim\|\mathbf{p}_{0}\|_{\infty},\quad S_{1}\lesssim\left(\frac{Q}{\varepsilon}\right)^{d/\beta}\log\left(\frac{Q}{\varepsilon}\right),$
such that $\|\boldsymbol{\phi}-\mathbf{p}_{0}\|_{\infty}\leq\varepsilon$. Furthermore, if $\beta>1$, we may choose $\gamma\lesssim Q$.
Lemma 8 indicates that transformer networks can uniformly approximate any function in the Hölder class $\mathcal{G}(\beta,Q)$ with explicit control over architectural complexity. Similar approximation results have been established in related settings, such as Theorem 8 in Jiao et al. (2024). Its proof builds on the observation by Gurevych et al. (2022) that transformer architectures, such as deep ReLU networks, can represent multiplication operations ans thus approximate polynomials. Since polynomials can can approximate continuous functions arbitrarily well, transformer networks can effectively approximate Hölder-smooth targets. Without loss of generality, we assume the output of the approximating transformer lies in $[0,1]$; otherwise, a projection layer can be applied without increasing the error, as the target functions $p_{k}^{0}$ are already supported on this interval.
Next, we apply Lemma 8 to approximate the conditional class probabilities in multiclass classification. Specifically, we consider the neural network class $\mathcal{F}(M,H,D,W_{1},B_{1},S_{1},\gamma,L,\\ W,S)$ defined in (3.4), which is constructed to approximate the ground-truth conditional class probabilities $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$ in the $\|\cdot\|_{\infty}$-norm, while ensuring that the output probability vectors are uniformly bounded away from zero. This is achieved by composing a transformer encoder with an MLP followed by a softmax transformation. The MLP processes the representations produced by the transformer encoder and computes a vector of logits, which are then mapped by the softmax function to class probability vectors. The resulting outputs are uniformly bounded away from zero and thus satisfy Assumption 2. A similar construction is used in Bos and Schmidt-Hieber (2022).
This transformer-MLP hybrid architecture is also motivated by its practical effectiveness. Transformers provide expressive, context-aware representations through self-attention mechanisms, while the MLP acts as an adaptive task-specific output head. Such modular designs are widely used in contemporary deep learning models and have demonstrated strong generalization performance across diverse applications. For simplicity, we henceforth denote the neural network classes $\mathcal{T}(M,H,D,W_{1},B_{1},S_{1},\gamma)$ defined in (3.3), $\mathcal{F}_{\mathrm{id}}(L,W,S)$ defined in (5), and $\mathcal{F}(M,H,D,W_{1},B_{1},S_{1},\gamma,L,W,S)$ defined in (3.4) as $\mathcal{T}$, $\mathcal{F}_{\mathrm{id}}$, and $\mathcal{F}$, respectively.
Theorem 9
Fix $\beta>0$ and $Q>0$. For any $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$ and any $\varepsilon>0$, there exists a neural network $\mathbf{\tilde{q}}=(\tilde{q}_{1},\cdots,\tilde{q}_{K})\in\mathcal{F}$ with
$M\lesssim\log\left(\frac{Q}{\varepsilon}\right),\quad H\lesssim\left(\frac{Q}{\varepsilon}\right)^{d/\beta},\quad D\lesssim 1,$
$W_{1}\lesssim\left(\frac{Q}{\varepsilon}\right)^{d/\beta},\quad B_{1}\lesssim\|\mathbf{p}\|_{\infty},\quad S_{1}\lesssim\left(\frac{Q}{\varepsilon}\right)^{d/\beta}\log\left(\frac{Q}{\varepsilon}\right),$
$L\lesssim\log\left(\frac{1}{\varepsilon}\right),\quad W\lesssim K\left(\frac{1}{\varepsilon}\right)^{1/\beta},\quad S\lesssim K\left(\frac{1}{\varepsilon}\right)^{1/\beta}\log\left(\frac{1}{\varepsilon}\right),$
such that
$\|{\tilde{q}}_{k}-{p}^{0}_{k}\|_{\infty}<2\varepsilon(K+1),\quad{\tilde{q}}_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2(K+1)}$
hold for all $k\in\{1,2,\dots,K\}$ and $\mathbf{Z}\in[0,1]^{d}$.
Theorem 9 shows that the conditional probability $\mathbf{p}_{0}$ can be uniformly approximated, with arbitrarily small error, by a neural network from $\mathcal{F}$. Its proof is provided in Appendix B. Based on this result, we derive a non-asymptotic upper bound for the minimal risk $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$, under uniform approximation and margin conditions.
Remark 10
The neural network $\mathbf{\tilde{q}}$ constructed in Theorem 9 inherently satisfies a natural lower bound on its output for every function $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$. Specifically, for any $\varepsilon>0$, the network’s outputs are bounded below by $\frac{\varepsilon}{2(K+1)}$, which is an intrinsic property of the architecture. This lower bound aligns with Assumption 2. In particular, by selecting an appropriate $\varepsilon$, the neural network $\mathbf{\tilde{q}}$ can be guaranteed to meet the lower bound condition on the probability functions, as specified by Assumption 2.
Theorem 11
Assume $\mathbf{p}_{0}\in\mathcal{G}_{\alpha}(\beta,Q)$ for some $\alpha\in[0,1]$. Let $\mathbf{p}=(p_{1},\ldots,p_{K}):[0,1]^{d}\to\mathcal{S}^{K}$ be an approximation of $\mathbf{p}_{0}$ such that
$\|\mathbf{p}-\mathbf{p}_{0}\|_{\infty}\leq 2(K+1)\varepsilon,\quad\min_{1\leq k\leq K}\inf_{\mathbf{Z}\in[0,1]^{d}}p_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2(K+1)}.$
Then we have
$R(\mathbf{p}_{0},\mathbf{p})\leq C\varepsilon^{1+\alpha}K^{4+\alpha}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right).$
If we choose $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$ with $C>0$ being a sufficiently small constant, then the risk $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$ is bounded as
$\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})\lesssim T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}K^{4+\alpha}\log(KT).$
Theorem 11 establishes a near-optimal convergence rate for in-context classification, recovering the classical nonparametric minimax bound when $\alpha=1$. The proof leverage $f$-divergence theory, bounding the KL divergence via the $\chi^{2}$-divergence (Nguyen et al., 2009; Cai and Lim, 2022). More details can be found in Appendix C.
Next, we deal with the uniform empirical covering entropy $\log(\mathcal{N}_{T})$ in the Oracle Inequality. Recall that
$\log(\mathcal{N}_{T})=\mathcal{V}_{\infty}\Big{(}\frac{1}{T},\log\mathcal{F},T\Big{)}=\sup_{\mathcal{Z}_{T}}\log\mathcal{N}\Big{(}\frac{1}{T},\log\mathcal{F},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\Big{)}.$
The following lemma provides an upper bound on the uniform empirical covering entropy $\log(\mathcal{N}_{T})$ of the class $\log(\mathcal{F})$ under the empirical sup-norm.
Lemma 12
For any $\delta>0$, and for sufficiently small $\delta$, the following inequality holds:
$\mathcal{V}_{\infty}\left(\delta,\log\mathcal{F},T\right)\leq\mathcal{V}_{\infty}\left(\tfrac{\delta}{4KL_{1}},\,\mathcal{T},\,T\right)+\mathcal{V}\left(\tfrac{\delta}{4K},\,\mathcal{F}_{\mathrm{id}},\,\|\cdot\|_{\infty}\right),$
where $\mathcal{V}_{\infty}\left(\delta,\,\mathcal{T},\,T\right)=\sup_{\mathcal{Z}_{T}}\log\mathcal{N}(\delta,\mathcal{T},\|\cdot\|_{\infty,\mathcal{Z}_{T}})$, $\mathcal{V}(\delta,\mathcal{F}_{\rm id},\|\cdot\|_{\infty})=\log\mathcal{N}(\delta,\mathcal{F}_{\rm id},\|\cdot\|_{\infty})$, and $L_{1}$ is the Lipschitz constant specified in Assumption 1.
The proof of Lemma 12 is provided in Appendix D. The argument decomposes $\log\mathcal{F}$ into a composition of the transformer class $\mathcal{T}$ and the MLP class $\mathcal{F}_{\mathrm{id}}$, and then applies standard covering number bounds for Lipschitz compositions to control the resulting entropy.
The next result provides an upper bound on the uniform empirical covering entropy of the transformer class under the empirical sup-norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$, where $\mathcal{Z}_{T}=\{\mathbf{Z}^{(1)},\dots,\mathbf{Z}^{(T)}\}$.
Lemma 13
For any $\delta_{1}>0$, the uniform empirical covering entropy of the transformer class $\mathcal{T}$ under the empirical sup-norm satisfies
$\mathcal{V}_{\infty}\left(\delta_{1},\mathcal{T},T\right)\lesssim KM^{2}S_{1}\log\left(\max\{M,H,W_{1}\}\right)\log\left(\frac{TB_{1}}{\delta_{1}}\right),$
where $\mathcal{T}=\mathcal{T}(M,H,D,W_{1},B_{1},S_{1},\gamma)$.
The proof of Lemma 13 builds on classical techniques from the neural network literature (Anthony and Bartlett, 2009; Bartlett et al., 2019), and it is closely related to recent analyses of transformer architectures (Gurevych et al., 2022; Jiao et al., 2024). A complete proof is provided in Appendix E.
Lemma 14
For any $\delta_{2}>0$, the covering entropy of the function class $\mathcal{F}_{\mathrm{id}}$ under the sup-norm $\|\cdot\|_{\infty}$ satisfies the following upper bound
$\mathcal{V}\left(\delta_{2},\,\mathcal{F}_{\mathrm{id}},\,\|\cdot\|_{\infty}\right)\leq 2SL\log\left(\frac{L(W+1)}{3\delta_{2}}\right),$
where $\mathcal{F}_{\mathrm{id}}=\mathcal{F}_{\mathrm{id}}(L,W,S)$ is defined in (5).
Lemma 14 gives an upper bound on the uniform covering entropy of the class $\mathcal{F}_{\mathrm{id}}(L,W,S)$, which consists of $S$-sparse ReLU networks with depth $L$, width $W$, and output dimension $K$, measured under the sup-norm over the input domain. Its proof is provided in Appendix F.
Theorem 15
Assume $\mathbf{p}_{0}\in\mathcal{G}_{\alpha}(\beta,Q)$ for some $\alpha\in[0,1]$. Let $\hat{\mathbf{p}}$ be an estimator of $\mathbf{p}_{0}$, taking values in the function class $\mathcal{F}(M,H,D,W_{1},B_{1},S_{1},\gamma,L,W,S)$, where the parameters of the class scale with the precision parameter $\varepsilon>0$ as follows:
$M\lesssim\log\left(\frac{Q}{\varepsilon}\right),\ H\lesssim\left(\frac{Q}{\varepsilon}\right)^{d/\beta},\ D\lesssim 1,$
$W_{1}\lesssim\left(\frac{Q}{\varepsilon}\right)^{d/\beta},\ B_{1}\lesssim\|\mathbf{p}_{0}\|_{\infty},\ S_{1}\lesssim\left(\frac{Q}{\varepsilon}\right)^{d/\beta}\log\left(\frac{Q}{\varepsilon}\right),$
$L\lesssim\log\left(\frac{1}{\varepsilon}\right),\ W\lesssim K\left(\frac{1}{\varepsilon}\right)^{1/\beta},\ S\lesssim K\left(\frac{1}{\varepsilon}\right)^{1/\beta}\log\left(\frac{1}{\varepsilon}\right).$
[…]
$\Delta_{T}(\mathbf{p}_{0},\widehat{\mathbf{p}})\lesssim T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}K^{4+\alpha},$
and we choose $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$ for some constant $C>0$, then the ICL risk is bounded by
$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}\right)\lesssim B^{2}T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}K^{4+\alpha}\log^{5}\left(KT\right).$
In particular, for the special case where $\alpha=1$, the ICL risk achieves the bound
$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}\right)\lesssim B^{2}T^{-\frac{2\beta}{2\beta+d}}K^{5}\log^{5}\left(KT\right).$
Theorem 15 establishes that the ICL risk attains the minimax-optimal convergence rate (up to logarithmic factors) when $\alpha=1$. This result highlights the importance of pretraining task diversity, as observed by Raventós et al. (2023). Our setup corresponds to the task-scaling regime formalized by Abedsoltan et al. (2024), where the risk decreases as the number of pretraining tasks $T$ increases, while the context length remains fixed. In our multiclass setting, the achievable risk is fundamentally determined by the number of pretraining tasks $T$. Since large language models are trained on virtually unlimited tasks, our bound provides a theoretical explanation for their strong in-context learning performance—even with only a small number of in-context examples at inference time.
Proof of Theorem 15. The proof is based on the oracle inequality (8), which decomposes the error into an approximation error $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$, controlled by Theorem 11, and a complexity term $(\log(\mathcal{N}_{T})\log^{2}T)/T$, bounded using Lemmas 12, 13, and 14, with $\delta=1/T$. The approximation scale $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$ is chosen to balance the approximation error and the entropy term, reflecting the trade-off between network expressiveness and statistical complexity. These bounds together yields our result. $\blacksquare$
5 Lower Bounds
To establish the optimality of the convergence rate in Theorem 15, we derive a matching minimax lower bound for the ICL risk. Our argument leverages the classical minimax theory for nonparametric estimation under the squared Hellinger distance translate these results to the KL divergence setting.
The squared Hellinger distance, defined as $H^{2}(P,Q)=\frac{1}{2}\int(\sqrt{dP}-\sqrt{dQ})^{2}$, is a fundamental tool for deriving statistical lower bounds due to its favorable metric properties (van der Vaart and Wellner, 1996; Geer, 2000). The rate $T^{-\frac{2\beta}{2\beta+d}}$ is known to be minimax-optimal for estimation under the squared Hellinger loss (Tsybakov, 2008, Chapter 2). The key to translating this result is the classical inequality $2H^{2}(P,Q)\leq\text{KL}(P,Q)$, which holds under bounded likelihood ratio conditions, see (7.5) in Birgé and Massart (1998) and Lemma 4 in Haussler and Opper (1997). Conversely, under some integrability assumptions, the KL divergence can be bounded by the Hellinger distance up to logarithmic factors (Wong and Shen, 1995, Theorem 5). This two-sided relationship enables the direct translation of minimax lower bounds from the Hellinger distance to the KL divergence. The following lemma formalizes this connection for the truncated KL divergence.
Lemma 16
Suppose $B\geq 2$. Then the squared Hellinger distance and the truncated KL divergence satisfy
$H^{2}(P,Q)\leq\tfrac{1}{2}\mathrm{KL}_{2}(P,Q)\leq\tfrac{1}{2}\mathrm{KL}_{B}(P,Q)\leq 5e^{B/2}H^{2}(P,Q).$
[…]
Theorem 17
Consider the same setting as in Theorem 15. Then, the minimax lower bound of the ICL risk is given by
$\inf_{\hat{\mathbf{p}}}\sup_{\mathbf{p}_{0}\in\mathcal{G}_{1}(\beta,Q)}R_{B}(\mathbf{p}_{0},\hat{\mathbf{p}})\gtrsim T^{-\frac{2\beta}{2\beta+d}}.$
This result aligns with the upper bound in Theorem 15 up to constant factors, thereby establishing the minimax-optimality of the rate $T^{-\frac{2\beta}{2\beta+d}}$ for ICL under the truncated KL loss when $\alpha=1$. The case of general $\alpha\in(0,1)$ remains open and will be studied in future work.
6 In-Context Learning Beyond Transformers
ICL is often regarded as a distinctive capability of transformer architectures, with explanations frequently grounded in constructions based on attention mechanisms (Akyürek et al., 2023; Von Oswald et al., 2023; Zhang et al., 2024; Reddy, 2024; Lu et al., 2024).
However, recent literature suggests that ICL is not exclusive to transformer architectures. Tong and Pehlevan (2025) empirically demonstrated that MLPs can also learn in-context on standard synthetic ICL taks. Similarly, Abedsoltan et al. (2024) showed that standard MLPs are capable of task-scaling in ICL settings. Moreover, Kratsios and Furuya (2025) theoretically established that MLPs possess universal approximation capabilities for ICL tasks.
In this section, we support these insights by analyzing the expected excess risk of MLP-based classifiers in the setting of multiclass ICL. Under appropriate assumptions, we demonstrate that MLPs can achieve minimax-optimal convergence rates for ICL. Our analysis of convergence rates provides theoretical support for recent empirical evidence on the effectiveness of MLPs. This finding suggests that, while attention mechanisms are powerful, they are not strictly necessary for enabling ICL.
Specifically, we consider a class of MLPs with ReLU activation functions, used as probabilistic classifiers. To accommodate input prompts in $\mathbb{R}^{d}$, we modify the network architecture introduced in (4) by increasing its input dimension from $K$ to $d$. Accordingly, we introduce a function class as follows:
$\mathcal{F}_{\rm MLP}(L,W,S):=\left\{\tilde{\mathbf{p}}(\mathbf{Z})=\boldsymbol{\Phi}\circ\mathbf{f}(\operatorname{vec}(\mathbf{Z}))\;\middle|\;\mathbf{f}\in\mathcal{F}_{\mathrm{id}}(L,W,S)\right\},$ (9)
where $\boldsymbol{\Phi}$ denotes the softmax function.
Lemma 18
For every $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$ and $0<\varepsilon<\min\{(K(4+C_{Q,\beta,d}))^{-1},(\beta+1)^{-\beta},(Q+1)^{-\beta/d}e^{-\beta}\}$ with $C_{Q,\beta,d}=(2Q+1)(1+d^{2}+\beta^{2})6^{d}+Q3^{\beta}$, there exist a neural network $\tilde{\mathbf{p}}\in\mathcal{F}_{\rm MLP}(L,W,S)$ with
$L\lesssim\log\left(\frac{1}{\varepsilon}\right),\quad W\lesssim K\varepsilon^{-d/\beta},\quad S\lesssim K\varepsilon^{-d/\beta}\log\left(\frac{1}{\varepsilon}\right),$
such that
$\tilde{p}_{k}\geq\frac{\varepsilon}{4},\quad\|\tilde{p}_{k}-p_{k}^{0}\|_{\infty}\leq\frac{1}{2}K(C_{Q,\beta,d}+4)\varepsilon,$
for all $k=1,\dots,K$.
This result slightly modifies Lemma 4.3 in Bos and Schmidt-Hieber (2022), which builds upon a line of work showing that ReLU MLPs can efficiently approximate $\beta$-Hölder smooth functions, with $L\lesssim\log(1/\varepsilon)$, $W\lesssim\varepsilon^{-d/\beta}$, $s\lesssim\varepsilon^{-d/\beta}\log(1/\varepsilon)$ (Yarotsky, 2017; Petersen and Voigtlaender, 2018; Schmidt-Hieber, 2020; Jiao et al., 2023; Chakraborty and Bartlett, 2024b).
Lemma 19
Consider the function class $\mathcal{F}_{\mathrm{MLP}}(L,W,S)$ defined in (9). For any $T\in\mathbb{N}$, the empirical covering entropy of the function class $\log(\mathcal{F}^{\prime}_{\Phi})$ under the empirical sup-norm satisfies
$\mathcal{V}_{\infty}\left(\tfrac{1}{T},\log(\mathcal{F}_{\rm MLP}(L,W,S)),T\right)\leq C\cdot SL\log S\cdot\log T,$
for some universal constant $C>0$.
The proof of Lemma 19 is provided in Appendix H. Building on Lemmas 18 and 19, we establish an upper bound on the expected excess risk for MLP-based classifiers in the ICL setting, stated as follows.
Theorem 20
Assume $\mathbf{p}_{0}\in\mathcal{G}_{\alpha}(\beta,Q)$ for some $\alpha\in[0,1]$. Let $\hat{\mathbf{p}}$ be an estimator of $\mathbf{p}_{0}$, taking values in the function class $\mathcal{F}_{\rm MLP}(L,W,S)$, where the parameters of the class scale with the precision parameter $\varepsilon>0$ as follows:
$L\lesssim\log\left(\frac{1}{\varepsilon}\right),\ W\lesssim K\varepsilon^{-d/\beta},\ S\lesssim K\varepsilon^{-d/\beta}\log\left(\frac{1}{\varepsilon}\right).$
[…]
$\Delta_{T}(\mathbf{p}_{0},\widehat{\mathbf{p}})\lesssim T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}K^{4+\alpha},$
and we choose the precision parameter $\varepsilon$ as
$\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}},$
for some constant $C>0$, then the ICL risk is bounded by
$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}\right)\lesssim B^{2}T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}K^{4+\alpha}\log^{4}\left(KT\right).$
In particular, for the special case where $\alpha=1$, the ICL risk achieves the bound
$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}\right)\lesssim B^{2}T^{-\frac{2\beta}{2\beta+d}}K^{5}\log^{4}\left(KT\right).$
Remark 21
We first show that the risk bounds in Theorems 20 and 15 both achieve the optimal rate $T^{-\frac{2\beta}{2\beta+d}}$, up to logarithmic factors. Recent works have empirically shown that MLPs can also exhibit ICL behavior (Tong and Pehlevan, 2025; Kratsios and Furuya, 2025; Abedsoltan et al., 2024). Building on this, we prove that in the task-scaling regime, where the number of tasks $T$ increases while the context length remains fixed, the convergence rate of MLPs matches that of transformer-based models. This provides a theoretical justification for the comparable performance of MLP-based ICL under realistic pretraining conditions.
Proof of Theorem 20. The approximation error is controlled as shown in Lemma 18. Using these bounds, we construct an estimator $\hat{\mathbf{p}}\in\mathcal{F}_{MLP}(L,W,S)$ such that $\|\hat{\mathbf{p}}-\mathbf{p}_{0}\|_{\infty}\leq\varepsilon$. Applying Theorems 711, and Lemma 19, and expressing the network parameters as functions of $\varepsilon$, we obtain the claimed in-context risk bound. $\blacksquare$
7 Discussions
The upper bound established in Theorem 15 reflects the curse of dimensionality, as the input prompt dimension $d$ explicitly appears in the convergence rate exponent. This phenomenon is well-known in nonparametric estimation under smoothness constraints, where the optimal minimax rate for estimating a $\beta$-Hölder function scales as $T^{-2\beta/(2\beta+d)}$ (Donoho and Johnstone, 1998). Our result demonstrates that this dependence persists in transformer-based ICL under general smoothness assumptions.
Nevertheless, the practical impact of dimensionality may be mitigated, as the number of tasks $T$ is typically large in realistic ICL scenarios. Furthermore, transformer architectures are believed to possess an inherent ability to adapt to latent low-dimensional structure within the input, potentially reducing their effective dependence on the prompt dimension.
Recent work has explored multiple strategies to overcome the curse of dimensionality. One strategy is to explore alternative function classes with more favorable approximation properties, such as mixed-Besov spaces (Suzuki, 2019) and Korobov spaces (Montanelli and Du, 2019). Another approach is to impose structural assumptions on the target function, including additive ridge functions (Fang and Cheng, 2023), hierarchical compositions (Schmidt-Hieber, 2020; Han et al., 2020), generalized single-index models (Bauer and Kohler, 2019), and multivariate adaptive regression splines (MARS) (Kohler et al., 2022).
Beyond structural assumptions on function classes, recent literature highlights the importance of leveraging the geometric structure of input data. The manifold hypothesis (Pope et al., 2021) suggests that although data appear in high-dimensional spaces, their intrinsic structure is governed by a low-dimensional manifold (Tenenbaum et al., 2000; Roweis and Saul, 2000). Under this assumption, deep neural networks have been shown to achieve significantly faster convergence rates by effectively adapting to the intrinsic dimensionality of data (Nakada and Imaizumi, 2020; Chen et al., 2022, 2019; Schmidt-Hieber, 2019; Cloninger and Klock, 2021). Recent studies further characterize the statistical properties of deep learning and generative models under low intrinsic dimensionality, providing refined generalization guarantees in these settings (Chakraborty and Bartlett, 2024a, b).
Despite these advances, the theoretical understanding of transformers in low-dimensional geometric settings remains limited. A recent contribution by Havrilla and Liao (2024) analyzes transformer approximation for a regression problem on low-dimensional manifolds, whose setting is quite different from ours. Extending Lemma 8 to manifold-supported functions deserve further study, as this would allow risk bounds to scale with the intrinsic dimension rather than the ambient dimension.
Appendix A Proof of Theorem 7
Proof  We begin with some notations that will be used in the proof.
  • The expected KL risk defined as:
    $R_{B}(\mathbf{p}_{0},\hat{\mathbf{p}})=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[\boldsymbol{p}_{0}(\mathbf{Z})^{\top}\left(B\wedge\log\left(\frac{\boldsymbol{p}_{0}(\mathbf{Z})}{\hat{\boldsymbol{p}}(\mathbf{Z})}\right)\right)\right].$
  • The expected empirical risk is given by:
    $R_{B,T}(\mathbf{p}_{0},\hat{\mathbf{p}}):=\mathbb{E}_{\mathcal{D}_{T}}\left[\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}^{(t)\top}\left(B\wedge\log\left(\frac{\mathbf{p}_{0}(\mathbf{Z}^{(t)})}{\hat{\mathbf{p}}(\mathbf{Z}^{(t)})}\right)\right)\right].$
  • The dataset is denoted by:
    $\mathcal{D}_{T}=\{(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\}_{t=1}^{T}.$
    Here, we omit the subscript $N+1$ from $\mathbf{y}_{N+1}^{(t)}$ for simplicity.
  • The empirical risk gap is defined as:
    $\Delta_{T}(\boldsymbol{p}_{0},\hat{\boldsymbol{p}}):=\mathbb{E}_{\mathcal{D}_{T}}\left[-\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}^{(t)\top}\log(\hat{\boldsymbol{p}}(\mathbf{Z}^{(t)}))-\min_{\boldsymbol{p}\in\mathcal{F}}\left(-\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}^{(t)\top}\log(\boldsymbol{p}(\mathbf{Z}^{(t)}))\right)\right].$
[…]
$\quad+\mathbb{E}_{\mathcal{D}_{T}}\mathbb{E}_{\mathbf{Z}}\sum_{k=1}^{K}p^{0}_{k}(\mathbf{Z})\left[B\wedge\log\left(\frac{p^{0}_{k}(\mathbf{Z})}{\hat{p}_{k}(\mathbf{Z})}\right)\right]\mathbb{I}\left(p^{0}_{k}(\mathbf{Z})<\frac{C_{T}}{T}\right)$
[…]
$\quad+\mathbb{E}_{\mathcal{D}_{T}}\frac{1}{T}\sum_{t=1}^{T}\sum_{k=1}^{K}y_{k}^{(t)}\left[B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z}^{(t)})}{\hat{p}_{k}(\mathbf{Z}^{(t)})}\right)\right]\mathbb{I}\left(p_{k}^{0}(\mathbf{Z}^{(t)})<\frac{C_{T}}{T}\right)$
[…]
where
[…]
For $W_{1}$, we introduce the truncated loss function
$g_{\mathbf{p}}(\mathbf{Z},\mathbf{y}):=\mathbf{y}^{\top}\left[\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\mathbf{p}(\mathbf{Z})}\right)_{\geq\frac{C_{T}}{T}}\right]=\sum_{k=1}^{K}y_{k}\left[B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z})}{p_{k}(\mathbf{Z})}\right)\right]\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})\geq\frac{C_{T}}{T}\right).$
With this notation, the quantity $W_{1}$ can be expressed as
[…]
$=\ \mathbb{E}_{\mathcal{D}_{T}}\,\mathbb{E}_{(\mathbf{Z},\mathbf{y})}\left[g_{\hat{\mathbf{p}}}(\mathbf{Z},\mathbf{y})\right]-2\,\mathbb{E}_{\mathcal{D}_{T}}\left[\frac{1}{T}\sum_{i=1}^{T}g_{\hat{\mathbf{p}}}(\mathbf{Z}^{(i)},\mathbf{y}^{(i)})\right]$
$\leq\ \int_{0}^{\infty}\mathbb{P}\left(\mathbb{E}_{(\mathbf{Z},\mathbf{y})}\,g_{\hat{\mathbf{p}}}(\mathbf{Z},\mathbf{y})-\frac{2}{T}\sum_{i=1}^{T}g_{\hat{\mathbf{p}}}(\mathbf{Z}^{(i)},\mathbf{y}^{(i)})>t\right)dt.$ (11)
Let $C_{T}$ be small enough such that $C_{T}\leq C_{1}T^{1-C_{2}}$ where $C_{1}>0$ and $0<C_{2}<1$. Then the application of Lemma 24 with $\epsilon=1/2$ and $a=b=t/2$ yields
$\ \mathbb{P}\left(\mathbb{E}_{(\mathbf{Z},\mathbf{y})}g_{\hat{\mathbf{p}}}(\mathbf{Z},\mathbf{y})-\frac{2}{T}\sum_{i=1}^{T}g_{\hat{\mathbf{p}}}(\mathbf{Z}^{(i)},\mathbf{y}^{(i)})>t\right)$
$\leq$ $\ \mathbb{P}\left(\exists\ \mathbf{p}\in\mathcal{F}:\mathbb{E}_{(\mathbf{Z},\mathbf{y})}g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})-\frac{2}{T}\sum_{i=1}^{T}g_{\mathbf{p}}(\mathbf{Z}^{(i)},\mathbf{y}^{(i)})>t\right)$
$=$ $\ \mathbb{P}\left(\exists\ \mathbf{p}\in\mathcal{F}:\mathbb{E}_{(\mathbf{Z},\mathbf{y})}g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})-\frac{1}{T}\sum_{i=1}^{T}g_{\mathbf{p}}(\mathbf{Z}^{(i)},\mathbf{y}^{(i)})>\frac{1}{2}\{t+\mathbb{E}_{(\mathbf{Z},\mathbf{y})}g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})\}\right)$
$\leq$ $\ 10\sup_{\mathcal{Z}_{T}}\mathcal{N}\left(\frac{t}{32},\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\exp\left(-\frac{tT}{480B_{T}^{2}}\right),$
where $B_{T}=B\vee\left|\log\left(C_{T}/T\right)\right|$ and $\|f-g\|_{\infty,\mathcal{Z}_{T}}=\max_{1\leq t\leq T}\max_{1\leq k\leq K}\left|f_{k}(\mathbf{Z}^{(t)})-g_{k}(\mathbf{Z}^{(t)})\right|$. Consequently,
$W_{1}\leq$ $\ \int_{0}^{\infty}\mathbb{P}\left(\mathbb{E}_{(\mathbf{Z},\mathbf{y})}g_{\hat{\mathbf{p}}}(\mathbf{Z},\mathbf{y})-\frac{2}{T}\sum_{i=1}^{T}g_{\hat{\mathbf{p}}}(\mathbf{Z}^{(i)},\mathbf{y}^{(i)})>t\right)dt$
$\leq$ $\ \alpha_{n}+\int_{\alpha_{n}}^{\infty}\mathbb{P}\left(\mathbb{E}_{(\mathbf{Z},\mathbf{y})}g_{\hat{\mathbf{p}}}(\mathbf{Z},\mathbf{y})-\frac{2}{T}\sum_{i=1}^{T}g_{\hat{\mathbf{p}}}(\mathbf{Z}^{(i)},\mathbf{y}^{(i)})>t\right)dt$
$\leq$ $\ \alpha_{n}+\int_{\alpha_{n}}^{\infty}10\sup_{\mathcal{Z}_{T}}\mathcal{N}\left(\frac{t}{32},\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\exp\left(-\frac{tT}{480B_{T}^{2}}\right)dt$
$\leq$ $\ \alpha_{n}+10\sup_{\mathcal{Z}_{T}}\mathcal{N}\left(\frac{\alpha_{n}}{32},\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\int_{\alpha_{n}}^{\infty}\exp\left(-\frac{tT}{480B_{T}^{2}}\right)dt$
$=$ $\ \alpha_{n}+10\sup_{\mathcal{Z}_{T}}\mathcal{N}\left(\frac{\alpha_{n}}{32},\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\frac{480B_{T}^{2}}{T}\exp\left(-\frac{\alpha_{n}T}{480B_{T}^{2}}\right).$
Now let $\alpha_{n}=(\log(10\mathcal{N}_{T}))480B_{T}^{2}/T$ with
$\mathcal{N}_{T}=\mathcal{N}_{\infty}\left(\frac{1}{T},\log(\mathcal{F}),T\right)=\sup_{\mathcal{Z}_{T}}\mathcal{N}\left(\frac{1}{T},\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right).$
It is easy to see that $1/T\leq\alpha_{n}/32$ and then
$\sup_{\mathcal{Z}_{T}}\mathcal{N}\left(\frac{\alpha_{n}}{32},\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\sup_{\mathcal{Z}_{T}}\mathcal{N}\left(\frac{1}{T},\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)=\mathcal{N}_{T}.$
Consequently,
$W_{1}\leq\log(10\mathcal{N}_{T})\frac{480B_{T}^{2}}{T}+\frac{480B_{T}^{2}}{T}=\{\log(10\mathcal{N}_{T})+1\}\frac{480B_{T}^{2}}{T}.$ (12)
[…]
$\leq$ $\ \frac{KC_{T}\{\log(T/C_{T})+B\}}{T},$ (13)
and
[…]
$\leq$ $\ \frac{KC_{T}\{\log(T/C_{T})+B\}}{T}.$ (14)
Combining (10), (12), (A), and (A), we have
[…]
$\leq$ $\ 2R_{B,T}(\mathbf{p}_{0},\widehat{\mathbf{p}})+\{\log(10\mathcal{N}_{T})+1\}\frac{480B_{T}^{2}}{T}+\frac{3KC_{T}\{\log(T/C_{T})+B\}}{T}.$ (15)
Recall that $B_{T}=B\vee\left|\log\left(\frac{C_{T}}{T}\right)\right|$, then the application of Lemma 23 yields the oracle inequality
[…]
$\quad+\frac{480B_{T}^{2}\{\log(10\mathcal{N}_{T})+1\}}{T}+\frac{3KC_{T}\{\log(T/C_{T})+B\}}{T}$
$\leq\ C\left(\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})+\Delta_{T}(\mathbf{p}_{0},\widehat{\mathbf{p}})+\frac{B^{2}\log(\mathcal{N}_{T})\log^{2}(T)}{T}\right).$
where $C>0$ is an universal constant. This completes the proof of Theorem 7.
A.1 Supporting Lemmas
Lemma 22 (Bound for Small Probabilities)
Let $\mathcal{F}$ be a class of conditional class probabilities, and let $\hat{\mathbf{p}}$ be an estimator of $\mathbf{p}_{0}$ taking values in $\mathcal{F}$. Suppose we have a sample $\mathcal{D}_{T}=\{(\mathbf{Z}^{(1)},\mathbf{y}^{(1)}),\cdots,(\mathbf{Z}^{(T)},\mathbf{y}^{(T)})\}$ with the same distribution as $(\mathbf{Z},\mathbf{y})$ and $(\mathbf{Z},\mathbf{y})$ is independent of $\mathcal{D}_{T}$. Given a constant $C_{T}$ satisfying $0<C_{T}\leq T/e$, the following bound holds:
$\bigg{|}\mathbb{E}_{\mathcal{D}_{T},(\mathbf{Z},\mathbf{y})}y_{k}\bigg{[}B\wedge\log\frac{p_{k}^{0}(\mathbf{Z})}{\widehat{p}_{k}(\mathbf{Z})}\bigg{]}\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})<\frac{C_{T}}{T}\right)\bigg{|}\leq\frac{C_{T}\{\log(T/C_{T})+B\}}{T}$
$\bigg{|}\mathbb{E}_{\mathcal{D}_{T}}\left[y_{k}^{(t)}\left(B\wedge\log\frac{p_{k}^{0}(\mathbf{Z}^{(t)})}{\hat{p}_{k}(\mathbf{Z}^{(t)})}\right)\mathbb{I}\left(p_{k}^{0}(\mathbf{Z}^{(t)})<\frac{C_{T}}{T}\right)\right]\bigg{|}\leq\frac{C_{T}\{\log(T/C_{T})+B\}}{T},$
for any $t\in\{1\dots T\}$ and $k\in\{1,\dots,K\}$.
Proof  Because $p^{0}_{k},\widehat{p}_{k}\in[0,1]$, we first observe that $\log(p_{k}^{0}(\mathbf{Z}))\leq B\wedge\log\frac{p_{k}^{0}(\mathbf{Z})}{\widehat{p}_{k}(\mathbf{Z})}\leq B$. Thus, taking absolute values, we have
$\bigg{|}\mathbb{E}_{\mathcal{D}_{T},(\mathbf{Z},\mathbf{y})}y_{k}\bigg{[}B\wedge\log\frac{p_{k}^{0}(\mathbf{Z})}{\widehat{p}_{k}(\mathbf{Z})}\bigg{]}\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})<\frac{C_{T}}{T}\right)\bigg{|}$
$\leq$ $\mathbb{E}_{\mathcal{D}_{T},(\mathbf{Z},\mathbf{y})}y_{k}\{|\log(p_{k}^{0}(\mathbf{Z}))|+B\}\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})<\frac{C_{T}}{T}\right)$
$\leq$ $\mathbb{E}_{\mathbf{Z}}\left[p_{k}^{0}(\mathbf{Z})\{|\log(p_{k}^{0}(\mathbf{Z}))|+B\}\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})<\frac{C_{T}}{T}\right)\right]$
$=$ $\mathbb{E}_{\mathbf{Z}}\left[p_{k}^{0}(\mathbf{Z})|\log(p_{k}^{0}(\mathbf{Z}))|\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})<\frac{C_{T}}{T}\right)\right]+B\mathbb{E}_{\mathbf{Z}}\left[p_{k}^{0}(\mathbf{Z})\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})\leq\frac{C_{T}}{T}\right)\right]$
$\leq$ $\mathbb{E}_{\mathbf{Z}}\left[p_{k}^{0}(\mathbf{Z})|\log(p_{k}^{0}(\mathbf{Z}))|\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})<\frac{C_{T}}{T}\right)\right]+B\frac{C_{T}}{T}.$
Since the function $u\mapsto u|\log(u)|$ is increasing on $(0,e^{-1})$ and $C_{T}/T\leq e^{-1}$, it follows that
$p_{k}^{0}(\mathbf{Z})|\log(p_{k}^{0}(\mathbf{Z}))|\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})<\frac{C_{T}}{T}\right)\leq\frac{C_{T}}{T}\log\frac{T}{C_{T}}.$
Consequently,
$\bigg{|}\mathbb{E}_{\mathcal{D}_{T},(\mathbf{Z},\mathbf{y})}y_{k}\bigg{[}B\wedge\log\frac{p_{k}^{0}(\mathbf{Z})}{\widehat{p}_{k}(\mathbf{Z})}\bigg{]}\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})<\frac{C_{T}}{T}\right)\bigg{|}\leq\frac{C_{T}\{\log(T/C_{T})+B\}}{T}.$
Similarly,
$\bigg{|}\mathbb{E}_{\mathcal{D}_{T}}\left[y_{k}^{(t)}\left(B\wedge\log\frac{p_{k}^{0}(\mathbf{Z}^{(t)})}{\hat{p}_{k}(\mathbf{Z}^{(t)})}\right)\mathbb{I}\left(p_{k}^{0}(\mathbf{Z}^{(t)})<\frac{C_{T}}{T}\right)\right]\bigg{|}$
$\leq$ $\mathbb{E}_{\mathcal{D}_{T}}\left[y_{k}^{(t)}\{|\log(p_{k}^{0}(\mathbf{Z}^{(t)}))|+B\}\mathbb{I}\left(p_{k}^{0}(\mathbf{Z}^{(t)})<\frac{C_{T}}{T}\right)\right]$
$=$ $\mathbb{E}_{\mathbf{Z}^{(t)}}\left[p_{k}^{0}(\mathbf{Z}^{(t)})\{|\log(p_{k}^{0}(\mathbf{Z}^{(t)}))|+B\}\mathbb{I}\left(p_{k}^{0}(\mathbf{Z}^{(t)})<\frac{C_{T}}{T}\right)\right]$
$\leq$ $\frac{C_{T}\{\log(T/C_{T})+B\}}{T}.$
This complete the proof of this lemma.
Lemma 23
For any estimator $\widehat{\mathbf{p}}\in\mathcal{F}$, the following inequality holds:
[…]
Proof  The first inequality follows directly from the definition, since for any $a,b\in\mathbb{R}$, we have $a\geq\min(a,b)$.
To show the second inequality, fix any arbitrary $\mathbf{p}^{*}\in\mathcal{F}$. Since $\Delta_{T}(\mathbf{p}_{0},\mathbf{p}^{*})\geq 0$, it follows that
[…]
where the equality follows from the law of total expectation and the fact that $\mathbb{E}[\mathbf{y}^{(t)}|\mathbf{Z}^{(t)}]=\mathbf{p}_{0}(\mathbf{Z}^{(t)})$. Note that
$\mathbb{E}_{\mathcal{D}_{T}}\left[\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}^{(t)\top}\log(\mathbf{p}_{0}(\mathbf{Z}^{(t)}))\right]=\mathbb{E}_{\mathbf{Z}}\left[\mathbf{p}_{0}(\mathbf{Z})^{\top}\log(\mathbf{p}_{0}(\mathbf{Z}))\right].$
Adding $\mathbb{E}_{\mathbf{Z}}[\mathbf{p}_{0}(\mathbf{Z})^{\top}\log(\mathbf{p}_{0}(\mathbf{Z}))]$ to both sides yields the desired inequality.
Recall that $\mathbf{p}(\mathbf{Z})=(p_{1}(\mathbf{Z}),\dots,p_{K}(\mathbf{Z}))$ and $g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})=\mathbf{y}^{\top}\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\mathbf{p}(\mathbf{Z})}\right)_{\geq C_{T}/T}$, where $B\geq 2$. We then have the following result, which is used to derive the oracle inequality.
Lemma 24
Suppose that for $\mathbf{p}\in\mathcal{F}$, $p_{k}(\mathbf{Z})\geq C_{1}/T^{C_{2}}$ for some constants $C_{1}>0$ and $0<C_{2}<1$, and $C_{T}\leq C_{1}T^{1-C_{2}}$. Then for each $\epsilon\in(0,\tfrac{1}{2}]$ and $a,b>0$, we have
$\mathbb{P}\Big{\{}\exists\ \mathbf{p}\in\mathcal{F}:\ \mathbb{E}g_{\bf p}(\mathbf{Z,y})-\frac{1}{T}\sum_{t=1}^{T}g_{\bf p}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\geq\epsilon\{a+b+\mathbb{E}g_{\bf p}(\mathbf{Z,y})\}\Big{\}}$
$\leq$ $10\sup_{\mathcal{Z}_{T}}\mathcal{N}\left(\frac{\epsilon b}{8},\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\exp\left(-\frac{\epsilon^{2}(1-\epsilon)aT}{15(1+\epsilon)B_{T}^{2}}\right).$
Proof  The proof of this Lemma follows the same line as Theorem 11.4 of Györfi et al. (2006) with some extra complications arising from dealing with the truncated loss function $g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})$ . We here provide a detailed proof for the sake of completeness. Recall that
$g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})=\mathbf{y}^{\top}\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\mathbf{p}(\mathbf{Z})}\right)_{\geq\frac{C_{T}}{T}}=\sum_{k=1}^{K}y_{k}\left[B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z})}{p_{k}(\mathbf{Z})}\right)\right]\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})\geq\frac{C_{T}}{T}\right).$
It is easy to see that
$g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})$ $\leq\ \sum_{k=1}^{K}y_{k}B=B,$
$g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})$ $\geq\ \sum_{k=1}^{K}y_{k}\left[B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z})}{p_{k}(\mathbf{Z})}\right)\right]\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})\geq\frac{C_{T}}{T}\right)$
$\geq\ \sum_{k=1}^{K}y_{k}\left[B\wedge\log\left(\frac{C_{T}}{T}\right)\right]=\log\left(\frac{C_{T}}{T}\right).$
Write $B_{T}=B\vee\left|\log(C_{T}/T)\right|$, it follows that $|g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})|\leq B_{T}$. The rest of the proof will be divided into six steps to enhance the readability.
Step 1. Symmetrization.
Let $\tilde{\mathcal{D}}_{T}=\{(\mathbf{\tilde{Z}}^{(t)},\mathbf{\tilde{y}}^{(t)})\}_{t=1}^{T}$ be i.i.d. random variables with same distribution as $(\mathbf{Z},\mathbf{y})$ and independent of $\mathcal{D}_{T}$. Consider the event
$\bigg{\{}\exists\ \mathbf{p}\in\mathcal{F}:\ \mathbb{E}g_{\bf p}(\mathbf{Z,y})-\frac{1}{T}\sum_{t=1}^{T}g_{\bf p}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\geq\epsilon[a+b+\mathbb{E}g_{\bf p}(\mathbf{Z,y})]\bigg{\}}.$
Let $\mathbf{p}^{T}\in\mathcal{F}$ be a function depending on $\mathcal{D}_{T}=\{(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\}_{t=1}^{T}$ such that
$\bigg{\{}\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]-\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}^{T}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\geq\epsilon\{a+b+\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]\}\bigg{\}}.$
Note that
$\mathbb{E}[g_{\mathbf{p}}^{2}(\mathbf{Z},\mathbf{y})]$ $=$ $\mathbb{E}\left(\sum_{k=1}^{K}y_{k}\left[B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z})}{p_{k}(\mathbf{Z})}\right)\right]\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})\geq\frac{C_{T}}{T}\right)\right)^{2}$
$=$ $\mathbb{E}\left(\sum_{k=1}^{K}y_{k}\left[B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z})}{p_{k}(\mathbf{Z})}\right)\right]^{2}\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})\geq\frac{C_{T}}{T}\right)\right)$
$=$ $\mathbb{E}\left(\sum_{k=1}^{K}p_{k}^{0}(\mathbf{Z})\left[B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z})}{p_{k}(\mathbf{Z})}\right)\right]^{2}\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})\geq\frac{C_{T}}{T}\right)\right)$
$\leq$ $\mathbb{E}\left(\sum_{k=1}^{K}p_{k}^{0}(\mathbf{Z})\left[B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z})}{p_{k}(\mathbf{Z})}\right)\right]^{2}\right)$
$\leq$ $B\mathbb{E}\left(\sum_{k=1}^{K}p_{k}^{0}(\mathbf{Z})\left[B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z})}{p_{k}(\mathbf{Z})}\right)\right]\right),$
where the second equation is due to the fact that $\mathbf{y}$ is one-hot encoded, and the second inequality follows from Lemma 25 in the following. Since for any $\mathbf{p}\in\mathcal{F}$, we have $p_{k}(\mathbf{Z})\geq C_{1}/T^{C_{2}}$ for some constants $C_{1}>0$ and $0<C_{2}<1$, and $C_{T}\leq C_{1}T^{1-C_{2}}$, it follows that
$\mathbb{E}\left(\sum_{k=1}^{K}p_{k}^{0}(\mathbf{Z})\left[B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z})}{p_{k}(\mathbf{Z})}\right)\right]\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})\leq\frac{C_{T}}{T}\right)\right)$
$\leq$ $\mathbb{E}\left(\sum_{k=1}^{K}p_{k}^{0}(\mathbf{Z})\left[B\wedge\log\left(\frac{C_{T}/T}{C_{1}/T^{C_{2}}}\right)\right]\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})\leq\frac{C_{T}}{T}\right)\right)$
$=$ $\mathbb{E}\left(\sum_{k=1}^{K}p_{k}^{0}(\mathbf{Z})\left[B\wedge\log\left(\frac{C_{T}}{C_{1}T^{1-C_{2}}}\right)\right]\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})\leq\frac{C_{T}}{T}\right)\right)\leq 0,$
where the last inequality follows from $\log(C_{T}/C_{1}T^{1-C_{2}})\leq B$. Consequently,
$\mathbb{E}[g_{\mathbf{p}}^{2}(\mathbf{Z,y})]$ $\leq$ $B\mathbb{E}\left(\sum_{k=1}^{K}p_{k}^{0}(\mathbf{Z})\left[B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z})}{p_{k}(\mathbf{Z})}\right)\right]\right)$
$\leq$ $B\mathbb{E}\left(\sum_{k=1}^{K}p_{k}^{0}(\mathbf{Z})\left[B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z})}{p_{k}(\mathbf{Z})}\right)\right]\mathbb{I}(p_{k}^{0}(\mathbf{Z})\geq C_{T}/T)\right)$
$\leq$ $B_{T}\mathbb{E}[g_{\mathbf{p}}(\mathbf{Z,y})].$
Combining this and Chebyshev’s inequality, we have
$\mathbb{P}\left(\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]-\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}^{T}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})\geq\frac{\epsilon}{2}\{a+b+\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]\}\Big{|}\mathcal{D}_{T}\right)$
$\leq$ $\frac{Var\{g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}\}}{\frac{T\epsilon^{2}}{4}\{a+b+\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]\}^{2}}$
$\leq$ $\frac{4\mathbb{E}[g^{2}_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}\}}{T\epsilon^{2}\{a+b+\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]\}^{2}}$
$\leq$ $\frac{4B_{T}\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}\}}{T\epsilon^{2}\{a+b+\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]\}^{2}}$
$\leq$ $\frac{4B_{T}}{T\epsilon^{2}}\frac{1}{4(a+b)}=\frac{B_{T}}{T(a+b)\epsilon^{2}},$
where the last inequality follows from the fact that $\frac{t}{(a+t)^{2}}\leq\frac{1}{4a}$ for all $t\geq 0$ and $a>0$. Therefore, for $T\geq\frac{8B_{T}}{(a+b)\epsilon^{2}}$,
$\mathbb{P}\left(\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]-\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}^{T}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})\leq\frac{\epsilon}{2}\{a+b+\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]\}\Big{|}\mathcal{D}_{T}\right)\geq\frac{7}{8}.$ (16)
Consequently,
$\ \mathbb{P}\left(\exists\ \mathbf{p}\in\mathcal{F}:\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})-\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\geq\frac{\epsilon}{2}\{a+b+\mathbb{E}g_{\mathbf{p}}(\mathbf{Z,y})\}\right)$
$\geq$ $\ \mathbb{P}\left(\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}^{T}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})-\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}^{T}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\geq\frac{\epsilon}{2}\{a+b+\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]\}\right)$
$\geq$ $\ \mathbb{P}\Big{(}\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]-\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}^{T}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\geq\epsilon\{a+b+\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]\},$
$\ \mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]-\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}^{T}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})\leq\frac{\epsilon}{2}\{a+b+\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]\}\Big{)}$
$=$ $\ \mathbb{E}\Big{\{}\mathbb{I}\Big{(}\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]-\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}^{T}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\geq\epsilon\{a+b+\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]\}\Big{)}$
$\ \times\mathbb{I}\Big{(}\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}]-\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}^{T}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})\leq\frac{\epsilon}{2}\{a+b+\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]\}\Big{)}\Big{\}}$
$=$ $\ \mathbb{E}\Bigg{\{}\mathbb{I}\Big{(}\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]-\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}^{T}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\geq\epsilon\{a+b+\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]\}\Big{)}$
$\ \times\mathbb{E}\Big{\{}\mathbb{I}\Big{(}\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]-\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}^{T}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})\leq\frac{\epsilon}{2}\{a+b+\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]\}\Big{)}\Big{|}\mathcal{D}_{T}\Big{\}}\Bigg{\}}$
$\geq$ $\ \frac{7}{8}\mathbb{P}\Big{\{}\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]-\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}^{T}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\geq\epsilon\{a+b+\mathbb{E}[g_{\mathbf{p}^{T}}(\mathbf{Z,y})|\mathcal{D}_{T}]\}\Big{\}}$
$=$ $\ \frac{7}{8}\mathbb{P}\Big{\{}\exists\ \mathbf{p}\in\mathcal{F}:\ \mathbb{E}g_{\bf p}(\mathbf{Z,y})-\frac{1}{T}\sum_{t=1}^{T}g_{\bf p}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\geq\epsilon\{a+b+\mathbb{E}g_{\bf p}(\mathbf{Z,y})\}\Big{\}},$
where the last inequality is due to (16). Thus, we obtain the following symmetrization inequality
$\ \mathbb{P}\Big{\{}\exists\ \mathbf{p}\in\mathcal{F}:\ \mathbb{E}g_{\bf p}(\mathbf{Z,y})-\frac{1}{T}\sum_{t=1}^{T}g_{\bf p}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\geq\epsilon\{a+b+\mathbb{E}g_{\bf p}(\mathbf{Z,y})\}\Big{\}}$
$\leq$ $\ \frac{8}{7}\mathbb{P}\left(\exists\ \mathbf{p}\in\mathcal{F}:\frac{1}{T}\sum_{t=1}^{T}\{g_{\mathbf{p}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})-g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\}\geq\frac{\epsilon}{2}\{a+b+\mathbb{E}g_{\mathbf{p}}(\mathbf{Z,y})\}\right)$ (17)
for all $T\geq\frac{8B_{T}}{(a+b)\epsilon^{2}}$.
Step 2. Replacement of the expectation $\mathbb{E}g_{\mathbf{p}}(\mathbf{Z,y})$ in (A.1) by a sample mean of $\tilde{\mathcal{D}}_{T}$. By (A.1), it follows that
$\mathbb{P}\left(\exists\ \mathbf{p}\in\mathcal{F}:\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})-\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\geq\frac{\epsilon}{2}\{a+b+\mathbb{E}g_{\mathbf{p}}(\mathbf{Z,y})\}\right)$ (18)
$\leq$ $\mathbb{P}\Bigg{(}\exists\ \mathbf{p}\in\mathcal{F}:\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})-\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\geq\frac{\epsilon}{2}\{a+b+\mathbb{E}g_{\mathbf{p}}(\mathbf{Z,y})\},$
$\frac{1}{T}\sum_{t=1}^{T}g^{2}_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-\mathbb{E}g^{2}_{\mathbf{p}}(\mathbf{Z,y})\leq\epsilon\{a+b+\frac{1}{T}\sum_{t=1}^{T}g^{2}_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})+\mathbb{E}g^{2}_{\mathbf{p}}(\mathbf{Z,y})\},$
$\frac{1}{T}\sum_{t=1}^{T}g^{2}_{\mathbf{p}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})-\mathbb{E}g^{2}_{\mathbf{p}}(\mathbf{Z,y})\leq\epsilon\{a+b+\frac{1}{T}\sum_{t=1}^{T}g^{2}_{\mathbf{p}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})+\mathbb{E}g^{2}_{\mathbf{p}}(\mathbf{Z,y})\}\Bigg{)}$
$+2\mathbb{P}\Bigg{(}\exists\ \mathbf{p}\in\mathcal{F}:\frac{\frac{1}{T}\sum_{t=1}^{T}g^{2}_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-\mathbb{E}g^{2}_{\mathbf{p}}(\mathbf{Z,y})}{a+b+\frac{1}{T}\sum_{t=1}^{T}g^{2}_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})+\mathbb{E}g^{2}_{\mathbf{p}}(\mathbf{Z,y})}>\epsilon\Bigg{)}$
$=:$ $I_{1}+2I_{2}.$
For $I_{2}$, it follows from Theorem 11.6 of Györfi et al. (2006) that
$I_{2}$ $=$ $\mathbb{P}\Bigg{(}\exists\ \mathbf{p}\in\mathcal{F}:\frac{\frac{1}{T}\sum_{t=1}^{T}g^{2}_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-\mathbb{E}g^{2}_{\mathbf{p}}(\mathbf{Z,y})}{a+b+\frac{1}{T}\sum_{t=1}^{T}g^{2}_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})+\mathbb{E}g^{2}_{\mathbf{p}}(\mathbf{Z,y})}>\epsilon\Bigg{)}$ (19)
$\leq$ $4\mathbb{E}\Big{(}\mathcal{N}(\frac{(a+b)\epsilon}{5},\mathcal{F}_{1},\|\cdot\|_{L_{1},\mathcal{D}_{T}})\exp(-\frac{3\epsilon^{2}(a+b)T}{40B_{T}^{2}})\Big{)},$
where $\mathcal{F}_{1}=\{g_{\mathbf{p}}(\mathbf{Z},\mathbf{y}):\mathbf{p}\in\mathcal{F}\}$ and $\|f-g\|_{L_{1},\mathcal{D}_{T}}=\frac{1}{T}\sum_{t=1}^{T}|f(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-g(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})|$.
Next we discuss the first probability $I_{1}$ in the decomposition (18). Note that the second and third inequalities in $I_{1}$ imply that
$(1-\epsilon)\frac{1}{T}\sum_{t=1}^{T}g^{2}_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-\epsilon(a+b)\leq(1+\epsilon)\mathbb{E}g^{2}_{\mathbf{p}}(\mathbf{Z,y}),$
$(1-\epsilon)\frac{1}{T}\sum_{t=1}^{T}g^{2}_{\mathbf{p}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})-\epsilon(a+b)\leq(1+\epsilon)\mathbb{E}g^{2}_{\mathbf{p}}(\mathbf{Z,y}).$
Combining these with the fact $\mathbb{E}[g_{\mathbf{p}}^{2}(\mathbf{Z,y})]\leq B_{T}\mathbb{E}[g_{\mathbf{p}}(\mathbf{Z,y})]$, we have
$I_{1}=$ $\ \mathbb{P}\Bigg{(}\exists\ \mathbf{p}\in\mathcal{F}:\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})-\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\geq\frac{\epsilon}{2}\left\{a+b+\mathbb{E}g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})\right\},$
$\quad\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}}^{2}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-\mathbb{E}g_{\mathbf{p}}^{2}(\mathbf{Z},\mathbf{y})\leq\epsilon\left\{a+b+\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}}^{2}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})+\mathbb{E}g_{\mathbf{p}}^{2}(\mathbf{Z},\mathbf{y})\right\},$
$\quad\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}}^{2}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})-\mathbb{E}g_{\mathbf{p}}^{2}(\mathbf{Z},\mathbf{y})\leq\epsilon\left\{a+b+\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}}^{2}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})+\mathbb{E}g_{\mathbf{p}}^{2}(\mathbf{Z},\mathbf{y})\right\}\Bigg{)}$
$\leq$ $\ \mathbb{P}\Bigg{\{}\exists\ \mathbf{p}\in\mathcal{F}:\frac{1}{T}\sum_{t=1}^{T}\left[g_{\mathbf{p}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})-g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\right]\geq\frac{\epsilon}{2}(a+b)$
$\quad+\frac{\epsilon}{2}\Bigg{\{}\frac{1-\epsilon}{2(1+\epsilon)B_{T}}\left(\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}}^{2}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})+\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}}^{2}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})\right)-\frac{2\epsilon(a+b)}{2(1+\epsilon)B_{T}}\Bigg{\}}\Bigg{\}}$
$=$ $\ \mathbb{P}\Bigg{\{}\exists\ \mathbf{p}\in\mathcal{F}:\frac{1}{T}\sum_{t=1}^{T}\left[g_{\mathbf{p}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})-g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\right]\geq\frac{\epsilon}{2}(a+b)$
$\quad+\frac{\epsilon(1-\epsilon)}{4(1+\epsilon)B_{T}}\cdot\frac{1}{T}\sum_{t=1}^{T}\left[g_{\mathbf{p}}^{2}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})+g_{\mathbf{p}}^{2}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})\right]-\frac{\epsilon^{2}(a+b)}{2(1+\epsilon)B_{T}}\Bigg{\}}.$ (20)
Step 3. Second symmetrization.
Now we discuss the symmetrization for the last probability in (A.1). Let $\{U_{t}:t=1,\cdots,T\}$ be an i.i.d sample following the uniform distribution on $\{-1,1\}$. It is easy to see that $g_{\mathbf{p}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})-g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})$ has the same distribution with $U_{t}\{g_{\mathbf{p}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})-g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\}$. Therefore,
$\mathbb{P}\Bigg{\{}\exists\ \mathbf{p}\in\mathcal{F}:\frac{1}{T}\sum_{t=1}^{T}\{g_{\mathbf{p}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})-g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\}\geq\frac{\epsilon}{2}(a+b)$ (21)
$+\frac{\epsilon(1-\epsilon)}{4(1+\epsilon)B_{T}}\frac{1}{T}\sum_{t=1}^{T}\{g^{2}_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})+g^{2}_{\mathbf{p}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})\}-\frac{\epsilon^{2}(a+b)}{2(1+\epsilon)B_{T}}\Bigg{\}}$
$=$ $\mathbb{P}\Bigg{\{}\exists\ \mathbf{p}\in\mathcal{F}:\frac{1}{T}\sum_{t=1}^{T}U_{t}\{g_{\mathbf{p}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})-g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\}\geq\frac{\epsilon}{2}(a+b)$
$+\frac{\epsilon(1-\epsilon)}{4(1+\epsilon)B_{T}}\frac{1}{T}\sum_{t=1}^{T}\{g^{2}_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})+g^{2}_{\mathbf{p}}(\tilde{\mathbf{Z}}^{(t)},\tilde{\mathbf{y}}^{(t)})\}-\frac{\epsilon^{2}(a+b)}{2(1+\epsilon)B_{T}}\Bigg{\}}$
$\leq$ $2\mathbb{P}\Bigg{\{}\exists\ \mathbf{p}\in\mathcal{F}:\Big{|}\frac{1}{T}\sum_{t=1}^{T}U_{t}g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\Big{|}\geq\frac{\epsilon}{4}(a+b)$
$+\frac{\epsilon(1-\epsilon)}{4(1+\epsilon)B_{T}}\frac{1}{T}\sum_{t=1}^{T}g^{2}_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-\frac{\epsilon^{2}(a+b)}{4(1+\epsilon)B_{T}}\Bigg{\}}.$
Now we investigate the last probability in (21) condition on $\mathcal{D}_{T}$. Fixing $\mathcal{D}_{T}=\{(\mathbf{Z}^{(t)},\mathbf{y}^{(t)}):t=1,\dots,T\}$, let $\delta>0$ and let $\mathcal{G}_{\delta}=\{g_{i}:i=1,\dots,M\}$ be a $\delta$-cover of $\mathcal{F}_{1}=\{g_{\mathbf{p}}:\mathbf{p}\in\mathcal{F}\}$ with respect to the norm $\|\cdot\|_{L_{1},\mathcal{D}_{T}}$ on the sample $\mathcal{D}_{T}$, where $M=N(\delta,\mathcal{F}_{1},\|\cdot\|_{L_{1},\mathcal{D}_{T}})$. Without loss of generalization, we assume that $\mathcal{G}_{\delta}\subset\mathcal{F}_{1}$. For a given $\mathbf{p}\in\mathcal{F}$, there exist $g_{i}\in\mathcal{F}_{1}$ such that
$\frac{1}{T}\sum_{t=1}^{T}\big{|}g_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\big{|}\leq\delta.$
This yields
$\Big{|}\frac{1}{T}\sum_{t=1}^{T}U_{t}g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\Big{|}$ $\leq$ $\Big{|}\frac{1}{T}\sum_{t=1}^{T}U_{t}g_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\Big{|}+\frac{1}{T}\sum_{t=1}^{T}\big{|}g_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\big{|}$
$\leq$ $\Big{|}\frac{1}{T}\sum_{t=1}^{T}U_{t}g_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\Big{|}+\delta.$
Note that $|g_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})|\leq B_{T}$ and $|g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})|\leq B_{T}$. Consequently,
$\frac{1}{T}\sum_{t=1}^{T}g^{2}_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})$ $=$ $\frac{1}{T}\sum_{t=1}^{T}g^{2}_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})+\frac{1}{T}\sum_{t=1}^{T}\{g^{2}_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-g^{2}_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\}$
$\geq$ $\frac{1}{T}\sum_{t=1}^{T}g^{2}_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-2B_{T}\frac{1}{T}\sum_{t=1}^{T}|g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-g_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})|$
$\geq$ $\frac{1}{T}\sum_{t=1}^{T}g^{2}_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-2\delta B_{T}.$
It follows that
$\mathbb{P}\Bigg{\{}\exists\ \mathbf{p}\in\mathcal{F}:\left|\frac{1}{T}\sum_{t=1}^{T}U_{t}g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\right|\geq\frac{\epsilon}{4}(a+b)+\frac{\epsilon(1-\epsilon)}{4(1+\epsilon)B_{T}}\cdot\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}}^{2}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})$
$\qquad\qquad\qquad\qquad-\frac{\epsilon^{2}(a+b)}{4(1+\epsilon)B_{T}}\ \Big{|}\ \mathcal{D}_{T}\Bigg{\}}$
$\leq\mathbb{P}\Bigg{\{}\exists\ g_{i}\in\mathcal{G}_{\delta}:\left|\frac{1}{T}\sum_{t=1}^{T}U_{t}g_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\right|\geq-\delta+\frac{\epsilon}{4}(a+b)-\frac{\epsilon^{2}(a+b)}{4(1+\epsilon)B_{T}}$
$\qquad\qquad\qquad\qquad+\frac{\epsilon(1-\epsilon)}{4(1+\epsilon)B_{T}}\cdot\left(\frac{1}{T}\sum_{t=1}^{T}g_{i}^{2}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-2\delta B_{T}\right)\ \Big{|}\ \mathcal{D}_{T}\Bigg{\}}$
$\leq N(\delta,\mathcal{F}_{1},\|\cdot\|_{L_{1},\mathcal{D}_{T}})\cdot\max_{g_{i}\in\mathcal{G}_{\delta}}\mathbb{P}\Bigg{\{}\left|\frac{1}{T}\sum_{t=1}^{T}U_{t}g_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\right|\geq-\delta+\frac{\epsilon}{4}(a+b)-\frac{\epsilon^{2}(a+b)}{4(1+\epsilon)B_{T}}$
$\qquad\qquad\qquad\qquad+\frac{\epsilon(1-\epsilon)}{4(1+\epsilon)B_{T}}\cdot\left(\frac{1}{T}\sum_{t=1}^{T}g_{i}^{2}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-2\delta B_{T}\right)\ \Big{|}\ \mathcal{D}_{T}\Bigg{\}}.$
Now set $\delta=\epsilon b/8$. Since $B_{T}\geq 2$ and $0<\epsilon\leq 1/2$, it follows that
$\delta+\frac{\epsilon b}{4}-\frac{\epsilon^{2}b}{4(1+\epsilon)B_{T}}-\frac{\epsilon(1-\epsilon)\delta}{2(1+\epsilon)}>0.$
Consequently,
$\mathbb{P}\Bigg{\{}\exists\ \mathbf{p}\in\mathcal{F}:\left|\frac{1}{T}\sum_{t=1}^{T}U_{t}g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\right|\geq\frac{\epsilon}{4}(a+b)+\frac{\epsilon(1-\epsilon)}{4(1+\epsilon)B_{T}}\cdot\frac{1}{T}\sum_{t=1}^{T}g_{\mathbf{p}}^{2}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})$
$\qquad\qquad\qquad\qquad-\frac{\epsilon^{2}(a+b)}{4(1+\epsilon)B_{T}}\ \Big{|}\ \mathcal{D}_{T}\Bigg{\}}$
$\leq N\left(\frac{\epsilon b}{8},\mathcal{F}_{1},\|\cdot\|_{L_{1},\mathcal{D}_{T}}\right)\cdot\max_{g_{i}\in\mathcal{G}_{\epsilon b/8}}\mathbb{P}\Bigg{\{}\left|\frac{1}{T}\sum_{t=1}^{T}U_{t}g_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\right|\geq\frac{\epsilon a}{4}-\frac{\epsilon^{2}a}{4(1+\epsilon)B_{T}}$
$\qquad\qquad\qquad\qquad+\frac{\epsilon(1-\epsilon)}{4(1+\epsilon)B_{T}}\cdot\frac{1}{T}\sum_{t=1}^{T}g_{i}^{2}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\ \Big{|}\ \mathcal{D}_{T}\Bigg{\}}.$ (22)
Step 4. Applying Bernstein’s inequality.
We write
$W_{t}=U_{t}g_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)}),\quad\sigma^{2}=\frac{1}{T}\sum_{t=1}^{T}var\{W_{t}|\mathbf{Z}^{(t)},\mathbf{y}^{(t)}\},$
$A_{1}=\frac{\epsilon a}{4}-\frac{\epsilon^{2}a}{4(1+\epsilon)B_{T}},\quad A_{2}=\frac{\epsilon(1-\epsilon)}{4(1+\epsilon)B_{T}}$
Since $g_{i}\in\mathcal{F}_{1}$, $B_{T}\geq 2$, and $0<\epsilon\leq 1/2$, it is easy to see that $|W_{t}|\leq B_{T}$ and $A_{1},A_{2}>0$. Applying Bernstein’s inequality, we have
$\mathbb{P}\Bigg{\{}\Big{|}\frac{1}{T}\sum_{t=1}^{T}U_{t}g_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\Big{|}\geq\frac{\epsilon a}{4}-\frac{\epsilon^{2}a}{4(1+\epsilon)B_{T}}+\frac{\epsilon(1-\epsilon)}{4(1+\epsilon)B_{T}}\frac{1}{T}\sum_{t=1}^{T}g^{2}_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\Big{|}\mathcal{D}_{T}\Bigg{\}}$ (23)
$=$ $\mathbb{P}\Bigg{\{}\Big{|}\frac{1}{T}\sum_{t=1}^{T}W_{t}\Big{|}\geq A_{1}+A_{2}\sigma^{2}\Big{|}\mathcal{D}_{T}\Bigg{\}}$
$\leq$ $2\exp\Bigg{(}-\frac{T(A_{1}+A_{2}\sigma^{2})^{2}}{2\sigma^{2}+\frac{2}{3}B_{T}(A_{1}+A_{2}\sigma^{2})}\Bigg{)}$
$=$ $2\exp\Bigg{(}-\frac{TA_{2}^{2}}{2B_{T}A_{2}/3}\cdot\frac{(A_{1}/A_{2}+\sigma^{2})^{2}}{A_{1}/A_{2}+(1+3/B_{T}A_{2})\sigma^{2}}\Bigg{)}.$
Some elementary calculations show that for any $c_{1},x>0$ and $c_{2}\geq 2$,
$\frac{(c_{1}+x)^{2}}{c_{1}+c_{2}x}\geq\frac{[c_{1}+c_{1}(c_{2}-2)/c_{2}]^{2}}{c_{1}+c_{2}\cdot c_{1}(c_{2}-2)/c_{2}}=4c_{1}\frac{c_{2}-1}{c_{2}^{2}}.$
Now let $c_{1}=A_{1}/A_{2}$, $x=\sigma^{2}$, and $c_{2}=1+3/B_{T}A_{2}$. Since $0<\epsilon\leq 1/2$, it follows that
$1+3/B_{T}A_{2}=1+\frac{12(1+\epsilon)}{\epsilon(1-\epsilon)}\geq 1+48>2.$
Consequently,
$\frac{TA_{2}^{2}}{2B_{T}A_{2}/3}\cdot\frac{(A_{1}/A_{2}+\sigma^{2})^{2}}{A_{1}/A_{2}+(1+3/B_{T}A_{2})\sigma^{2}}$ $\geq$ $\frac{TA_{2}^{2}}{2B_{T}A_{2}/3}\cdot 4\frac{A_{1}}{A_{2}}\frac{3/B_{T}A_{2}}{(1+3/B_{T}A_{2})^{2}}$
$=$ $\frac{18TA_{1}A_{2}}{(3+B_{T}A_{2})^{2}}.$
Also note that $A_{1}=\frac{\epsilon a}{4}-\frac{\epsilon^{2}a}{4(1+\epsilon)B_{T}}$ and $A_{2}=\frac{\epsilon(1-\epsilon)}{4(1+\epsilon)B_{T}}$. As $B_{T}\geq 2$ and $0<\epsilon\leq 1/2$, it follows that
$A_{1}=\frac{\epsilon a}{4}\Big{(}1-\frac{\epsilon}{(1+\epsilon)B_{T}}\Big{)}\geq\frac{\epsilon a}{4}\frac{3}{4}=\frac{3\epsilon a}{16}.$
Then we obtain that
$\frac{18TA_{1}A_{2}}{(3+B_{T}A_{2})^{2}}$ $\geq$ $18T\cdot\frac{3\epsilon a}{16}\cdot\frac{\epsilon(1-\epsilon)}{4(1+\epsilon)B_{T}}\cdot\frac{1}{(3+\frac{\epsilon(1-\epsilon)}{4(1+\epsilon)})^{2}}$
$\geq$ $\frac{27T\epsilon^{2}(1-\epsilon)a}{32(1+\epsilon)B_{T}}\cdot\frac{1}{(3+1/16)^{2}}$
$=$ $\frac{27\cdot 16^{2}}{32\cdot 49^{2}}\frac{T\epsilon^{2}(1-\epsilon)a}{(1+\epsilon)B_{T}}$
$\geq$ $\frac{\epsilon^{2}(1-\epsilon)}{12(1+\epsilon)B_{T}}aT.$
Combining this with (23), we obtain
$\mathbb{P}\Bigg{\{}\Big{|}\frac{1}{T}\sum_{t=1}^{T}U_{t}g_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\Big{|}\geq\frac{\epsilon a}{4}-\frac{\epsilon^{2}a}{4(1+\epsilon)B_{T}}+\frac{\epsilon(1-\epsilon)}{4(1+\epsilon)B_{T}}\frac{1}{T}\sum_{t=1}^{T}g^{2}_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\Big{|}\mathcal{D}_{T}\Bigg{\}}$
$\leq$ $2\exp\Bigg{(}-\frac{\epsilon^{2}(1-\epsilon)}{12(1+\epsilon)B_{T}}aT\Bigg{)}.$
Applying this result with (A.1), (21), and (A.1), we obtain
$I_{1}\leq 4\mathbb{E}\Bigg{(}N(\epsilon b/8,\mathcal{F}_{1},\|\cdot\|_{L_{1},\mathcal{D}_{T}})\exp\Big{(}-\frac{\epsilon^{2}(1-\epsilon)}{12(1+\epsilon)B_{T}}aT\Big{)}\Bigg{)}.$ (24)
Step 5. Bounding the covering numbers in (19) and (24).
Recall that $\mathcal{F}_{1}=\{g_{\mathbf{p}}(\mathbf{Z},\mathbf{y}):\mathbf{p}\in\mathcal{F}\}$, where
$g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})=\mathbf{y}^{\top}\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\mathbf{p}(\mathbf{Z})}\right)_{\geq\frac{C_{T}}{T}}=\sum_{k=1}^{K}y_{k}\left[B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z})}{p_{k}(\mathbf{Z})}\right)\right]\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})\geq\frac{C_{T}}{T}\right).$
Given $\mathcal{Z}_{T}=\{\mathcal{Z}_{1},\dots,\mathcal{Z}_{T}\}$, let $\log(\mathbf{p}^{1}),\dots,\log(\mathbf{p}^{M})$ with $M=\mathcal{N}(\delta,\log({\mathcal{F}}),\|\cdot\|_{\infty,\mathcal{Z}_{T}})$ be a $\delta$-covering of $\log(\mathcal{F})$ with respect to the norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$. Without loss of generality, we assume that $\mathbf{p}^{i}\in\mathcal{F}$. For any $\mathbf{p}\in\mathcal{F}$, there exists an $\mathbf{p}^{i}$ such that
$\|\log(\mathbf{p})-\log(\mathbf{p}^{i})\|_{\infty,\mathcal{Z}_{T}}\leq\delta,$
where
$\|\log(\mathbf{p})-\log(\mathbf{p}^{i})\|_{\infty,\mathcal{Z}_{T}}=\max_{1\leq t\leq T}\max_{1\leq k\leq K}|\log(\mathbf{p}_{k}(Z^{(t)}))-\log(\mathbf{p}^{i}_{k}(Z^{(t)}))|.$
Consequently,
$\|g_{\mathbf{p}}-g_{\mathbf{p}^{i}}\|_{L_{1},\mathcal{D}_{T}}$ $=$ $\frac{1}{T}\sum_{t=1}^{T}|g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-g_{\mathbf{p}^{i}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})|$
$=$ $\frac{1}{T}\sum_{t=1}^{T}\Bigg{|}\sum_{k=1}^{K}y_{k}^{(t)}\left[B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z}^{(t)})}{p_{k}(\mathbf{Z}^{(t)})}\right)\right]\mathbb{I}\left(p_{k}^{0}(\mathbf{Z}^{(t)})\geq\frac{C_{T}}{T}\right)$
$-\sum_{k=1}^{K}y_{k}^{(t)}\left[B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z}^{(t)})}{p^{i}_{k}(\mathbf{Z}^{(t)})}\right)\right]\mathbb{I}\left(p_{k}^{0}(\mathbf{Z}^{(t)})\geq\frac{C_{T}}{T}\right)\Bigg{|}$
$\leq$ $\frac{1}{T}\sum_{t=1}^{T}\sum_{k=1}^{K}y_{k}^{(t)}\Bigg{|}B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z}^{(t)})}{p_{k}(\mathbf{Z}^{(t)})}\right)-B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z}^{(t)})}{p^{i}_{k}(\mathbf{Z}^{(t)})}\right)\Bigg{|}$
$\leq$ $\frac{1}{T}\sum_{t=1}^{T}\sum_{k=1}^{K}y_{k}^{(t)}\Big{|}\log\left(p_{k}(\mathbf{Z}^{(t)})\right)-\log\left(p^{i}_{k}(\mathbf{Z}^{(t)})\right)\Big{|}$
$\leq$ $\|\log(\mathbf{p})-\log(\mathbf{p}^{i})\|_{\infty,\mathcal{Z}_{T}}\leq\delta.$
This implies that
$\mathcal{N}(\delta,\mathcal{F}_{1},\|\cdot\|_{L_{1},\mathcal{D}_{T}})\leq\mathcal{N}(\delta,\log({\mathcal{F}}),\|\cdot\|_{\infty,\mathcal{Z}_{T}}).$
Step 6. Conclusion.
The assertions (A.1), (19) and (24) imply, for $T\geq\frac{8B_{T}}{(a+b)\epsilon^{2}}$,
$\mathbb{P}\Big{\{}\exists\ \mathbf{p}\in\mathcal{F}:\ \mathbb{E}g_{\bf p}(\mathbf{Z,y})-\frac{1}{T}\sum_{t=1}^{T}g_{\bf p}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\geq\epsilon\{a+b+\mathbb{E}g_{\bf p}(\mathbf{Z,y})\}\Big{\}}$
$\leq$ $\frac{32}{7}\mathbb{E}\Bigg{(}\mathcal{N}(\epsilon b/8,\mathcal{F}_{1},\|\cdot\|_{L_{1},\mathcal{D}_{T}})\exp\Big{(}-\frac{\epsilon^{2}(1-\epsilon)}{12(1+\epsilon)B_{T}}aT\Big{)}\Bigg{)}$
$+\frac{32}{7}\mathbb{E}\Bigg{(}\mathcal{N}(\frac{(a+b)\epsilon}{5},\mathcal{F}_{1},\|\cdot\|_{L_{1},\mathcal{D}_{T}})\exp(-\frac{3\epsilon^{2}(a+b)T}{40B_{T}^{2}})\Bigg{)}$
$\leq$ $\frac{32}{7}\mathbb{E}\Bigg{(}\mathcal{N}(\epsilon b/8,\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}})\exp\Big{(}-\frac{\epsilon^{2}(1-\epsilon)}{12(1+\epsilon)B_{T}}aT\Big{)}\Bigg{)}$
$+\frac{32}{7}\mathbb{E}\Bigg{(}\mathcal{N}(\frac{(a+b)\epsilon}{5},\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}})\exp(-\frac{3\epsilon^{2}(a+b)T}{40B_{T}^{2}})\Bigg{)}$
$\leq$ $2\cdot\frac{32}{7}\sup_{\mathcal{Z}_{T}}\mathcal{N}(\epsilon b/8,\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}})\exp\Bigg{(}-\frac{\epsilon^{2}(1-\epsilon)aT}{15(1+\epsilon)B_{T}^{2}}\Bigg{)}$
$\leq$ $10\sup_{\mathcal{Z}_{T}}\mathcal{N}(\frac{\epsilon b}{8},\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}})\exp\Bigg{(}-\frac{\epsilon^{2}(1-\epsilon)aT}{15(1+\epsilon)B_{T}^{2}}\Bigg{)}.$
For $T<\frac{8B_{T}}{(a+b)\epsilon^{2}}$, it is easy to see that
$\exp\Bigg{(}-\frac{\epsilon^{2}(1-\epsilon)aT}{15(1+\epsilon)B_{T}^{2}}\Bigg{)}\geq\exp\Bigg{(}-\frac{\epsilon^{2}(1-\epsilon)a}{15(1+\epsilon)B_{T}^{2}}\cdot\frac{8B_{T}}{(a+b)\epsilon^{2}}\Bigg{)}\geq\exp(-\frac{8}{30})\geq\frac{1}{10}.$
Then for all $T$,
$\mathbb{P}\Big{\{}\exists\ \mathbf{p}\in\mathcal{F}:\ \mathbb{E}g_{\bf p}(\mathbf{Z,y})-\frac{1}{T}\sum_{t=1}^{T}g_{\bf p}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\geq\epsilon\{a+b+\mathbb{E}g_{\bf p}(\mathbf{Z,y})\}\Big{\}}$
$\leq$ $10\sup_{\mathcal{Z}_{T}}\mathcal{N}(\frac{\epsilon b}{8},\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}})\exp\Bigg{(}-\frac{\epsilon^{2}(1-\epsilon)aT}{15(1+\epsilon)B_{T}^{2}}\Bigg{)}.$
This completes the proof of this Lemma.
The following result is the Lemma 3.7 in Bos and Schmidt-Hieber (2022), which is used in Step 1 of the proof of Lemma 24 with $m=2$. We call $(p_{1},\dots,p_{K})\in\mathbb{R}^{K}$ a probability vector if $\sum_{i=1}^{K}p_{i}=1$ and $p_{i}\geq 0$ for all $1\leq i\leq K$.
Lemma 25
Given $B>2$ and the integer $m\geq 2$. Then, for any two probability vectors $(p_{1},\dots,p_{K})$ and $(q_{1},\dots,q_{K})$, we have
$\sum_{k=1}^{K}p_{k}\left|B\wedge\log\frac{p_{k}}{q_{k}}\right|^{m}\leq\max\left\{m!,\frac{B^{m}}{B-1}\right\}\sum_{k=1}^{K}p_{k}\left(B\wedge\log\frac{p_{k}}{q_{k}}\right).$
Appendix B Proof of Theorem 9
Proof  By composing the neural networks established in Lemma 8 and Lemma 27, we construct a neural network $\mathbf{G}(\boldsymbol{\phi})=\big{(}G(\phi_{1}),\dots,G(\phi_{K})\big{)}$ such that, for any $\varepsilon\in(0,1)$,
$\|\boldsymbol{\phi}-\mathbf{p}_{0}\|_{\infty}\leq\varepsilon\quad{\rm and}\quad\|e^{\mathbf{G}(\boldsymbol{\phi})}-\boldsymbol{\phi}\|_{\infty}\leq\varepsilon,$
where $\boldsymbol{\phi}=(\phi_{1},\phi_{2},\dots,\phi_{K})^{\top}\in\mathcal{T}(M,D,H,W_{1},B_{1},S_{1},\gamma)$ and $\mathbf{G}\in\mathcal{F}(L,KW,KS)$. Consequently, for any $k=1,\dots,K$, we have
$\left\|e^{G(\phi_{k})}-p_{k}^{0}\right\|_{\infty}\leq\left\|e^{G(\phi_{k})}-\phi_{k}\right\|_{\infty}+\left\|\phi_{k}-p_{k}^{0}\right\|_{\infty}\leq\varepsilon+\varepsilon=2\varepsilon.$
Now, define the vector-valued function $\widetilde{\mathbf{q}}=\big{(}\widetilde{q}_{1},\cdots,\widetilde{q}_{K}\big{)}^{\top}$ component-wisely as
$\widetilde{q}_{k}(\mathbf{Z})=\frac{e^{G(\phi_{k}(\mathbf{Z}))}}{\sum_{j=1}^{K}e^{G(\phi_{j}(\mathbf{Z}))}},\quad k=1,\dots,K.$
To bound $\left\|\widetilde{q}_{k}-p_{k}^{0}\right\|_{\infty}$, note that $\mathbf{p}_{0}=\big{(}p_{1}^{0},\dots,p_{K}^{0}\big{)}$ is a probability vector, and apply the triangle inequality to obtain
$\left\|\widetilde{q}_{k}-p_{k}^{0}\right\|_{\infty}$ $\leq$ $\left\|e^{G(\phi_{k})}\left(\frac{1}{\sum_{j=1}^{K}e^{G(\phi_{j})}}-1\right)\right\|_{\infty}+\left\|e^{G(\phi_{k})}-p_{k}^{0}\right\|_{\infty}$
$=$ $\left\|e^{G(\phi_{k})}\left(\frac{\sum_{\ell=1}^{K}p_{\ell}^{0}}{\sum_{j=1}^{K}e^{G(\phi_{j})}}-1\right)\right\|_{\infty}+\left\|e^{G(\phi_{k})}-p_{k}^{0}\right\|_{\infty}$
$\leq$ $\sum_{\ell=1}^{K}\left\|p_{\ell}^{0}-e^{G(\phi_{\ell})}\right\|_{\infty}\left\|\frac{e^{G(\phi_{k})}}{\sum_{j=1}^{K}e^{G(\phi_{j})}}\right\|_{\infty}+\left\|e^{G(\phi_{k})}-p_{k}^{0}\right\|_{\infty}$
$\leq$ $2K\varepsilon+2\varepsilon=2\varepsilon(K+1).$
Next, we investigate the low bound of $\widetilde{q}_{k}(\mathbf{Z})$. According to Lemma  27, we have $e^{G(\phi_{k})}\geq\varepsilon$. Also note that
$\left\|\sum_{j=1}^{K}e^{G(\phi_{j}(\mathbf{Z}))}-1\right\|_{\infty}=\left\|\sum_{j=1}^{K}\{e^{G(\phi_{j}(\mathbf{Z}))}-p_{j}^{0}\}\right\|_{\infty}\leq 2K\varepsilon.$
Consequently,
$\widetilde{q}_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2K\varepsilon+1}\geq\frac{\varepsilon}{2K+1}\geq\frac{\varepsilon}{2(K+1)}.$
This completes the proof of Theorem 9.
The approximation capabilities of deep ReLU networks have been extensively studied, establishing sharp rates for various function classes (Yarotsky, 2017; DeVore et al., 2021; Elbrächter et al., 2021; Berner et al., 2021; Siegel, 2023; Schmidt-Hieber, 2020; Shen et al., 2019; Lu et al., 2021; Shen et al., 2022; Jiao et al., 2023). A key component in our subsequent analysis is the efficient approximation of the logarithm function by a scalar-valued neural network with controlled complexity. To this end, we present the following lemma, which directly adapts the construction from Theorem 4.1 of Bos and Schmidt-Hieber (2022).
Lemma 26 (Scalar approximation of $\log x$)
For any $\varepsilon\in(0,1)$ and $\beta>0$, there exists a scalar-valued ReLU neural network $G:[0,1]\to\mathbb{R}$ such that
$\left|e^{G(x)}-x\right|\leq\varepsilon,\quad\text{and}\quad G(x)\geq\log(\varepsilon),\quad\text{for all }x\in[0,1],$
with network depth $L$, width $W$, and size $S$ (total number of non-zero parameters) satisfying
$L\lesssim\log\left(\frac{1}{\varepsilon}\right),\quad W\lesssim\left(\frac{1}{\varepsilon}\right)^{\frac{1}{\beta}},\quad S\lesssim\left(\frac{1}{\varepsilon}\right)^{\frac{1}{\beta}}\log\left(\frac{1}{\varepsilon}\right).$
Lemma 27 (Vectorized approximation via coordinate-wise application)
Let $K\geq 1$, and let $G$ be a scalar-valued neural network with depth $L$, width $W$, and at most $S$ nonzero parameters. Define the vector-valued network
$\mathbf{G}(\boldsymbol{\phi}(\mathbf{Z})):=\bigl{(}G(\phi_{1}(\mathbf{Z})),\dots,G(\phi_{K}(\mathbf{Z}))\bigr{)}^{\top},$
where $\boldsymbol{\phi}(\mathbf{Z})=(\phi_{1}(\mathbf{Z}),\dots,\phi_{K}(\mathbf{Z}))^{\top}\in[0,1]^{K}$. Then $\mathbf{G}\in\mathcal{F}_{\mathrm{id}}(L,KW,KS)$, and for all $k\in[K]$ and $\mathbf{Z}\in[0,1]^{d}$,
$\left|e^{G(\phi_{k}(\mathbf{Z}))}-\phi_{k}(\mathbf{Z})\right|\leq\varepsilon,\quad\text{and}\quad G(\phi_{k}(\mathbf{Z}))\geq\log(\varepsilon).$
Proof  Although $\mathcal{F}_{\mathrm{id}}(L,W,S)$ as defined in (5) refers to vector-valued neural networks mapping $0,1]^{K}\to\mathbb{R}^{K}$, we temporarily allow the notation to refer to scalar-valued networks when the context is clear.
To prove the lemma, it suffices to construct a scalar-valued ReLU neural network $G:[0,1]\to\mathbb{R}$ such that
$\left|e^{G(x)}-x\right|\leq\varepsilon\quad\text{and}\quad G(x)\geq\log(\varepsilon)\quad\text{for all }x\in[0,1].$
Given such a function $G$, we define the vector-valued network $\mathbf{G}:[0,1]^{K}\to\mathbb{R}^{K}$ by
$\mathbf{G}(\mathbf{x}):=\bigl{(}G(x_{1}),G(x_{2}),\dots,G(x_{K})\bigr{)}^{\top}.$
By construction, for each $k\in[K]$, the coordinate function satisfies
$\left|e^{G(x_{k})}-x_{k}\right|\leq\varepsilon,\quad\text{and}\quad G(x_{k})\geq\log(\varepsilon),\quad\text{for all }x_{k}\in[0,1],$
so $\mathbf{G}$ satisfies the desired approximation and positivity conditions coordinate-wise. Since the same network $G$ is applied independently to each coordinate, the resulting network $\mathbf{G}$ lies in the class $\mathcal{F}_{\mathrm{id}}(L,KW,KS)$, completing the proof.
Appendix C Proof of Theorem 11
To establish an upper bound for the risk $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$ in terms of the Kullback–Leibler (KL) divergence, we first provide two Lemmas that will be used in the proof of Theorem 11.
Lemma 28
Let $\mathbf{p}(\mathbf{Z})=(p_{1}(\mathbf{Z}),\dots,p_{K}(\mathbf{Z}))^{\top}$ be a conditional class probability function. Suppose there exist constants $\alpha\geq 0$ and $C<\infty$ such that
$\mathbb{P}_{\mathbf{Z}}\bigl{(}p_{k}(\mathbf{Z})\leq t\bigr{)}\leq C\,t^{\alpha},\quad\forall t\in[0,1],\,k\in\{1,\dots,K\}.$
Then for any $h\in(0,1]$, it holds that
$\int_{\{p_{k}(\mathbf{z})\geq h\}}\frac{1}{p_{k}(\mathbf{z})}\,d\mathbb{P}_{\mathbf{Z}}(\mathbf{z})\;\leq\;\begin{cases}\dfrac{C\,h^{\alpha-1}}{1-\alpha},&0\leq\alpha<1,\\[6.0pt] C\bigl{(}1-\log h\bigr{)},&\alpha\geq 1.\end{cases}$
Proof  Using the standard layer-cake representation, we have for a nonnegative measurable function $g$:
$\int g(\mathbf{z})\,d\mathbb{P}_{\mathbf{Z}}(\mathbf{z})=\int_{0}^{\infty}\mathbb{P}_{\mathbf{Z}}\bigl{(}g(\mathbf{Z})\geq u\bigr{)}\,du.$
Set
$g(\mathbf{z})=\frac{1}{p_{k}(\mathbf{z})}\mathbf{1}_{\{p_{k}(\mathbf{z})\geq h\}},$
thus
$\int_{\{p_{k}(\mathbf{z})\geq h\}}\frac{1}{p_{k}(\mathbf{z})}\,d\mathbb{P}_{\mathbf{Z}}(\mathbf{z})=\int_{0}^{\infty}\mathbb{P}_{\mathbf{Z}}\left(\frac{1}{p_{k}(\mathbf{Z})}\mathbb{I}_{\{p_{k}(\mathbf{Z})\geq h\}}\geq u\right)\,du.$
Since
$\frac{1}{p_{k}(\mathbf{Z})}\mathbf{1}_{\{p_{k}(\mathbf{Z})\geq h\}}\geq u>0\quad\Longrightarrow\quad h\leq p_{k}(\mathbf{Z})\leq\frac{1}{u},$
it follows that $u\in(0,1/h]$ and
$\int_{\{p_{k}(\mathbf{z})\geq h\}}\frac{1}{p_{k}(\mathbf{z})}\,d\mathbb{P}_{\mathbf{Z}}(\mathbf{z})\leq\int_{0}^{1/h}\mathbb{P}_{\mathbf{Z}}\bigl{(}p_{k}(\mathbf{Z})\leq\tfrac{1}{u}\bigr{)}\,du.$
To complete the proof, we analyze two cases separately:
Case 1: $0\leq\alpha<1$. Note that $C\geq 1$. Then, for all $u\in(0,1/h]$, we have
$\mathbb{P}_{\mathbf{Z}}\bigl{(}p_{k}(\mathbf{Z})\leq\tfrac{1}{u}\bigr{)}\leq C\,u^{-\alpha},$
and thus
$\int_{\{p_{k}(\mathbf{z})\geq h\}}\frac{1}{p_{k}(\mathbf{z})}\,d\mathbb{P}_{\mathbf{Z}}(\mathbf{z})\leq\int_{0}^{1/h}C\,u^{-\alpha}\,du=\frac{C\,h^{\alpha-1}}{1-\alpha}.$
Case 2: $\alpha\geq 1$. It follows that $\mathbb{P}_{\mathbf{Z}}(p_{k}(\mathbf{Z})\leq t)\leq Ct^{\alpha}\leq Ct$ for all $t\in[0,1]$ and $C\geq 1$. Then we can bound the integral as follows:
$\int_{\{p_{k}(\mathbf{z})\geq h\}}\frac{d\mathbb{P}_{\mathbf{Z}}(\mathbf{z})}{p_{k}(\mathbf{z})}$ $\leq\int_{0}^{1/h}\mathbb{P}_{\mathbf{Z}}\left(p_{k}(\mathbf{Z})\leq\frac{1}{u}\right)du$
$\leq\int_{0}^{1/h}\min\left(1,\frac{C}{u}\right)du$
$=\int_{0}^{C}1\,du+\int_{C}^{1/h}\frac{C}{u}\,du$
$=C+C\left(\log\left(\frac{1}{h}\right)-\log C\right)$
$=C\left(1-\log h-\log C\right)$
$\leq C(1-\log h),$
where the last inequality is due to $C\geq 1$. Combining both cases completes the proof.
The next lemma provides the relationship between the KL divergence and the Chi-square divergence.
Lemma 29 (KL-Chi-square Divergence Inequality)
Let $P$ and $Q$ be two probability measures satisfying $P\ll Q$ with the Radon-Nikodym derivative
$g(z)=\frac{dP}{dQ}(z).$
Then, the following inequality holds:
$\mathrm{KL}(P\|Q)\leq\chi^{2}(P\|Q),$
where $\mathrm{KL}(P\|Q)=\mathbb{E}_{Q}[g(z)\log g(z)]$ and $\chi^{2}(P\|Q)=\mathbb{E}_{Q}[(g(z)-1)^{2}]$.
Proof  Recall the elementary inequality
$\log(x)\leq x-1,\quad\text{for all }x>0.$
Multiplying both sides by the non-negative function $g(z)$, we obtain
$g(z)\log(g(z))\leq g(z)(g(z)-1).$
Expanding the right-hand side, we have
$g(z)(g(z)-1)=(g(z)-1)^{2}+(g(z)-1).$
Taking expectation with respect to $Q$ on both sides yields
$\mathbb{E}_{Q}[g(z)\log(g(z))]\leq\mathbb{E}_{Q}[(g(z)-1)^{2}]+\mathbb{E}_{Q}[g(z)-1].$
Since $\mathbb{E}_{Q}[g(z)]=1$, the second expectation vanishes. Consequently,
$\mathbb{E}_{Q}[g(z)\log(g(z))]\leq\mathbb{E}_{Q}[(g(z)-1)^{2}].$
This completes the proof of Lemma 29.
Proof  Recall that $R(\mathbf{p}_{0},\mathbf{p})=\mathbb{E}_{\mathbf{Z}}\left[\mathbf{p}_{0}(\mathbf{Z})^{\top}\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\mathbf{p}(\mathbf{Z})}\right].$ Applying Lemma 29 to our setting yields
$R(\mathbf{p}_{0},\mathbf{p})=\mathbb{E}_{\mathbf{Z}}\left[KL(\mathbf{p}_{0}(\mathbf{Z}),\mathbf{p}(\mathbf{Z}))\right]\leq\mathbb{E}_{\mathbf{Z}}\left[\sum_{k=1}^{K}\frac{(p_{k}^{0}(\mathbf{Z})-p_{k}(\mathbf{Z}))^{2}}{p_{k}(\mathbf{Z})}\right].$
Recall that
$\|\mathbf{p}-\mathbf{p}_{0}\|_{\infty}\leq 2(K+1)\varepsilon,\quad\text{and}\quad\min_{1\leq k\leq K}\inf_{\mathbf{Z}\in[0,1]^{d}}p_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2(K+1)}.$
Therefore, for any $k\in\{1,\dots,K\}$ and any $\mathbf{Z}\in[0,1]^{d}$, it holds that
$p_{k}(\mathbf{Z})\geq\max\left\{p_{k}^{0}(\mathbf{Z})-2(K+1)\varepsilon,\;\frac{\varepsilon}{2(K+1)}\right\}.$
Furthermore, if
$p_{k}^{0}(\mathbf{Z})-2(K+1)\varepsilon\geq\frac{\varepsilon}{2(K+1)},$
then we have
$p_{k}^{0}(\mathbf{Z})-2(K+1)\varepsilon\geq p_{k}^{0}(\mathbf{Z})\left(1-\frac{2(K+1)}{2(K+1)+\frac{1}{2(K+1)}}\right)=\frac{p_{k}^{0}(\mathbf{Z})}{1+4(K+1)^{2}}.$
This gives rise to the following upper bound:
$\begin{split}\frac{\left(p_{k}^{0}(\mathbf{Z})-p_{k}(\mathbf{Z})\right)^{2}}{p_{k}(\mathbf{Z})}&\leq 8(K+1)^{3}\varepsilon\cdot\mathbb{I}\left\{p_{k}^{0}(\mathbf{Z})\leq 2(K+1)\varepsilon+\frac{\varepsilon}{2(K+1)}\right\}\\ &+\frac{4(K+1)^{2}\left[1+4(K+1)^{2}\right]\varepsilon^{2}}{p_{k}^{0}(\mathbf{Z})}\cdot\mathbb{I}\left\{p_{k}^{0}(\mathbf{Z})>2(K+1)\varepsilon+\frac{\varepsilon}{2(K+1)}\right\}.\end{split}$
Taking the expectation on both sides with respect to $\mathbf{Z}$, we obtain
$\mathbb{E}_{\mathbf{Z}}\left[\frac{\left(p_{k}^{0}(\mathbf{Z})-p_{k}(\mathbf{Z})\right)^{2}}{p_{k}(\mathbf{Z})}\right]\leq 8(K+1)^{3}\varepsilon\,\mathbb{P}\left\{p_{k}^{0}(\mathbf{Z})\leq 2(K+1)\varepsilon+\frac{\varepsilon}{2(K+1)}\right\}$
$+4(K+1)^{2}\left[1+4(K+1)^{2}\right]\varepsilon^{2}\,\mathbb{E}_{\mathbf{Z}}\left[\frac{\mathbb{I}\left\{p_{k}^{0}(\mathbf{Z})>2(K+1)\varepsilon+\frac{\varepsilon}{2(K+1)}\right\}}{p_{k}^{0}(\mathbf{Z})}\right].$
By applying the $\alpha$-SVB condition to bound the first term and using Lemma 28 to bound the second term, we obtain
$\mathbb{E}_{\mathbf{Z}}\left[\frac{\left(p_{k}^{0}(\mathbf{Z})-p_{k}(\mathbf{Z})\right)^{2}}{p_{k}(\mathbf{Z})}\right]$
$\leq$ $C\,\varepsilon^{1+(\alpha\wedge 1)}\,8(K+1)^{3}\left(\frac{4(K+1)^{2}+1}{2(K+1)}\right)^{\alpha\wedge 1}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right),$
where $\alpha\wedge\beta=\min\{\alpha,\beta\}$. Also note that $0\leq\alpha\leq 1$ and $\frac{4(K+1)^{2}+1}{2(K+1)}\leq 3(K+1)$. Consequently,
$R(\mathbf{p}_{0},\mathbf{p})$ $=\mathbb{E}_{\mathbf{Z}}\left[\sum_{k=1}^{K}\frac{\left(p_{k}^{0}(\mathbf{Z})-p_{k}(\mathbf{Z})\right)^{2}}{p_{k}(\mathbf{Z})}\right]$
$\leq C\varepsilon^{1+(\alpha\wedge 1)}K(K+1)^{3}\left(\frac{4(K+1)^{2}+1}{2(K+1)}\right)^{\alpha\wedge 1}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\!\left(\frac{1}{\varepsilon}\right)\right)$
$\leq C\varepsilon^{1+(\alpha\wedge 1)}K(K+1)^{3+(\alpha\wedge 1)}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right)$
$=C\varepsilon^{1+\alpha}K(K+1)^{3+\alpha}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right)$
$\leq C\varepsilon^{1+\alpha}K^{4+\alpha}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right).$
Now choose $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}\,K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$ with $C>0$ being a sufficiently small constant. This yields
$\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})\lesssim T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}K^{4+\alpha}\log(KT).$
This concludes the proof of Theorem  11.
Appendix D Proof of Lemma 12
Proof  We first establish a useful Lipschitz continuity property of the log-softmax function. This intermediate result is critical for bounding covering numbers in Lemma 12.
Lemma 30 (Lipschitz Continuity of Log-Softmax)
Let $\mathbf{y}=(y_{1},\dots,y_{K})\in\mathbb{R}^{K}$, and define the softmax function $\mathbf{\Phi}:\mathbb{R}^{K}\to\Delta^{K-1}$ by
$\mathbf{\Phi}(\mathbf{y}):=\frac{1}{\sum_{j=1}^{K}\exp(y_{j})}\begin{bmatrix}\exp(y_{1})\\ \vdots\\ \exp(y_{K})\end{bmatrix}$,
where $\Delta^{K-1}$ is the probability simplex in $\mathbb{R}^{K}$. Then $\log\circ\mathbf{\Phi}$ is $K$-Lipschitz with respect to the $\|\cdot\|_{\infty}$-norm, i.e.,
$\|\log(\mathbf{\Phi}(\mathbf{y}_{1}))-\log(\mathbf{\Phi}(\mathbf{y}_{2}))\|_{\infty}\leq K\|\mathbf{y}_{1}-\mathbf{y}_{2}\|_{\infty},\quad\forall\,\mathbf{y}_{1},\mathbf{y}_{2}\in\mathbb{R}^{K}.$
Proof  Consider the log-softmax function
$\log(\mathbf{\Phi}(\mathbf{y}))_{i}=\mathbf{y}_{i}-\log\left(\sum_{j=1}^{K}e^{\mathbf{y}_{j}}\right).$
Taking derivatives, for $i,k\in\{1,\dots,K\}$, we have
$\frac{\partial}{\partial\mathbf{y}_{k}}\log(\mathbf{\Phi}(\mathbf{y}))_{i}=\delta_{ik}-\mathbf{\Phi}(\mathbf{y})_{k},$
where $\delta_{ik}$ is the Kronecker delta. As $0\leq\mathbf{\Phi}(\mathbf{y})_{k}\leq 1$, the Jacobian’s entries $\delta_{ik}-\mathbf{\Phi}(\mathbf{y})_{k}$ are bounded by 1 in absolute value. Thus, the Lipschitz constant under the $\ell_{\infty}$-norm is at most $K$.
We now define an auxiliary function class excluding the softmax layer:
$\mathcal{F}^{\prime}=\{\mathbf{f}\circ\boldsymbol{\phi}(\mathbf{Z}):\mathbf{f}\in\mathcal{F}_{id}(L,W,S),\;\boldsymbol{\phi}\in\mathcal{T}(M,D,H,W,S,B_{1},B_{2},\gamma)\}.$
This allows us to represent the original function class conveniently as
$\log\mathcal{F}=\log(\mathbf{\Phi}(\mathcal{F}^{\prime})).$
The next lemma establish the covering number bound of $log\mathcal{F}$ with respect to the norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$.
Lemma 31
Given a sample $\mathcal{Z}_{T}=\{\mathbf{Z}^{(1)},\dots,\mathbf{Z}^{(T)}\}\subset[0,1]^{d}$, then for any $\delta>0$, we have
$\mathcal{N}\left(\delta,\log\mathcal{F},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\mathcal{N}\left(\tfrac{\delta}{2K},\mathcal{F}^{\prime},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right).$
Proof  Fix $\delta>0$. Let $\{\mathbf{f}_{j}\}_{j=1}^{J}$ be the centers of a minimal $\delta/(2K)$-cover of $\mathcal{F}^{\prime}$ under the empirical sup-norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$. That is, for every $\mathbf{f}\in\mathcal{F}^{\prime}$, there exists $j\in\{1,\dots,J\}$ such that
$\|\mathbf{f}-\mathbf{f}_{j}\|_{\infty,\mathcal{Z}_{T}}\leq\tfrac{\delta}{2K}.$
For each center $\mathbf{f}_{j}$, choose $\hat{\mathbf{f}}_{j}\in\mathcal{F}^{\prime}$ such that
$\|\mathbf{f}_{j}-\hat{\mathbf{f}}_{j}\|_{\infty,\mathcal{Z}_{T}}\leq\tfrac{\delta}{2K}.$
By the triangle inequality, for any $\mathbf{f}\in\mathcal{F}^{\prime}$, there exists $\hat{\mathbf{f}}_{j}\in\mathcal{F}^{\prime}$ such that
$\|\mathbf{f}-\hat{\mathbf{f}}_{j}\|_{\infty,\mathcal{Z}_{T}}\leq\|\mathbf{f}-\mathbf{f}_{j}\|_{\infty,\mathcal{Z}_{T}}+\|\mathbf{f}_{j}-\hat{\mathbf{f}}_{j}\|_{\infty,\mathcal{Z}_{T}}\leq\tfrac{\delta}{K}.$
Now consider an arbitrary $\mathbf{g}\in\log(\mathcal{F})$, so that $\mathbf{g}=\log\circ(\mathbf{\Phi}\circ\mathbf{f})$ for some $\mathbf{f}\in\mathcal{F}^{\prime}$. Then for the associated $\hat{\mathbf{f}}_{j}$, we have
$\|\mathbf{g}-\log\circ(\mathbf{\Phi}\circ\hat{\mathbf{f}}_{j})\|_{\infty,\mathcal{Z}_{T}}=\|\log\circ(\mathbf{\Phi}\circ\mathbf{f})-\log\circ(\mathbf{\Phi}\circ\hat{\mathbf{f}}_{j})\|_{\infty,\mathcal{Z}_{T}}\leq K\|\mathbf{f}-\hat{\mathbf{f}}_{j}\|_{\infty,\mathcal{Z}_{T}}\leq\delta,$
where we used the $K$-Lipschitz continuity of $\log\circ\mathbf{\Phi}$ (Lemma 30) with respect to $\|\cdot\|_{\infty}$-norm. Since $\mathbf{g}\in\log(\mathcal{F})$ was arbitrary and each $\hat{\mathbf{f}}_{j}\in\mathcal{F}^{\prime}$, the set $\{\log\circ(\mathbf{\Phi}\circ\hat{\mathbf{f}}_{j})\}_{j=1}^{J}\subset\log(\mathcal{F})$ forms a $\delta$-cover of $\log(\mathcal{F})$ under $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$. The claim of Lemma 31 follows.
Lemma 32 (Empirical Covering Bound for Composed Class)
Let $\mathcal{T}$ be a class of functions from $0,1]^{d}$ to $\mathbb{R}^{K}$, and let $\mathcal{F}_{\mathrm{id}}$ be a class of functions from $\mathbb{R}^{K}$ to $\mathbb{R}^{K}$, such that every $\mathbf{f}\in\mathcal{F}_{\mathrm{id}}$ is $L_{1}$-Lipschitz with respect to the sup-norm $\|\cdot\|_{\infty}$. Define the composed function class
$\mathcal{F}^{\prime}:=\left\{\mathbf{f}\circ\boldsymbol{\phi}:\mathbf{f}\in\mathcal{F}_{\mathrm{id}},\;\boldsymbol{\phi}\in\mathcal{T}\right\}.$
Then, for any empirical sample $\mathcal{Z}_{T}=\{\mathbf{Z}^{(1)},\ldots,\mathbf{Z}^{(T)}\}\subset[0,1]^{d}$, we have
$\mathcal{N}\left(\tfrac{\delta}{2K},\mathcal{F}^{\prime},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\mathcal{N}\left(\tfrac{\delta}{4KL_{1}},\mathcal{T},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\cdot\mathcal{N}\left(\tfrac{\delta}{4K},\mathcal{F}_{\mathrm{id}},\|\cdot\|_{\infty}\right).$
Proof  Fix $\delta>0$. Let $\widehat{\mathcal{T}}$ be a $\tfrac{\delta}{4KL_{1}}$-cover of $\mathcal{T}$ with respect to the empirical sup-norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$, and let $\widehat{\mathcal{F}}_{\mathrm{id}}$ be a $\tfrac{\delta}{4K}$-cover of $\mathcal{F}_{\mathrm{id}}$ with respect to the sup-norm $\|\cdot\|_{\infty}$. Define the composite approximation set
$\widehat{\mathcal{F}^{\prime}}:=\left\{\hat{\mathbf{f}}\circ\hat{\boldsymbol{\phi}}:\hat{\mathbf{f}}\in\widehat{\mathcal{F}}_{\mathrm{id}},\;\hat{\boldsymbol{\phi}}\in\widehat{\mathcal{T}}\right\}$.
Next we show that $\widehat{\mathcal{F}^{\prime}}$ is a $\tfrac{\delta}{2K}$-cover of $\mathcal{F}^{\prime}$ with respect to the empirical sup-norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$. Indeed, for any $\mathbf{f}\in\mathcal{F}_{\mathrm{id}}$ and $\boldsymbol{\phi}\in\mathcal{T}$, there exist $\hat{\mathbf{f}}\in\widehat{\mathcal{F}}_{\mathrm{id}}$ and $\hat{\boldsymbol{\phi}}\in\widehat{\mathcal{T}}$ such that
$\|\boldsymbol{\phi}-\hat{\boldsymbol{\phi}}\|_{\infty,\mathcal{Z}_{T}}\leq\tfrac{\delta}{4KL_{1}}\quad\text{and}\quad\|\mathbf{f}-\hat{\mathbf{f}}\|_{\infty}\leq\tfrac{\delta}{4K}.$
Then for each $t\in\{1,\ldots,T\}$, we have
$\|\mathbf{f}(\boldsymbol{\phi}(\mathbf{Z}^{(t)}))-\hat{\mathbf{f}}(\hat{\boldsymbol{\phi}}(\mathbf{Z}^{(t)}))\|_{\infty}$ $\leq\|\mathbf{f}(\boldsymbol{\phi}(\mathbf{Z}^{(t)}))-\mathbf{f}(\hat{\boldsymbol{\phi}}(\mathbf{Z}^{(t)}))\|_{\infty}+\|\mathbf{f}(\hat{\boldsymbol{\phi}}(\mathbf{Z}^{(t)}))-\hat{\mathbf{f}}(\hat{\boldsymbol{\phi}}(\mathbf{Z}^{(t)}))\|_{\infty}$
$\leq L_{1}\|\boldsymbol{\phi}(\mathbf{Z}^{(t)})-\hat{\boldsymbol{\phi}}(\mathbf{Z}^{(t)})\|_{\infty}+\tfrac{\delta}{4K}$
$\leq L_{1}\tfrac{\delta}{4KL_{1}}+\tfrac{\delta}{4K}=\tfrac{\delta}{2K}.$
Taking the maximum over $t=1,\ldots,T$ yields
$\|\mathbf{f}\circ\boldsymbol{\phi}-\hat{\mathbf{f}}\circ\hat{\boldsymbol{\phi}}\|_{\infty,\mathcal{Z}_{T}}\leq\tfrac{\delta}{2K},$
showing that $\widehat{\mathcal{F}^{\prime}}$ is indeed a $\tfrac{\delta}{2K}$-cover of $\mathcal{F}^{\prime}$. Consequently,
$\mathcal{N}\left(\tfrac{\delta}{2K},\mathcal{F}^{\prime},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq|\widehat{\mathcal{T}}|\cdot|\widehat{\mathcal{F}}_{\mathrm{id}}|=\mathcal{N}\left(\tfrac{\delta}{4KL_{1}},\mathcal{T},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\cdot\mathcal{N}\left(\tfrac{\delta}{4K},\mathcal{F}_{\mathrm{id}},\|\cdot\|_{\infty}\right).$
This completes the proof of Lemma 32.
Finally, by Lemma 31 and Lemma 32, we obtain
$\mathcal{N}\left(\delta,\log\mathcal{F},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)$ $\leq$ $\mathcal{N}\left(\tfrac{\delta}{2K},\mathcal{F}^{\prime},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)$
$\leq$ $\mathcal{N}\left(\tfrac{\delta}{4KL_{1}},\mathcal{T},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\cdot\mathcal{N}\left(\tfrac{\delta}{4K},\mathcal{F}_{\mathrm{id}},\|\cdot\|_{\infty}\right).$
Applying the logarithm to both sides and taking the supremum over all empirical samples $\mathcal{Z}_{T}\subset[0,1]^{d}$, we obtain
$\mathcal{V}_{\infty}\left(\delta,\,\log\mathcal{F},\,T\right)\leq\mathcal{V}_{\infty}\left(\tfrac{\delta}{4KL_{1}},\,\mathcal{T},\,T\right)+\mathcal{V}\left(\tfrac{\delta}{4K},\,\mathcal{F}_{\mathrm{id}},\,\|\cdot\|_{\infty}\right).$
Thus, we complete the proof of Lemma 12 in the main text.
Appendix E Proof of Lemma 13
Before presenting the detailed proof of Lemma 13, we first introduce the notion of pseudo-dimension which will be used in the proof of Lemma  13.
Definition 33
(Pseudo-dimension) Let $\mathcal{H}$ be a real-valued function class defined on $\mathbb{R}^{d}$. The pseudo-dimension of $\mathcal{H}$, denoted by ${\rm Pdim}(\mathcal{H})$, is the largest integer $m$ for which there exist $\{\mathbf{z}_{1},\dots,\mathbf{z}_{m}\}\subset\mathbb{R}^{d}$ and $\{y_{1},\dots,y_{m}\}\subset\mathbb{R}$ such that
$\left|\{(\operatorname{sgn}(g(\mathbf{z}_{1})-y_{1}),\dots,\operatorname{sgn}(g(\mathbf{z}_{m})-y_{m})):g\in\mathcal{H}\}\right|=2^{m}.$
Proof [Proof of Lemma 13.] Our proof for Lemma 13 is similar to that of Lemma 30 in Jiao et al. (2024), Lemma 8 in Gurevych et al. (2022), and Theorem 6 in Bartlett et al. (2019). The main idea is to bound the pseudo-dimension of subsets of $\mathcal{T}$ and then utilize well-known results developed by Anthony and Bartlett (2009) to bound the covering entropy by pseudo-dimensions. Here $\mathcal{T}=\mathcal{T}(M,H,D,W_{1},B_{1},S_{1},\gamma)$ defined in (3.3) in the main text.
Recall that $\mathcal{T}$ consists of functions mapping from $\mathbb{R}^{d}$ to $\mathbb{R}^{K}$. Our objective is to derive an upper bound on the pseudo-dimension of subsets of $\mathcal{T}$, in which the positions of the nonzero parameters are fixed. We first establish the bound for the spacial case $K=1$ and then generalize the result to arbitrary $K\geq 1$.
According to the definition of $\mathcal{T}$ with $K=1$, each function $g\in\mathcal{T}$ has at most $S_{1}$ nonzero parameters. Fixing the positions of these nonzero parameters, we denote their values by $\boldsymbol{\theta}\in\mathbb{R}^{S_{1}}$ and define the function class:
$\mathcal{G}=\{g(\cdot,\boldsymbol{\theta}):\mathbb{R}^{d}\to\mathbb{R}\mid g\in\mathcal{T},\boldsymbol{\theta}\in\mathbb{R}^{S_{1}}\},$
To bound the pseudo-dimension of $\mathcal{G}$, consider a set of input-output pairs $\{(\mathbf{Z}^{(i)},y_{i})\}_{i=1}^{m}\subseteq\mathbb{R}^{d}\times\mathbb{R}$ such that
$\left|\left\{\left(\operatorname{sgn}(g(\mathbf{Z}^{(1)},\boldsymbol{\theta})-y_{1}),\dots,\operatorname{sgn}(g(\mathbf{Z}^{(m)},\boldsymbol{\theta})-y_{m})\right):g\in\mathcal{G}\right\}\right|=2^{m}.$
Fixing the inputs $\{\mathbf{Z}^{(i)}\}_{i=1}^{m}$ and viewing the network parameters $\boldsymbol{\theta}$ as a set of $S_{1}$ real-valued variables, we aim to bound such an integer $m$, which gives an upper bound of $\operatorname{Pdim}(\mathcal{G})$.
Specifically, we define a sequence of partitions $\mathcal{S}_{0},\mathcal{S}_{1},\mathcal{S}_{2},\dots,\mathcal{S}_{M},\mathcal{S}_{M+1}$ of the parameter domain $\mathbb{R}^{S_{1}}$, constructed iteratively through successive refinement, such that in the final partition, for any region $P\in\mathcal{S}_{M+1}$, the functions
$g(\mathbf{Z}^{(1)},\boldsymbol{\theta}),\dots,g(\mathbf{Z}^{(m)},\boldsymbol{\theta})$
are polynomials in $\boldsymbol{\theta}$ of degree at most $10^{M+2}$ for all $\boldsymbol{\theta}\in P$. Note that for the partition $\mathcal{S}_{M+1}$, we have the following bound
$2^{m}$ $=$ $\left|\left\{\left(\operatorname{sgn}(g(\mathbf{Z}^{(1)},\boldsymbol{\theta})-y_{1}),\dots,\operatorname{sgn}(g(\mathbf{Z}^{(m)},\boldsymbol{\theta})-y_{m})\right):g\in\mathcal{G}\right\}\right|$ (25)
$\leq$ $\sum_{P\in\mathcal{S}_{M+1}}\left|\left\{\left(\operatorname{sgn}(g(\mathbf{Z}^{(1)},\boldsymbol{\theta})-y_{1}),\dots,\operatorname{sgn}(g(\mathbf{Z}^{(m)},\boldsymbol{\theta})-y_{m})\right):\boldsymbol{\theta}\in P\right\}\right|.$
Lemma 34 in the following ensures that each term in the sum of (25) can be effectively bounded, which provides an upper bound for $m$.
Next, we construct the sequence of partitions $\mathcal{S}_{0},\mathcal{S}_{1},\mathcal{S}_{2},\dots,\mathcal{S}_{M},\mathcal{S}_{M+1}$ layer by layer. For the input layer, define $\mathcal{S}_{0}=\{\mathbb{R}^{S_{1}}\}$. Then all components of $\mathbf{Z}_{0}=\mathbf{Z}$ are polynomial as functions of $\boldsymbol{\theta}$, each with degree 0, for all $\boldsymbol{\theta}\in\mathbb{R}^{S_{1}}$.
Consider $r\in\{1,\ldots,M\}$. Suppose inductively that for each set $S\in\mathcal{S}_{r-1}$, all components of $Z_{r-1}$ are polynomial functions in $\boldsymbol{\theta}$ of degree at most $10^{r}$ for $\boldsymbol{\theta}\in S$. Then any component of
$\boldsymbol{q}_{r-1,s,i}=\mathbf{W}_{Q,r,s}\boldsymbol{z}_{r-1,i},\quad\boldsymbol{k}_{r-1,s,i}=\mathbf{W}_{K,r,s}\boldsymbol{z}_{r-1,i}$
is a polynomial function in $\mathbf{\theta}$ of degree at most $10^{r}+1$ on each set $S\in\mathcal{S}_{r-1}$. Consequently, for each $S\in\mathcal{S}_{r-1}$, the inner product
$\langle\boldsymbol{q}_{r-1,s,i},\boldsymbol{k}_{r-1,s,j}\rangle$
is a polynomial function in $\boldsymbol{\theta}$ with degree at most $2(10^{r}+1)$. By applying Lemma 34, the differences
$\left\{\langle\boldsymbol{q}_{r-1,s,i},\boldsymbol{k}_{r-1,s,j_{1}}\rangle-\langle\boldsymbol{q}_{r-1,s,i},\boldsymbol{k}_{r-1,s,j_{2}}\rangle:s=1,\ldots,H,\ i,j_{1},j_{2}=1,\ldots,N+1\right\}$
have at most
$\Delta_{1}=2\left(\frac{2eH(N+1)^{3}(2\cdot 10^{r}+2)}{S_{1}}\right)^{S_{1}}$
distinct sign patterns.
We refine each set in $\mathcal{S}_{r-1}$ into at most $\Delta_{1}$ subsets so that within each subset, these polynomial differences maintain consistent signs. Consequently, on each subset, every component of
$\mathbf{W}_{O,r,s}\left(\mathbf{W}_{V,r,s}\mathbf{Z}_{r-1}\right)\left[(\mathbf{W}_{K,r,s}\mathbf{Z}_{r-1})^{\top}(\mathbf{W}_{Q,r,s}\mathbf{Z}_{r-1})\odot\sigma_{H}\left((\mathbf{W}_{K,r,s}\mathbf{Z}_{r-1})^{\top}(\mathbf{W}_{Q,r,s}\mathbf{Z}_{r-1})\right)\right]$
is a polynomial with degree at most $3\cdot 10^{r}+4$. Therefore, all components of
$\mathbf{Y}_{r}=F^{(SA)}(\mathbf{Z}_{r-1})$
are polynomials with degree at most $3\cdot 10^{r}+4$ within each refined region. On each subset within the new partition, every component of
$\mathbf{W}_{F_{1}}F^{(SA)}(\mathbf{Z}_{r-1})$
is represented by a polynomial with degree at most $3\cdot 10^{r}+5$. By invoking Lemma 34 again, each subset of this partition can be further refined into
$\Delta_{2}=2\left(\frac{2eW_{1}\left(3\cdot 10^{r}+5\right)}{S_{1}}\right)^{S_{1}}$
subsets, ensuring all components in $\mathbf{W}_{F_{1}}F^{(SA)}(\mathbf{Z}_{r-1})$ maintain consistent sign patterns within this refined partition. The resulting partition, obtained by two successive refinements of $\mathcal{S}_{r-1}$, is denoted as $\mathcal{S}_{r}$. Since within each subset of $\mathcal{S}_{r}$, the signs of all components remain unchanged, we conclude that every component of
$\sigma\left(\mathbf{W}_{F_{1}}F^{(SA)}(\mathbf{Z}_{r-1})\right)$
is either identically zero or represented by a polynomial of degree at most $3\cdot 10^{r}+5$. Consequently, within each subset in $\mathcal{S}_{r}$, every component of the output
$\mathbf{Z}_{r}=F^{(\mathrm{FF})}\left(F^{(SA)}(\mathbf{Z}_{r-1})\right)=F^{(SA)}(\mathbf{Z}_{r-1})+\mathbf{W}_{F_{2}}\sigma\left(\mathbf{W}_{F_{1}}F^{(SA)}(\mathbf{Z}_{r-1})\right)$
is given by a polynomial whose degree at most
$3\cdot 10^{r}+6\leq 10^{r+1}.$
Repeat the above steps, we obtain a partition $\mathcal{S}_{M}$ of $\mathbb{R}^{S_{1}}$ such that on each set $S\in\mathcal{S}_{M}$ all components of
$\mathbf{Z}_{M}$
are polynomials as functions of $\boldsymbol{\theta}\in S$ of degree no more than $10^{M+1}$. Consequently, for all $k\in\{1,\ldots,m\}$
$g(\mathbf{Z}^{(k)},\boldsymbol{\theta})-y_{k}$
are polynomials of degree at most $10^{M+1}+1\leq 10^{M+2}$ in $\boldsymbol{\theta}$ for $\boldsymbol{\theta}\in S$.
According to the above refinement process, we have
$\left|\mathcal{S}_{M}\right|=\prod_{r=1}^{M}\frac{\left|\mathcal{S}_{r}\right|}{\left|\mathcal{S}_{r-1}\right|}\leq\prod_{r=1}^{M}2\left(\frac{2eH(N+1)^{3}\left(2\cdot 10^{r}+2\right)}{S_{1}}\right)^{S_{1}}2\left(\frac{2eW_{1}\left(3\cdot 10^{r}+5\right)}{S_{1}}\right)^{S_{1}}.$
Note that
$\begin{aligned} &\left|\left\{\left(\operatorname{sgn}\left(g(\mathbf{Z}^{(1)},\boldsymbol{\theta})-y_{1}\right),\ldots,\operatorname{sgn}\left(g(\mathbf{Z}^{(m)},\boldsymbol{\theta})-y_{m}\right)\right):g\in\mathcal{G}\right\}\right|\\ &\leq\sum_{S\in\mathcal{S}_{M}}\left|\left\{\left(\operatorname{sgn}\left(g(\mathbf{Z}^{(1)},\boldsymbol{\theta})-y_{1}\right),\ldots,\operatorname{sgn}\left(g(\mathbf{Z}^{(m)},\boldsymbol{\theta})-y_{m}\right)\right):\boldsymbol{\theta}\in S\right\}\right|.\end{aligned}$
It follows from Lemma 34 that
$\begin{aligned} 2^{m}&=\left|\left\{\left(\operatorname{sgn}\left(g(\mathbf{Z}^{(1)},\boldsymbol{\theta})-y_{1}\right),\ldots,\operatorname{sgn}\left(g(\mathbf{Z}^{(m)},\boldsymbol{\theta})-y_{m}\right)\right):g\in\mathcal{G}\right\}\right|\\ &\leq 2\left(\frac{2em10^{M+2}}{S_{1}}\right)^{S_{1}}|\mathcal{S}_{M}|\\ &\leq 2\left(\frac{2em10^{M+2}}{S_{1}}\right)^{S_{1}}\prod_{r=1}^{M}2\left(\frac{2eH(N+1)^{3}\left(2\cdot 10^{r}+2\right)}{S_{1}}\right)^{S_{1}}2\left(\frac{2eW_{1}\left(3\cdot 10^{r}+5\right)}{S_{1}}\right)^{S_{1}}\\ &\leq 2^{2M+1}\left(\frac{m6e(\max\{H(N+1),W_{1}\})^{3}10^{M+2}}{(2M+1)S_{1}}\right)^{(2M+1)S_{1}}.\end{aligned}$
When $m\geq(2M+1)S_{1}$, by Lemma 35 in the following, we have
$m$ $\leq$ $(2M+1)+(2M+1)S_{1}\log_{2}\Big{\{}12e(\max\{H(N+1),W_{1}\})^{3}10^{M+2}$
$\times\log_{2}\Big{(}6e(\max\{H(N+1),W_{1}\})^{3}10^{M+2}\Big{)}\Big{\}}$
$\leq$ $CS_{1}M^{2}\log\big{(}\max\{H(N+1),W_{1}\}\big{)}$
$\leq$ $CS_{1}NM^{2}\log\big{(}\max\{H,W_{1}\}\big{)},$
where $C$ is an universal constant. When $m\leq(2M+1)S_{1}$, it is readily seen that $m\leq CS_{1}NM^{2}\log\big{(}\max\{H,W_{1}\}\big{)}$ also holds for some positive constant $C$. This implies
$\operatorname{Pdim}(\mathcal{G})\leq CS_{1}NM^{2}\log\big{(}\max\{H,W_{1}\}\big{)}.$
Applying Lemma 36, the covering number of the function class $\mathcal{G}$ with respect to $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$-norm is bounded as
$\mathcal{N}(\delta_{1},\mathcal{G},\|\cdot\|_{\infty,\mathcal{Z}_{T}})\leq\left(\frac{eTB_{1}}{\delta_{1}\operatorname{Pdim}(\mathcal{G})}\right)^{\operatorname{Pdim}(\mathcal{G})},\ \text{for}\ T\geq\operatorname{Pdim}(\mathcal{G}).$
Next, we consider the case $T\leq\operatorname{Pdim}(\mathcal{G})$. Write $\mathcal{B}_{B_{1}}=\{\mathbf{x}=(x_{1},\dots,x_{T})^{\top}\in\mathbb{R}^{T}:\|\mathbf{x}\|_{\infty}\leq B_{1}\}$, where $\|\mathbf{x}\|_{\infty}=\max_{1\leq i\leq T}|x_{i}|$. Similar to the argument of Lemma 1.18 in Rigollet and Hütter (2023), we have
$\mathcal{N}(\delta_{1},\mathcal{B}_{B_{1}},\|\cdot\|_{\infty})\leq\left(\frac{3B_{1}}{\delta_{1}}\right)^{T}.$
Since $\sup_{\mathbf{Z}}|g(\mathbf{Z},\mathbf{\theta})|\leq B_{1}$, it follows that
$\mathcal{N}(\delta_{1},\mathcal{G},\|\cdot\|_{\infty,\mathcal{Z}_{T}})\leq\mathcal{N}(\delta_{1},\mathcal{B}_{B_{1}},\|\cdot\|_{\infty})\leq\left(\frac{3B_{1}}{\delta_{1}}\right)^{T}\leq\left(\frac{3B_{1}}{\delta_{1}}\right)^{{\rm Pdim}(\mathcal{G})},\ \text{for}\ T<{\rm Pdim}(\mathcal{G}).$
Therefore, in all cases, we have
$\mathcal{N}(\delta_{1},\mathcal{G},\|\cdot\|_{\infty,\mathcal{Z}_{T}})\leq\left(\frac{3eTB_{1}}{\delta_{1}}\right)^{{\rm Pdim}(\mathcal{G})}.$
Note that the functions in $\mathcal{T}$ depend on at most $CMH^{2}\left(\max\{D,W_{1},N+1\}\right)^{3}$ parameters, of which at most $S_{1}$ parameters can be nonzero. Thus, the number of ways to choose these parameters is bounded by
$\binom{CMH^{2}\left(\max\{D,W_{1},N+1\}\right)^{3}}{S_{1}}\leq\left(CMH^{2}\left(\max\{D,W_{1},N+1\}\right)^{3}\right)^{S_{1}}.$
Consequently,
$\mathcal{N}(\delta_{1},\mathcal{T},\|\cdot\|_{\infty,\mathcal{Z}_{T}})\leq\left(CMH^{2}\left(\max\{D,W_{1},N+1\}\right)^{3}\right)^{S_{1}}\left(\frac{3eTB_{1}}{\delta_{1}}\right)^{{\rm Pdim}(\mathcal{G})}.$
Recall that $\operatorname{Pdim}(\mathcal{G})\leq CS_{1}NM^{2}\log\big{(}\max\{H,W_{1}\}\big{)}$. It follows that
$\log\mathcal{N}(\delta_{1},\mathcal{T},\|\cdot\|_{\infty,\mathcal{Z}_{T}})$
$\leq$ ${\rm Pdim}(\mathcal{G})\log\left(\frac{3eTB_{1}}{\delta_{1}}\right)+S_{1}\log\left(CMH^{2}\left(\max\{D,W_{1},N+1\}\right)^{3}\right)$
$\leq$ $CS_{1}NM^{2}\log(\max\{H,W_{1}\})\log\left(\frac{3eTB_{1}}{\delta_{1}}\right)+S_{1}\log\left(CMH^{2}(\max\{D,W_{1},N+1\})^{3}\right)$
$\lesssim$ $S_{1}M^{2}\log(\max\{M,H,W_{1}\})\log\left(\frac{TB_{1}}{\delta_{1}}\right).$
Consequently,
$\mathcal{V}_{\infty}(\delta_{1},\mathcal{T},T)\lesssim S_{1}M^{2}\log(\max\{M,H,W_{1}\})\log\left(\frac{TB_{1}}{\delta_{1}}\right).$
Then we establish the upper bound of $\mathcal{V}_{\infty}(\delta_{1},\mathcal{T},T)$ for the special case $K=1$.
For arbitrary $K\geq 1$, note that the covering number of $\mathcal{T}$ under the $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$-norm can be bounded by combining the covering numbers of the $K$ scalar-valued function classes. Thus, we obtain
$\mathcal{V}_{\infty}(\delta_{1},\mathcal{T},T)\lesssim KS_{1}M^{2}\log(\max\{M,H,W_{1}\})\log\left(\frac{TB_{1}}{\delta_{1}}\right).$
This complete the proof of Lemma 13.
In the end of this subsection, we present some lemmas that are used in the proof of Lemma  13.
Lemma 34 (Theorem 8.3 of Anthony and Bartlett (2009))
Suppose $W\leq m$, and consider polynomials $f_{1},\ldots,f_{m}$ of degree at most $D$ in $W$ variables. Define the quantity
$K:=\left|\left\{\left(\operatorname{sgn}\left(f_{1}(\boldsymbol{x})\right),\dots,\operatorname{sgn}\left(f_{m}(\boldsymbol{x})\right)\right):\boldsymbol{x}\in\mathbb{R}^{W}\right\}\right|.$
Then the following bound holds:
$K\leq 2\left(\frac{2emD}{W}\right)^{W}.$
Lemma 35 (Lemma 16 of Bartlett et al. (2019))
Suppose that $2^{m}\leq 2^{L}\left(\frac{mR}{w}\right)^{w}$ for some $R\geq 16$ and $m\geq w\geq L\geq 0$. Then
$m\leq L+w\log_{2}\left(2R\log_{2}R\right).$
Lemma 36 (Theorem 12.2 of Anthony and Bartlett (2009))
Let $\mathcal{F}$ be a class of real-valued functions defined on a domain $\mathcal{X}$ with range bounded in the interval $[0,B]$. Let $\epsilon>0$, and suppose that the pseudo-dimension of $\mathcal{F}$ is $d$. Then the $\ell_{\infty}$ covering number satisfies
$\mathcal{N}_{\infty}(\epsilon,\mathcal{F},m)\leq\sum_{i=1}^{d}\binom{m}{i}\left(\frac{B}{\epsilon}\right)^{i},$
which is further bounded by
$\left(\frac{emB}{\epsilon d}\right)^{d}\quad\text{for }m\geq d.$
Appendix F Proof of Lemma 14
Proof [Proof of Lemma 14] For clarity in our proof, we restate the functional form of networks, i.e., any network $\mathbf{f}\in\mathcal{F}_{id}(L,W,S)$ is rewritten as
$\mathbf{f}(\mathbf{x})=\left(W^{(L)}\sigma(\cdot)+b^{(L)}\right)\circ\dots\circ\left(W^{(1)}\mathbf{x}+b^{(1)}\right),$
where $\sigma(\cdot)$ denotes the ReLU activation function. Define the intermediate mappings
$\mathcal{A}_{k}(\mathbf{f})(\mathbf{x})$ $=$ $\sigma\circ\left(W^{(k-1)}\sigma(\cdot)+b^{(k-1)}\right)\circ\dots\circ\left(W^{(1)}\mathbf{x}+b^{(1)}\right)\quad\text{for }k=2,\dots,L,$
$\mathcal{B}_{k}(\mathbf{f})(\mathbf{x})$ $=$ $\left(W^{(L)}\sigma(\cdot)+b^{(L)}\right)\circ\dots\circ\left(W^{(k)}\sigma(\mathbf{x})+b^{(k)}\right)\quad\text{for }k=1,\dots,L.$
For completeness, set $\mathcal{A}_{1}(\mathbf{f})(\mathbf{x})=\mathbf{x}$ and $\mathcal{B}_{L+1}(\mathbf{f})(\mathbf{x})=\mathbf{x}$. Then we can express
$\mathbf{f}(\mathbf{x})=\mathcal{B}_{k+1}(\mathbf{f})\circ(W^{(k)}\cdot+b^{(k)})\circ\mathcal{A}_{k}(\mathbf{f})(\mathbf{x}),\quad\forall\ k=1,\dots,L.$
Let $\mathbf{f},\mathbf{g}\in\mathcal{F}_{id}(L,W,S)$ be two networks given by
$\mathbf{f}(\mathbf{x})$ $=$ $\left(W^{(L)}\sigma(\cdot)+b^{(L)}\right)\circ\dots\circ\left(W^{(1)}\mathbf{x}+b^{(1)}\right)$
$\mathbf{g}(\mathbf{x})$ $=$ $\left(W^{{}^{\prime}(L)}\sigma(\cdot)+b^{{}^{\prime}(L)}\right)\circ\dots\circ\left(W^{{}^{\prime}(1)}\mathbf{x}+b^{{}^{\prime}(1)}\right)$
with parameters satisfying $\|W^{(k)}-{W^{{}^{\prime}(k)}}\|_{\infty}\leq\delta_{2}$ and $\|b^{(k)}-{b^{{}^{\prime}(k)}}\|_{\infty}\leq\delta_{2}$ for all $1\leq k\leq L$. Note that
$\mathcal{B}_{k+1}(\mathbf{g})\circ(W^{(k)}\cdot+b^{(k)})\circ\mathcal{A}_{k}(\mathbf{f})(\mathbf{x})=\mathcal{B}_{k+2}(\mathbf{g})\circ({W^{{}^{\prime}(k+1)}}\cdot+{b^{{}^{\prime}(k+1)}})\circ\mathcal{A}_{k}(\mathbf{f})(\mathbf{x}),$
for $k=1,\dots,L-1$. This yields
$\|\mathbf{f}(\mathbf{x})-\mathbf{g}(\mathbf{x})\|_{\infty}$
$=$ $\left\|\sum_{k=1}^{L}\mathcal{B}_{k+1}(\mathbf{g})\circ(W^{(k)}\cdot+b^{(k)})\circ\mathcal{A}_{k}(\mathbf{f})(\mathbf{x})-\mathcal{B}_{k+1}(\mathbf{g})\circ({W^{{}^{\prime}(k)}}\cdot+{b^{{}^{\prime}(k)}})\circ\mathcal{A}_{k}(\mathbf{f})(\mathbf{x})\right\|_{\infty}$
$\leq$ $\sum_{k=1}^{L}\|\mathcal{B}_{k+1}(\mathbf{g})\|_{Lip}\|(W^{(k)}\cdot+b^{(k)})\circ\mathcal{A}_{k}(\mathbf{f})(\mathbf{x})-({W^{{}^{\prime}(k)}}\cdot+{b^{{}^{\prime}(k)}})\circ\mathcal{A}_{k}(\mathbf{f})(\mathbf{x})\|_{\infty}.$
Next, we explicitly derive the following uniform bound for each intermediate mapping. According the definition of $\mathcal{F}_{\rm id}(L,W,S)$ given in (5), we have
$\|\mathcal{A}_{k}(\mathbf{f})\|_{\infty}$ $\leq$ $\max_{j}\|W^{(k-1)}_{j,:}\|_{1}\|\mathcal{A}_{k-1}(\mathbf{f})\|_{\infty}+\|b^{(k-1)}\|_{\infty}$
$\leq$ $W\|\mathcal{A}_{k-1}(\mathbf{f})\|_{\infty}+1,$
where $W^{(k-1)}_{j,:}$ is the $j$-row of the matrix $W^{(k-1)}$. Consequently,
$\|\mathcal{A}_{k}(\mathbf{f})\|_{\infty}\leq(W+1)^{k-1}.$
Similarly, we can show that the Lipschitz constant of $\mathcal{B}_{k+1}(\mathbf{g})$ is bounded by
$\|\mathcal{B}_{k+1}(\mathbf{g})\|_{Lip}\leq W^{L-k}.$
Using these bounds, we obtain
$\|\mathbf{f}(\mathbf{y})-\mathbf{g}(\mathbf{y})\|_{\infty}\leq\sum_{k=1}^{L}W^{L-k}[W(W+1)^{k-1}+1]\delta_{2}\leq L(W+1)^{L}\delta_{2}.$
Similar to the arguments for Lemma  13, we fix the positions of these nonzero parameters of the network $F_{id}(L,W,S)$ and then its covering number with respect to the $\|\cdot\|_{\infty}$-norm is bounded by
$\left(\frac{3\delta_{2}}{L(W+1)^{L}}\right)^{-S}.$
Recall that $\mathcal{F}_{id}(L,W,S)$ depend on at most $(W+1)^{L}$ parameters, of which at most $S$ parameters can be nonzero. Thus, the number of different chosen patterns of these nonzero parameters is $\binom{(W+1)^{L}}{S}\leq(W+1)^{LS}$. Therefore, we obtain
$\mathcal{N}(\delta_{2},\mathcal{F}_{id}(L,W,S),\|\cdot\|_{\infty})\leq(W+1)^{LS}\left(\frac{3\delta_{2}}{L(W+1)^{L}}\right)^{-S}=\left(\frac{L(W+1)^{2L}}{3\delta_{2}}\right)^{S}.$
Taking logarithms, we have
$\log\mathcal{N}(\delta_{2},\mathcal{F}_{id}(L,W,S),\|\cdot\|_{\infty})\leq 2LS\log\left(\frac{L(W+1)}{3\delta_{2}}\right).$
This completes the proof of Lemma 14.
Appendix G Proof of Lemma 16
The Kullback–Leibler divergence and the Hellinger distance are two classical measures of discrepancy between probability distributions; see Tsybakov (2008, Section 2.4) and Birgé and Massart (1998, Lemma 5). These two measures are closely related and satisfy several useful inequalities, which are fundamental in nonparametric estimation theory. Recall that the squared Hellinger distance between $P$ and $Q$ is defined as
$H^{2}(P,Q)=\left(1-\int\sqrt{dP\,dQ}\right),$
as given in Birgé and Massart (1998, Section 7.3).
We proceed to establish the inequality stated in Lemma 16.
Proof [Proof of Lemma 16] We first show that $H^{2}(P,Q)\leq\frac{1}{2}\mathrm{KL}_{2}(P,Q)\leq\frac{1}{2}\mathrm{KL}_{B}(P,Q)$. As the second inequality follows from the definition of the truncated KL-divergence, we only show $H^{2}(P,Q)\leq\frac{1}{2}\mathrm{KL}_{2}(P,Q)$. Let $P=P^{a}+P^{s}$ be the Lebesgue decomposition of $P$ with respect to $Q$, where $P^{a}\ll Q$ and $P^{s}\perp Q$. Let $\mu=\frac{1}{2}(P+Q)$ be the dominating measure and let $p,p^{a},p^{s},q$ denote the corresponding densities of $P,P^{a},P^{s},Q$ with respect to $\mu$, respectively. Since $p^{s}q=0$, it follows that
$H^{2}(P,Q)=1-\int\sqrt{pq}d\mu=\int(p^{a}+p^{s}-\sqrt{p^{a}q})\,d\mu.$
Since $x\leq-\log(1+x)$ for x>1x>-1, it follows that
$\mathrm{KL}_{2}(P,Q)$ $=$ $\int\left(2\wedge\log\frac{dP}{dQ}\right)dP$
$=$ $\int\left(2\wedge 2\log\sqrt{p/q}\right)pd\mu$
$=$ $2\int\left[1\wedge\left(-\log\left(1+\sqrt{p/q}-1\right)\right)\right]pd\mu$
$\geq$ $2\int\left[1\wedge\left(1-\sqrt{q/p}\right)\right]pd\mu$
$=$ $2\int(p-\sqrt{pq})d\mu$
$=$ $2H^{2}(P,Q).$
Thus, we obtain $H^{2}(P,Q)\leq\frac{1}{2}KL_{2}(P,Q)$.
The proof of the upper bound on $\mathrm{KL}_{B}(P,Q)$ in terms of $H^{2}(P,Q)$ follows a similar approach to that of Theorem 5 of Wong and Shen (1995) and Lemma 3.4 of Bos and Schmidt-Hieber (2022). For completeness, we also provide a detailed derivation. Note that by second order Taylor expansion, we have $x\log(x)\leq x-1+\frac{1}{2}(x-1)^{2}/(x\wedge 1)$ for $x>0$. Consequently,
$x\log x=2\sqrt{x}[\sqrt{x}\log(\sqrt{x})]\leq 2(x-\sqrt{x})+(\sqrt{x}\vee 1)(\sqrt{x}-1)^{2},\quad\forall\ x>0.$
Applying this inequality to $x=p^{a}/q$, we obtain that on the set $\{p^{a}/q\leq e^{B}\}$,
$\frac{p^{a}}{q}\log\left(\frac{p^{a}}{q}\right)\leq 2\left(\frac{p^{a}}{q}-\sqrt{\frac{p^{a}}{q}}\right)+e^{B/2}\left(\sqrt{\frac{p^{a}}{q}}-1\right)^{2}.$
This implies
$\mathrm{KL}_{B}(P,Q)$ $=$ $\int_{p^{a}/q\leq e^{B}}\frac{p^{a}}{q}\log\left(\frac{p^{a}}{q}\right)qd\mu+B\int_{\frac{dP}{dQ}>e^{B}}dP$
$\leq$ $2\int_{p^{a}/q\leq e^{B}}(p^{a}-\sqrt{p^{a}q})d\mu+e^{B/2}\int_{p^{a}/q\leq e^{B}}(\sqrt{p^{a}}-\sqrt{q})^{2}d\mu+B\int_{\frac{dP}{dQ}>e^{B}}dP$
$\leq$ $2\int_{p^{a}/q\leq e^{B}}(p^{a}-\sqrt{p^{a}q})d\mu+2e^{B/2}H^{2}(P,Q)+B\int_{\frac{dP}{dQ}>e^{B}}dP,$
where the last inequality is due to $p^{s}q=0$ and $H^{2}(P,Q)=\int(1-\sqrt{pq})d\mu$. Next, we consider two cases to complete the proof.
Case 1: $\int_{p^{a}/q\leq e^{B}}(p^{a}-\sqrt{p^{a}q})d\mu\leq 0$. Note that
$H^{2}(P,Q)=\frac{1}{2}\int(\sqrt{p}-\sqrt{q})^{2}d\mu\geq\frac{1}{2}\int_{p/q\geq e^{B}}p(1-e^{-B/2})^{2}d\mu=\frac{1}{2}(1-e^{-B/2})^{2}\int_{\frac{dP}{dQ}>e^{B}}dP.$
Consequently,
$\mathrm{KL}_{B}(P,Q)\leq 2\left(e^{B/2}+B(1-e^{-B/2})^{-2}\right)H^{2}(P,Q).$
Case 2: $\int_{p^{a}/q\leq e^{B}}(p^{a}-\sqrt{p^{a}q})d\mu>0$. Some elementary calculations show that
$B\int_{\frac{dP}{dQ}>e^{B}}dP\leq B(1-e^{-B/2})^{-1}\int_{p^{a}/q>e^{B}}(p-\sqrt{pq})d\mu.$
It follows that
$\mathrm{KL}_{B}(P,Q)\leq$ $\ 2\int_{p^{a}/q\leq e^{B}}(p^{a}-\sqrt{p^{a}q})d\mu+2e^{B/2}H^{2}(P,Q)+B\int_{\frac{dP}{dQ}>e^{B}}dP$
$\leq$ $\ 2e^{B/2}H^{2}(P,Q)+B(1-e^{-B/2})^{-1}\int_{p^{a}/q\leq e^{B}}(p-\sqrt{pq})d\mu$
$\ +B(1-e^{-B/2})^{-1}\int_{p^{a}/q>e^{B}}(p-\sqrt{pq})d\mu$
$=$ $\ 2e^{B/2}H^{2}(P,Q)+B(1-e^{-B/2})^{-1}\int(p-\sqrt{pq})d\mu$
$=$ $\ \left(2e^{B/2}+B(1-e^{-B/2})^{-1}\right)H^{2}(P,Q).$
In both cases, since $B\geq 2$, it follows that
$B(1-e^{-B/2})^{-1}\leq 2e^{B/2},\quad B(1-e^{-B/2})^{-2}\leq 4e^{B/2}.$
Thus, we obtain $\mathrm{KL}_{B}(P,Q)\leq 10e^{B/2}H^{2}(P,Q)$. Consequently,
$\frac{1}{2}\mathrm{KL}_{B}(P,Q)\leq 5e^{B/2}H^{2}(P,Q).$
This completes the proof of Lemma 16.
Appendix H Proof of Lemma 19
Proof [Proof of Lemma 19] Let $\mathcal{F}_{\mathrm{MLP}}(L,W,S)$ be defined as in (9), where $\boldsymbol{\Phi}$ denotes the softmax mapping. For any empirical input set $\mathcal{Z}_{T}\subset[0,1]^{d}$, we aim to bound the covering number of the log-transformed class $\log\big{(}\mathcal{F}_{\mathrm{MLP}}(L,W,S)\big{)}$ under the empirical sup-norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$.
By Lemma 31, for any $\delta>0$, we have
$\mathcal{N}\left(\delta,\log\left(\mathcal{F}_{\mathrm{MLP}}(L,W,S)\right),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\mathcal{N}\left(\tfrac{\delta}{2K},\mathcal{F}_{\mathrm{id}}(L,W,S),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right).$
Setting $\delta=\tfrac{1}{T}$, it follows that
$\mathcal{N}\left(\tfrac{1}{T},\log\left(\mathcal{F}_{\mathrm{MLP}}(L,W,S)\right),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\mathcal{N}\left(\tfrac{1}{2KT},\mathcal{F}_{\mathrm{id}}(L,W,S),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right).$
Applying the standard pseudo-dimension-based covering bound (see Lemma 36, Theorem 12.2 of Anthony and Bartlett (2009)), we obtain
$\mathcal{N}\left(\tfrac{1}{2KT},\mathcal{F}_{\mathrm{id}}(L,W,S),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\left(\frac{2eKC\log T\cdot T^{2}}{\operatorname{Pdim}(\mathcal{F}_{\mathrm{id}})}\right)^{\operatorname{Pdim}(\mathcal{F}_{\mathrm{id}})},$
where the output range of $\mathcal{F}_{\mathrm{id}}$ is bounded by $C\log T$ for some universal constant $C>0$.
Furthermore, Theorems 3 and 6 of Bartlett et al. (2019) imply that there exist universal constants $\underline{c},\overline{c}>0$ such that
$\underline{c}\cdot SL\log(S/L)\leq\operatorname{Pdim}(\mathcal{F}_{\mathrm{id}}(L,W,S))\leq\overline{c}\cdot SL\log S.$
Taking the logarithm and the supremum over all $\mathcal{Z}_{T}\subset[0,1]^{d}$, we conclude that
$\mathcal{V}_{\infty}\left(\tfrac{1}{T},\log\left(\mathcal{F}_{\mathrm{MLP}}(L,W,S)\right),T\right)\leq C\cdot SL\log S\cdot\log T,$
for some universal constant $C>0$.

### v2

@@
 Abstract
-This paper investigates the expected excess risk of In-Context Learning (ICL) for multi-class classification.
-We model each task as a sequence of labeled prompt samples and a query input, where a pre-trained model estimates the conditional class probabilities of the query.
-The expected excess risk is defined as the average truncated Kullback–Leibler (KL) divergence between the predicted and ground-truth conditional class distributions, averaged over a specified family of tasks.
-We establish a new oracle inequality for the expected excess risk based on KL divergence in multiclass classification.
-This allows us to derive tight upper and lower bounds for the expected excess risk in transformer-based models, demonstrating that the ICL estimator achieves the minimax optimal rate—up to a logarithmic factor—for conditional probability estimation.
-From a technical standpoint, our results introduce a novel method for controlling generalization error using the uniform empirical covering entropy of the log-likelihood function class.
-Furthermore, we show that multilayer perceptrons (MLPs) can also perform ICL and achieve this optimal rate under specific assumptions, suggesting that transformers may not be the exclusive architecture capable of effective ICL.
-Keywords: expected excess risk, in-context learning, multi-class classification, multilayer perceptrons, transformers
+This paper investigates the expected excess risk of in-context learning (ICL) for multiclass classification.
+We formalize each task as a sequence of labeled examples followed by a query input; a pretrained model then estimates the query’s conditional class probabilities.
+The expected excess risk is defined as the average truncated Kullback-Leibler (KL) divergence between the predicted and true conditional class distributions over a specified family of tasks.
+We establish a new oracle inequality for this risk, based on KL divergence, in multiclass classification.
+This yields tight upper and lower bounds for transformer-based models, showing that the ICL estimator achieves the minimax optimal rate (up to logarithmic factors) for conditional probability estimation.
+From a technical standpoint, our results introduce a novel method for controlling generalization error via uniform empirical entropy.
+We further demonstrate that multilayer perceptrons (MLPs) can also perform ICL and attain the same optimal rate (up to logarithmic factors) under suitable assumptions, suggesting that effective ICL need not be exclusive to transformer architectures.
+Keywords: expected excess risk, in-context learning, multiclass classification, multilayer perceptrons, transformers
 1 Introduction
@@
 Brown et al.
-(2020) observed that when a model reaches a certain scale, it can emerge the remarkable capability of in-context learning (ICL).
-Specifically, pre-trained LLMs can perform novel tasks with just a few labeled examples in the prompt, without any parameter updates or fine-tuning.
-This capability facilitates exceptional few-shot generalization in LLMs while establishing a novel paradigm for few-shot learning (Brown et al., 2020; Garcia et al., 2023).
-Consequently, it accelerates the deployment of deep learning models, enabling a foundational shift toward LLMs as instruction-following systems.
+(2020) observed that, once scaled sufficiently, LLMs exhibit the remarkable capability of in-context learning (ICL).
+Specifically, pretrained LLMs can perform novel tasks with just a few labeled examples in the prompt, without any parameter updates or fine-tuning.
+This capability enables strong few-shot generalization and has established a new paradigm for few-shot learning (Brown et al., 2020; Garcia et al., 2023).
+Consequently, this paradigm accelerates the deployment of deep learning models, enabling a foundational shift toward LLMs as instruction-following systems.
 A recent line of research has quantified ICL by studying transformer models pretrained on linear regression tasks with Gaussian priors (Garg et al., 2022; Akyürek et al., 2023; Li et al., 2023; Raventós et al., 2023; Wu et al., 2024).
-Concurrently, a growing body of research has revealed a close connection between the forward pass of multi-layer transformer models and multi-step gradient descent algorithms (Akyürek et al., 2023; Von Oswald et al., 2023; Bai et al., 2023; Ahn et al., 2023; Zhang et al., 2024), suggesting that transformers may implement ICL by implicitly simulating gradient descent.
-Recent empirical findings show that transformers fail to exhibit ICL unless they are pretrained on a sufficiently diverse set of tasks (Raventós et al., 2023), highlighting the crucial role of task diversity in pretraining.
+Concurrently, a growing body of work has revealed a close connection between the forward pass of multi-layer transformer models and multi-step gradient descent algorithms (Akyürek et al., 2023; Von Oswald et al., 2023; Bai et al., 2023; Ahn et al., 2023; Zhang et al., 2024), suggesting that transformers may implement ICL by implicitly simulating gradient descent.
+Recent empirical findings indicate that transformers fail to exhibit ICL unless they are pretrained on a sufficiently diverse set of tasks (Raventós et al., 2023), underscoring the crucial role of task diversity in pretraining.
 The pioneering experimental study by Garg et al.
-(2022) demonstrated that transformers can in-context learn a variety of fundamental function classes, thereby sparking theoretical analyses of ICL in statistics.
+(2022) demonstrated that transformer models can in-context learn a variety of fundamental function classes, thereby motivating theoretical analyses of ICL in statistics.
 Wu et al.
@@
 Kim et al.
-(2024) considered deep neural network with one-layer linear transformer structure and established the minimax optimality of the in-context learner from the view of nonparametric regression.
-Although substantial research has analyzed the ICL capabilities of large language models, achieving significant progress in understanding this mechanism, there are still notable limitations and challenges in its theoretical analysis.
-First, a significant gap exists between the limit assumptions underlying theoretical analyses and the scenarios typically encountered in practical applications.
-Most existing studies focus on the combined regime of task scaling and context scaling (Abedsoltan et al., 2024), where both the number of pretraining tasks and the context length are assumed to be large.
-Some analyses further consider the asymptotic limit as both the task scaling and the context scaling approach infinity.
+(2024) analyzed a model consisting of a deep neural network and a single-layer linear transformer, and established the minimax optimality of the in-context learner from the perspective of nonparametric regression.
+Although substantial research has analyzed the ICL capabilities of large language models and achieved significant progress in understanding this mechanism, notable limitations and challenges remain in its theoretical analysis.
+First, a significant gap exists between the assumptions of theoretical frameworks and the constraints of practical applications.
+As noted by Abedsoltan et al.
+(2024), there are two different aspects of generalization in ICL: task-scaling, where model performance improves as the number of pretraining tasks $T$ increases while keeping the context length $N$ fixed, and context-scaling, where model performance improves as $N$ increases while $T$ remains fixed.
+Most existing studies focus on the combined regime of task-scaling and context-scaling, where both $T$ and $N$ are assumed to be large.
+Some analyses further consider the asymptotic limit as both $T$ and $N$ approach infinity.
 However, the practical importance of few-shot scenarios is neglected, as inference is typically constrained by limited context length.
-For example, GPT-3 was evaluated under few-shot settings  (Brown et al., 2020), where the context length only allowed one or two labeled examples per class in the prompt.
-To our best knowledge, Wu et al.
-(2024) rigorously analyzed the impact of task scaling on the prediction risk of ICL at a fixed context length.
-Second, there are substantial limitations in both task settings and model architecture, as current theoretical results do not cover multi-classification tasks or complete transformer architecture.
-Most theoretical studies on ICL concentrate on simplified transformer architectures, such as linear transformers or single-head attention, to facilitate analytical tractability (Zhang et al., 2024; Huang et al., 2024; Kim and Suzuki, 2024; Li et al., 2024a).
-However, empirical research demonstrates that in a wide range of ICL tasks, multi-head dot-product attention mechanisms offers substantial performance gains over its single-head counterpart (Cui et al., 2024; Xing et al., 2024).
-For regression problems, some progresses have been made in understanding ICL based on multi-head dot-product attention (Chen et al., 2024a, b; Deora et al., 2024).
-Li et al.
-(2024b) and Li et al.
-(2024c) considered the binary classification and analyzed the training dynamics of single-layer transformer models.
-Frei and Vardi (2024) investigated in-context classifcation of Gaussian mixtures using a linear transformer model.
-Meanwhile, multiclass classification, which is fundamental and widely used in deep learning, has not received adequate theoretical attention.
+For example, GPT-3 was evaluated under few-shot settings (Brown et al., 2020), where the context length $N$ only allowed one or two labeled examples per class in the prompt.
+To our knowledge, Wu et al.
+(2024) provide the first theoretical analysis of the effect of task-scaling on the prediction risk of ICL under a fixed context length.
+Second, existing theoretical analyses are constrained by simplified task settings and model architectures.
+The literature predominantly focuses on linear regression or binary classification, neglecting the widespread practical use of multiclass classification.
+Architecturally, to maintain analytical tractability, most studies resort to simplified transformer variants, such as linear transformers or single-head attention mechanisms (Zhang et al., 2024; Huang et al., 2024; Kim and Suzuki, 2024; Li et al., 2024a).
+This stands in stark contrast to empirical evidence, which consistently demonstrates the superior performance of standard multi-head dot-product attention in a wide range of ICL tasks (Cui et al., 2024; Xing et al., 2024).
+Although some recent work has begun to address multi-head attention for regression problems (Chen et al., 2024a, b; Deora et al., 2024) and binary classification  (Li et al., 2024b, c), the theoretical understanding of multiclass classification remains underdeveloped.
 Shen et al.
-(2024) studied the training dynamics of a single-layer linear transformer model for multi-classification of Gaussian mixtures in the context of ICL.
-To address these limitations, we focus on multi-classification tasks and develop a theoretical framework for bounding the expected excess risk of ICL under the task-scaling regime.
-Specifically, we derive non-asymptotic error bounds under the assumption that the number of pretraining tasks is much larger than the context length.
-We further analyze asymptotic behavior when the number of pretraining tasks tends to infinite while the context length remains fixed.
-From the perspective of modeling, we frame ICL as a nonparametric conditional probability estimation problem, where the model estimates conditional class probabilities of a query input given labeled prompt examples.
-We use a deep transformer model with multi-head dot-product attention and analyze its generalization performance in ICL.
+(2024) studied multiclass classification of Gaussian mixtures in the context of ICL, but their framework was still limited to single-layer linear transformer models.
+To address these limitations, we develop a theoretical framework for multiclass in-context learning that operates under a task-scaling regime with a deep transformer architecture.
+Specifically, we derive non-asymptotic error bounds on the expected excess risk under the assumption that the number of pretraining tasks $T$ is much larger than the context length $N$.
+We further analyze the asymptotic behavior as the number of pretraining tasks $T$ tends to infinity while the context length $N$ remains fixed.
+From a modeling perspective, we frame ICL as a nonparametric conditional probability estimation problem, where the model estimates conditional class probabilities of a query input given labeled prompt examples.
+To this end, we use a deep transformer model with multi-head dot-product attention and analyze its generalization performance in ICL.
 Our contributions are summarized as follows:
   1. 1.
-    Our work extends existing theoretical frameworks by relaxing main assumptions in experimental and practical settings.
-Specifically, we relax constraints commonly adopted in previous studies, such as focusing solely on regression task, using simplified transformer architecture, or assuming asymptotic regimes for context scaling.
-This enables closer alignment with empirical deployments while preserving analytical rigor.
-Our work bridges critical gaps between theoretical analysis and real-world implementation.
+    Our work extends existing theoretical frameworks by relaxing restrictive assumptions that misalign with practical usage.
+Unlike prior analyses where the context length $N$ is assumed to be large, we focus on few-shot settings which are typically operated in real-world deployments.
+We also relax other constraints common in previous studies, such as focusing on simplified transformer architectures.
+These generalizations preserve analytical rigor while enabling closer alignment with real-world deployments, thereby bridging critical gaps between theoretical analysis and practical implementation.
   2. 2.
     We establish a new oracle inequality for the expected excess risk based on Kullback–Leibler (KL) divergence in multiclass classification.
-To the best of our knowledge, this is the first result that decomposes the expected excess risk under KL divergence into approximation, optimization, and generalization errors, where the generalization error is controlled by the uniform empirical covering entropy of the log-likelihood function class.
+To the best of our knowledge, this is the first result that decomposes the expected excess risk under KL divergence into approximation, optimization, and generalization errors, with the generalization error controlled by uniform empirical entropy.
+While prior work such as Schmidt-Hieber (2020) or Bos and Schmidt-Hieber (2022) establishes oracle inequalities for deep neural networks via $L^{\infty}$-metric entropy over the entire covariate space.
+Our oracle inequality provides a new methodology for risk analysis and can also be applied to other settings of multiclass classification.
   3. 3.
-    Our analysis finds that transformers are not necessarily the uniquely optimal architecture for ICL.
-By comparing the convergence rates of transformer and MLP architectures, we demonstrate that both can achieve minimax-optimal rates under certain conditions.
-This reveals that transformers are not the only viable architecture foundation for ICL, which is also corroborated by recent empirical studies (Tong and Pehlevan, 2025; Kratsios and Furuya, 2025).
+    Our risk analysis finds that transformers are not necessarily the uniquely optimal architecture for ICL.
+By comparing the convergence rates of transformer and MLP architectures, we demonstrate that both can achieve minimax-optimal rates (up to logarithmic factors) under certain conditions.
+This reveals that transformers may be not the only viable architectural foundation for ICL, which is also corroborated by recent empirical studies (Tong and Pehlevan, 2025; Kratsios and Furuya, 2025).
 2 Related Work
-We briefly review related work on ICL and theoretical analyses of transformer architectures across various settings.
+We provide a brief review of related work on ICL and theoretical analyses of transformer architectures in various settings.
 2.1 Theoretical Analyses of Transformers in ICL
 Recent theoretical studies have rigorously analyzed various aspects of transformer-based models in ICL.
-For linear function classes, the statistical complexity and representation power of transformers have been precisely characterized (Zhang et al., 2024; Wu et al., 2024; Cheng et al., 2024), with further extensions to mixture models and broader function classes explored in (Jin et al., 2024; Guo et al., 2024; Collins et al., 2024).
+For linear function classes, the statistical complexity and representational power of transformers have been precisely characterized (Zhang et al., 2024; Wu et al., 2024; Cheng et al., 2024), with extensions to mixture models and broader function classes explored in Jin et al.
+(2024); Guo et al.
+(2024); Collins et al.
+(2024).
 In terms of optimization, convergence properties and landscape analyses of transformer-based ICL objectives have been investigated primarily under simplified settings, focusing on linear attention and regression tasks (Zhang et al., 2024; Huang et al., 2024; Kim et al., 2024; Nichani et al., 2024; Kim and Suzuki, 2024).
 More recently, Chen et al.
-(2024a) have studied single-layer transformers with multi-head attention in nonlinear regimes, although their analysis remains restricted to shallow architectures and regression.
+(2024a) studied single-layer transformers with multi-head attention in nonlinear regimes, although their analysis is restricted to shallow architectures and regression.
 In contrast, we establish expected excess risk bounds for deep transformers with multi-head attention in the multiclass classification setting, thereby substantially extending existing theoretical results.
 2.2 Mechanistic and Algorithmic Perspectives of Transformers in ICL
-Recent work has revealed that transformers can emulate classical learning algorithms such as gradient descent and its variants (Akyürek et al., 2023; Von Oswald et al., 2023; Bai et al., 2023; Fu et al., 2024).
-Studies have also examined global minimizers and critical points that emerge from adaptive gradient descent methods within transformer architectures (Ahn et al., 2023; Mahankali et al., 2024).
-Additionally, Bayesian analyses suggest that transformers implicitly approximate Bayesian inference, further enhancing the interpretability of their ICL mechanisms (Xie et al., 2022; Müller et al., 2022; Zhang et al., 2022; Panwar et al., 2024; Jeon et al., 2024).
+Recent work has shown that transformers can emulate standard learning algorithms such as gradient descent and its variants (Akyürek et al., 2023; Von Oswald et al., 2023; Bai et al., 2023; Fu et al., 2024).
+Additional studies have examined global minimizers and critical points arising in adaptive gradient descent methods within transformer architectures (Ahn et al., 2023; Mahankali et al., 2024).
+Bayesian analyses further suggest that transformers implicitly approximate Bayesian inference, offering a probabilistic explanation for ICL mechanisms (Xie et al., 2022; Müller et al., 2022; Zhang et al., 2022; Panwar et al., 2024; Jeon et al., 2024).
 Transformers’ ability to perform ICL on data generated by Markov chains has also been investigated (Collins et al., 2024; Edelman et al., 2024; Makkuva et al., 2024), as has their potential for in-context decision-making within reinforcement learning frameworks (Lin et al., 2024; Sinii et al., 2024).
-Broader theoretical perspectives concerning the generalization behavior, algorithmic interpretations, and meta-learning properties of transformer-based models have been extensively studied by Li et al.
+Broader theoretical perspectives on the generalization behavior, algorithmic interpretations, and meta-learning properties of transformer-based models have been extensively studied by Li et al.
 (2023), Dai et al.
@@
 Analyses of multi-head attention mechanisms, from both representational and Bayesian viewpoints, have further illuminated the learning dynamics in these models (An et al., 2020; Mahdavi et al., 2024).
-A detailed discussion of these results, however, lies beyond the scope of this paper.
+A detailed discussion of these results lies beyond the scope of this paper.
 2.3 ICL Beyond Transformers
-Recent literature has begun to investigate whether models beyond transformers can exhibit in-context learning behavior.
-Tong and Pehlevan (2025) empirically demonstrated that multi-layer perceptrons (MLPs) can achieve ICL performance comparable to transformers on synthetic tasks.
-Complementing this, Kratsios and Furuya (2025) proved that MLPs possess universal approximation capabilities in an ICL setting.
-Our theoretical results further support these empirical and theoretical findings by showing that, under comparable conditions, MLPs attain minimax-optimal rates for multiclass classification.
-This provides evidence that attention-based architectures may not be essential for ICL.
+Recent literature has begun to investigate whether architectures beyond transformers can exhibit in-context learning behavior.
+Tong and Pehlevan (2025) empirically demonstrated that multilayer perceptrons (MLPs) can match the ICL performance of transformers on synthetic tasks.
+Complementing this, Kratsios and Furuya (2025) showed that MLPs are universal approximators in an ICL setting.
+Our theoretical results further support these findings by demonstrating that, under appropriate conditions, MLPs attain minimax optimal rates for multiclass classification.
+These results provide evidence that attention-based architectures may not be strictly necessary for ICL.
 3 Preliminaries
@@
 The set of natural numbers including zero is denoted by $\mathbb{N}:=\mathbb{Z}_{\geq 0}$.
-For a vector $\mathbf{x}=(x_{1},\cdots,x_{p})^{\top}\in\mathbb{R}^{p}$ and a matrix $\mathbf{X}=(X_{ij})\in\mathbb{R}^{p\times N}$, we define the entrywise max-norms $\|\mathbf{x}\|_{\infty}:=\max_{i}|x_{i}|$ and $\|\mathbf{X}\|_{\infty}:=\max_{i,j}|X_{ij}|$, the Euclidean norm $\|\mathbf{x}\|:=(\sum_{i}x_{i}^{2})^{1/2}$, and the Frobenius norm $\|\mathbf{X}\|_{F}:=(\sum_{i,j}X_{ij}^{2})^{1/2}$.
+For a vector $\mathbf{x}=(x_{1},\cdots,x_{p})^{\top}\in\mathbb{R}^{p}$ and a matrix $\mathbf{X}=(X_{ij})\in\mathbb{R}^{p\times N}$, we define the entrywise max-norms $\|\mathbf{x}\|_{\infty}:=\max_{i}|x_{i}|$, $\|\mathbf{X}\|_{\infty}:=\max_{i,j}|X_{ij}|$, the Euclidean norm $\|\mathbf{x}\|:=(\sum_{i}x_{i}^{2})^{1/2}$, and the Frobenius norm $\|\mathbf{X}\|_{F}:=(\sum_{i,j}X_{ij}^{2})^{1/2}$.
 The counting norm $\|\mathbf{x}\|_{0}$ denotes the number of nonzero entries in $\mathbf{x}$, and analogously for matrices.
-For vector-valued functions $\mathbf{f},\mathbf{g}:\Omega\subset\mathbb{R}^{d}\to\mathbb{R}^{K}$, we write $\|\mathbf{f}-\mathbf{g}\|_{\infty}:=\sup_{\mathbf{z}\in\Omega}\max_{j\in[K]}|f_{j}(\mathbf{z})-g_{j}(\mathbf{z})|$, omitting $\Omega$ when clear.
+For a vector-valued function $\mathbf{f}=(f_{1},\dots,f_{K})^{\top}:\Omega\subset\mathbb{R}^{d}\to\mathbb{R}^{K}$, we define the $L^{\infty}$-norm as $\|\mathbf{f}\|_{\infty}:=\sup_{\mathbf{z}\in\Omega}\max_{j\in[K]}|f_{j}(\mathbf{z})|$, omitting $\Omega$ when clear.
 For vectors $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{K}$, define $\log(\mathbf{x}/\mathbf{x}^{\prime}):=\log(\mathbf{x})-\log(\mathbf{x}^{\prime})$, applied coordinatewise.
@@
 A standard basis vector in $\mathbb{R}^{K}$ has the form $(0,\ldots,0,1,0,\ldots,0)^{\top}$.
-The (K1)(K{-}1)-simplex is $\mathcal{S}^{K}:=\{\mathbf{p}\in\mathbb{R}^{K}:\sum_{k}p_{k}=1,\;p_{k}\geq 0\}$.
+The (K1)(K-1)-simplex is $\mathcal{S}^{K}:=\{\mathbf{p}=(p_{1},\dots,p_{K})^{\top}\in\mathbb{R}^{K}:\sum_{k}p_{k}=1,\;p_{k}\geq 0\}$.
 For probability measures $P$ and $Q$, the Kullback–Leibler divergence is $\mathrm{KL}(P\|Q):=\int\log(dP/dQ)\,dP$ if $P\ll Q$, and $\infty$ otherwise.
-We use the notation $f(x)\lesssim g(x)$ to signify that $f(x)\leq Cg(x)$ for some constant $C>0$.
+We use the notation $f(x)\lesssim g(x)$ to signify that $f(x)\leq Cg(x)$ for all $x$ and some constant $C>0$.
 Similarly, $f(x)\gtrsim g(x)$ signifies $f(x)\geq Cg(x)$.
+For $\mathbf{x}\in\mathbb{R}^{p}$, we write $\|\mathbf{x}\|_{1}=\sum_{i}|x_{i}|$.
 The hardmax operator is $\sigma_{H}(\mathbf{x}):=\lim_{c\to\infty}\exp(c\mathbf{x})/\|\exp(c\mathbf{x})\|_{1}$, applied columnwise for matrices.
 The Hadamard product $\odot$ denotes entrywise multiplication.
-The metric entropy of the function class $\mathcal{F}$ with respect to the metric $\rho$ is defined as $\mathcal{V}(\delta,\mathcal{F},\rho):=\log\mathcal{N}(\delta,\mathcal{F},\rho)$, where $\mathcal{N}$ is the $\delta$-covering number of $\mathcal{F}$ under the metric $\rho$.
-Throughout this paper, we use $C$ to denote a generic constant, which may differ from line to line.
+The metric entropy of the function class $\mathcal{F}$ with respect to the metric $\rho$ is defined as $\mathcal{V}(\delta,\mathcal{F},\rho):=\log\mathcal{N}(\delta,\mathcal{F},\rho)$, where $\mathcal{N}(\delta,\mathcal{F},\rho)$ is the $\delta$-covering number of $\mathcal{F}$ under the metric $\rho$.
+Throughout this paper, we use $C$ to denote a generic constant, which may differ across appearances.
 3.1 In-Context Learning
 In-context learning (ICL) refers to the phenomenon where a model performs prediction by conditioning on a sequence of input-output examples, known as a prompt, without updating its parameters (Brown et al., 2020; Garg et al., 2022).
-A prompt consists of a sequence $(\mathbf{x}_{1},\mathbf{y}_{1},\ldots,\mathbf{x}_{N},\mathbf{y}_{N},\mathbf{x}_{N+1})$, where each $\mathbf{y}_{i}=f(\mathbf{x}_{i})$ for some unknown function $f$, and $\mathbf{x}_{N+1}$ is a query input.
-The goal is to output a prediction $\hat{\mathbf{y}}(\mathbf{x}_{N+1})$ such that $\hat{\mathbf{y}}(\mathbf{x}_{N+1})\approx f(\mathbf{x}_{N+1})$.
-In ICL, each task is specified by a joint distribution $P$ over $(\mathbf{x},\mathbf{y})\in[0,1]^{p}\times[K]$, where $P$ is drawn from a family $\mathcal{P}$.
+A prompt consists of $N$ labeled examples together with an unlabeled query input, i.e., $(\mathbf{x}_{1},\mathbf{y}_{1},\ldots,\mathbf{x}_{N},\mathbf{y}_{N},\mathbf{x}_{N+1})$, where each $\mathbf{y}_{i}=f(\mathbf{x}_{i})$ for some unknown function $f$.
+The goal is to output a prediction $\hat{\mathbf{y}}(\mathbf{x}_{N+1})$ for the query input $\mathbf{x}_{N+1}$ such that $\hat{\mathbf{y}}(\mathbf{x}_{N+1})\approx f(\mathbf{x}_{N+1})$.
+In ICL for multiclass classification, each task is specified by a joint distribution $P$ over $(\mathbf{x},\mathbf{y})\in[0,1]^{p}\times\{0,1\}^{K}$, where $\mathbf{y}$ is a one-hot label vector.
+The distribution $P$ is drawn from a family of probability measures $\mathcal{P}$ and may vary across tasks.
 For a given task, the model observes $N$ i.i.d.
-examples from $P$ as a prompt and is then required to predict the label $\mathbf{y}_{N+1}$ for a query input $\mathbf{x}_{N+1}$, where $(\mathbf{x}_{N+1},\mathbf{y}_{N+1})\sim P$ independently of the prompt.
-For a given task, we draw $(\mathbf{x}_{1},\mathbf{y}_{1}),\ldots,(\mathbf{x}_{N+1},\mathbf{y}_{N+1})\overset{\text{i.i.d.}}{\sim}P$, where the first $N$ pairs form the prompt and $\mathbf{x}_{N+1}$ is the query.
+examples $(\mathbf{x}_{1},\mathbf{y}_{1}),\ldots,(\mathbf{x}_{N},\mathbf{y}_{N})$ from $P$ and is then required to predict the label $\mathbf{y}_{N+1}$ for a query input $\mathbf{x}_{N+1}$, where $(\mathbf{x}_{N+1},\mathbf{y}_{N+1})\sim P$ independently of the first $N$ labeled examples.
 Following Bai et al.
-(2023), we encode the prompt and the query into a matrix $\mathbf{Z}\in\mathbb{R}^{(p+K)\times(N+1)}$ as
-$\mathbf{Z}=\begin{bmatrix}\mathbf{X}&\mathbf{x}_{N+1}\\ \mathbf{Y}&\mathbf{o}_{K}\end{bmatrix}$,
+(2023), we encode the prompt $(\mathbf{x}_{1},\mathbf{y}_{1},\ldots,\mathbf{x}_{N},\mathbf{y}_{N},\mathbf{x}_{N+1})$ into a matrix $\mathbf{Z}\in\mathbb{R}^{(p+K)\times(N+1)}$:
+$\mathbf{Z}=\begin{bmatrix}\mathbf{X}&\mathbf{x}_{N+1}\\ \mathbf{Y}&\mathbf{0}_{K}\end{bmatrix}$,
 where $\mathbf{X}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})\in\mathbb{R}^{p\times N}$, $\mathbf{Y}=(\mathbf{y}_{1},\ldots,\mathbf{y}_{N})\in\mathbb{R}^{K\times N}$, and $\mathbf{o}_{K}\in\mathbb{R}^{K}$ is a zero vector placeholder for the label of the query input $\mathbf{x}_{N+1}$, which the model is expected to predict based on the prompt.
@@
 Consider task $t$ with prompt $\mathbf{Z}^{(t)}=(\mathbf{X}^{(t)},\mathbf{Y}^{(t)},\mathbf{x}_{N+1}^{(t)})$, where $(\mathbf{X}^{(t)},\mathbf{Y}^{(t)})$ denotes the set of $N$ in-context examples and $\mathbf{x}_{N+1}^{(t)}$ is the query input.
-Let $\mathbf{y}_{N+1}^{(t)}\in\{0,1\}^{K}$ be the one-hot encoded label of the query, with $y_{N+1,k}^{(t)}$ denoting its $k$-th component.
+Let $\mathbf{y}_{N+1}^{(t)}\in\{0,1\}^{K}$ be the one-hot label vector of the query, with $y_{N+1,k}^{(t)}$ denoting its $k$-th component.
 The true conditional class probabilities are defined as
@@
 and the full vector of probabilities is given by
-$\mathbf{p}_{0}(\mathbf{Z}^{(t)}):=\left(p_{1}^{0}(\mathbf{Z}^{(t)}),\dots,p_{K}^{0}(\mathbf{Z}^{(t)})\right)\in\mathcal{S}^{K},$
-where $\mathcal{S}^{K}$ denotes the $K$-dimensional probability simplex.
+$\mathbf{p}_{0}(\mathbf{Z}^{(t)}):=\left(p_{1}^{0}(\mathbf{Z}^{(t)}),\dots,p_{K}^{0}(\mathbf{Z}^{(t)})\right)^{\top}\in\mathcal{S}^{K},$
+where $\mathcal{S}^{K}$ denotes the $K$-dimensional probability simplex and $\top$ denotes the transpose.
 To estimate the conditional class probabilities, a standard approach in deep learning is to minimize the empirical negative log-likelihood via (stochastic) gradient descent.
-Let $\mathcal{D}_{T}=\{(\mathbf{Z}^{(t)},\mathbf{y}_{N+1}^{(t)})\}_{t=1}^{T}$ be a dataset of $T$ tasks, where each $\mathbf{Z}^{(t)}$ denotes the prompt input and $\mathbf{y}_{N+1}^{(t)}\in\{0,1\}^{K}$ is a one-hot encoded label vector.
-Given a conditional class probability vector $\mathbf{p}(\mathbf{Z})=(p_{1}(\mathbf{Z}),\dots,p_{K}(\mathbf{Z}))$, the likelihood over $\mathcal{D}_{T}$ is defined as $\mathcal{L}(\mathbf{p}\mid\mathcal{D}_{T})=\prod_{t=1}^{T}\prod_{k=1}^{K}\left(p_{k}(\mathbf{Z}^{(t)})\right)^{y_{N+1,k}^{(t)}}$.
+Let $\mathcal{D}_{T}=\{(\mathbf{Z}^{(t)},\mathbf{y}_{N+1}^{(t)})\}_{t=1}^{T}$ be a data set of $T$ tasks.
+Given a conditional class probability vector $\mathbf{p}(\mathbf{Z})=(p_{1}(\mathbf{Z}),\dots,p_{K}(\mathbf{Z}))^{\top}$, the likelihood over $\mathcal{D}_{T}$ is defined as $\mathcal{L}(\mathbf{p}\mid\mathcal{D}_{T})=\prod_{t=1}^{T}\prod_{k=1}^{K}\left(p_{k}(\mathbf{Z}^{(t)})\right)^{y_{N+1,k}^{(t)}}$.
 Taking the negative logarithm yields the empirical negative log-likelihood (or cross-entropy loss):
 $\ell(\mathbf{p},\mathcal{D}_{T})=-\frac{1}{T}\sum_{t=1}^{T}\sum_{k=1}^{K}y_{N+1,k}^{(t)}\log p_{k}(\mathbf{Z}^{(t)})=-\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}_{N+1}^{(t)\,\top}\log\mathbf{p}(\mathbf{Z}^{(t)}),$
-where the logarithm is applied element-wise and denotes the transpose.
+where the logarithm is applied element-wise.
 This loss measures the average discrepancy between the predicted and true class distributions across tasks.
@@
 The resulting estimator $\hat{\mathbf{p}}\in\mathcal{F}$ is obtained by optimizing over a function class $\mathcal{F}$, typically parameterized by a deep neural network and constrained to produce valid probability vectors.
-We aim to investigate a non-asymptotic bound of the expected excess risk of the estimator $\hat{\mathbf{p}}$, as measured by the expected cross-entropy loss:
-$\mathbb{E}_{\mathcal{D}_{T},(\mathbf{Z},\mathbf{y}_{N+1})}\left[\mathbf{y}_{N+1}^{\top}\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\hat{\mathbf{p}}(\mathbf{Z})}\right]=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[\mathbf{p}_{0}(\mathbf{Z})^{\top}\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\hat{\mathbf{p}}(\mathbf{Z})}\right]=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[{\rm KL}(\mathbf{p}_{0}(\mathbf{Z})\|\hat{\mathbf{p}}(\mathbf{Z}))\right],$
-where $(\mathbf{Z},\mathbf{y}_{N+1})$ is independent of the data set $\mathcal{D}_{T}$, $\mathbf{p}_{0}(\mathbf{Z})$ is the true conditional distribution, and $\hat{\mathbf{p}}(\mathbf{Z})$ is the predicted distribution, and ${\rm KL}(\mathbf{p}_{0}(\mathbf{Z})\|\hat{\mathbf{p}}(\mathbf{Z}))$ denotes the Kullback–Leibler (KL) divergence between the two distributions, conditional on $\mathbf{Z}$ and $(\mathcal{D}_{T},\mathbf{Z})$, respectively.
+We aim to investigate a non-asymptotic bound of the expected excess risk of the estimator $\widehat{\mathbf{p}}$, as measured by the expected cross-entropy loss:
+[…]
+$=$ $\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[\mathbf{p}_{0}(\mathbf{Z})^{\top}\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\widehat{\mathbf{p}}(\mathbf{Z})}\right]$
+$=$ $\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[{\rm KL}(\mathbf{p}_{0}(\mathbf{Z})\|\widehat{\mathbf{p}}(\mathbf{Z}))\right],$
+where $(\mathbf{Z},\mathbf{y}_{N+1})$ is independent of the data set $\mathcal{D}_{T}$, $\mathbf{p}_{0}(\mathbf{Z})$ is the true conditional distribution, and $\widehat{\mathbf{p}}(\mathbf{Z})$ is the predicted distribution, and ${\rm KL}(\mathbf{p}_{0}(\mathbf{Z})\|\widehat{\mathbf{p}}(\mathbf{Z}))$ denotes the Kullback-Leibler (KL) divergence between the two distributions $\mathbf{p}_{0}(\mathbf{Z})$ and $\widehat{\mathbf{p}}(\mathbf{Z})$, conditional on $\mathbf{Z}$ and $(\mathcal{D}_{T},\mathbf{Z})$, respectively.
 However, the expected KL divergence may diverge when the likelihood ratio becomes unbounded (see Lemma 2.1 in Bos and Schmidt-Hieber (2022)).
 To address this limitation, following the suggestion of Bos and Schmidt-Hieber (2022), we consider a truncated version of the KL divergence at level $B>0$:
-$\mathrm{KL}_{B}(\mathbf{p}_{0}(\mathbf{Z})\,\|\,\hat{\mathbf{p}}(\mathbf{Z}))=\mathbf{p}_{0}(\mathbf{Z})^{\top}\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\hat{\mathbf{p}}(\mathbf{Z})}\right),$
-where $a\wedge b=\min\{a,b\}$.
-Therefore, the expected excess risk of the estimator $\hat{\mathbf{p}}$, with respect to the expected cross-entropy loss, is defined as the expected truncated Kullback-Leibler divergence:
-$R_{B}(\mathbf{p}_{0},\hat{\mathbf{p}}):=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[\mathrm{KL}_{B}(\mathbf{p}_{0}(\mathbf{Z})\,\|\,\hat{\mathbf{p}}(\mathbf{Z}))\right]=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[\mathbf{p}_{0}(\mathbf{Z})^{\top}\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\hat{\mathbf{p}}(\mathbf{Z})}\right)\right].$
+$\mathrm{KL}_{B}(\mathbf{p}_{0}(\mathbf{Z})\,\|\,\widehat{\mathbf{p}}(\mathbf{Z}))=\mathbf{p}_{0}(\mathbf{Z})^{\top}\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\widehat{\mathbf{p}}(\mathbf{Z})}\right),$
+where $a\wedge b=\min\{a,b\}$. Therefore, the expected excess risk of the estimator $\hat{\mathbf{p}}$ is defined as the expected truncated Kullback-Leibler divergence:
+$R_{B}(\mathbf{p}_{0},\widehat{\mathbf{p}}):=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[\mathrm{KL}_{B}(\mathbf{p}_{0}(\mathbf{Z})\,\|\,\widehat{\mathbf{p}}(\mathbf{Z}))\right]=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[\mathbf{p}_{0}(\mathbf{Z})^{\top}\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\widehat{\mathbf{p}}(\mathbf{Z})}\right)\right].$
 […]
-[…]
-$R(\mathbf{p}_{0},\hat{\mathbf{p}})=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[KL(\mathbf{p}_{0}(\mathbf{Z}),\hat{\mathbf{p}}(\mathbf{Z}))\right].$
-Due to the non-convexity of the function class $\mathcal{F}$, we introduce a parameter to quantify the gap between the empirical risk achieved by an estimator $\hat{\mathbf{p}}$ and the global minimum within $\mathcal{F}$.
-Formally, this empirical risk gap is defined as:
-$\Delta_{T}(\mathbf{p}_{0},\hat{\mathbf{p}}):=\mathbb{E}_{\mathcal{D}_{T}}\left[-\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}_{N+1}^{(t)\,\top}\log(\hat{\mathbf{p}}(\mathbf{Z}^{(t)}))-\min_{\mathbf{p}\in\mathcal{F}}\left(-\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}_{N+1}^{(t)\,\top}\log(\mathbf{p}(\mathbf{Z}^{(t)}))\right)\right].$ (1)
+The practice of truncating the log-likelihood ratio to ensure boundedness has been well established, see, for example, Wong and Shen (1995) and Bos and Schmidt-Hieber (2022).
+Truncation at level $B>0$ regularizes the KL divergence by controlling the influence of near-zero predictions, thereby preventing instability in the risk.
+Furthermore, due to the non-convexity of the function class $\mathcal{F}$, we introduce a parameter to quantify the gap between the expected empirical risk achieved by an estimator $\hat{\mathbf{p}}$ and the global minimum within $\mathcal{F}$.
+Formally, this expected empirical risk gap is defined as:
+$\Delta_{T}(\mathbf{p}_{0},\widehat{\mathbf{p}}):=\mathbb{E}_{\mathcal{D}_{T}}\left[-\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}_{N+1}^{(t)\,\top}\log(\widehat{\mathbf{p}}(\mathbf{Z}^{(t)}))-\min_{\mathbf{p}\in\mathcal{F}}\left(-\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}_{N+1}^{(t)\,\top}\log(\mathbf{p}(\mathbf{Z}^{(t)}))\right)\right].$ (1)
 This quantity measures the expected difference between the empirical risk of $\hat{\mathbf{p}}$ and the optimal empirical risk achievable within the class $\mathcal{F}$.
@@
 We begin by defining the model classes for our estimator, employing a transformer encoder for prompt representation and a subsequent MLP for processing the encoded output.
-This hybrid architecture—utilizing a transformer as a feature extractor with an MLP classifier atop—constitutes a prevalent design paradigm in practice.
+This hybrid architecture–utilizing a transformer as a feature extractor with an MLP classifier atop–constitutes a prevalent design paradigm in practice.
 Transformers excel at capturing complex contextual dependencies through self-attention mechanisms, generating enriched latent representations of sequential data.
@@
 Transformer architecture.
-Adapting the architecture proposed in previous theoretical work on transformer-based models Gurevych et al.
-(2022); Bai et al.
-(2023); Jiao et al.
-(2024), we adopt a transformer encoder to process the input prompt.
-Let $\mathbf{Z}\in\mathbb{R}^{D\times(N+1)}$ denote the input prompt, where $D=p+K$ is the feature dimension and $N+1$ is the in-context length.
+Adapting the architecture proposed in previous theoretical work on transformer-based models (Gurevych et al., 2022; Bai et al., 2023; Jiao et al., 2024), we adopt a transformer encoder to process the input prompt.
+Let $\mathbf{Z}\in\mathbb{R}^{D\times(N+1)}$ denote the input prompt, where $D=p+K$ is the feature dimension and $N$ is the in-context length.
 A transformer network $\boldsymbol{\phi}$ maps $\mathbf{Z}$ to an output of the same dimension through a sequence of $M$ transformer blocks, each comprising a self-attention (SA) and a feed-forward (FF) sublayer.
@@
 Each self-attention (SA) sublayer applies a multi-head attention mechanism with a residual connection:
-$F^{(\mathrm{SA})}(\mathbf{Z})=\mathbf{Z}+\sum_{s=1}^{H}\mathbf{W}_{O,s}\left(\mathbf{W}_{V,s}\mathbf{Z}\right)\left[\left(\mathbf{W}_{K,s}\mathbf{Z}\right)^{\top}\left(\mathbf{W}_{Q,s}\mathbf{Z}\right)\right]\odot\sigma_{H}\left(\left(\mathbf{W}_{K,s}\mathbf{Z}\right)^{\top}\left(\mathbf{W}_{Q,s}\mathbf{Z}\right)\right),$
+$F^{(\mathrm{SA})}(\mathbf{Z})=\mathbf{Z}+\sum_{s=1}^{H}\mathbf{W}_{O,s}\left(\mathbf{W}_{V,s}\mathbf{Z}\right)\left(\left(\mathbf{W}_{K,s}\mathbf{Z}\right)^{\top}\left(\mathbf{W}_{Q,s}\mathbf{Z}\right)\right)\odot\sigma_{H}\left(\left(\mathbf{W}_{K,s}\mathbf{Z}\right)^{\top}\left(\mathbf{W}_{Q,s}\mathbf{Z}\right)\right),$
 where $\sigma_{H}$ denotes the hardmax activation, and each attention head is parameterized by learnable projection matrices $\mathbf{W}_{Q,s},\mathbf{W}_{K,s},\mathbf{W}_{V,s},\mathbf{W}_{O,s}\in\mathbb{R}^{D\times D}$.
@@
 $\ \max(D,D^{\prime}_{r})\leq W_{1},\ \forall\ r=1,\dots,M\big{\}},$ (3)
-where $d=(p+K)(N+1)$ and $\|\cdot\|_{0}$ counts the number of nonzero elements.
+where $d=(p+K)(N+1)$, $D=p+K$, and $\|\cdot\|_{0}$ counts the number of nonzero elements.
 MLP architecture.
@@
 $\mathcal{F}_{\mathrm{id}}(L,W,S):=\bigg{\{}\mathbf{f}\in\mathcal{F}(L,W)\Big{|}\max_{j\in[L]}\left(\|\mathbf{W}_{j}\|_{\infty}\vee\|\boldsymbol{\nu}_{j}\|_{\infty}\right)\leq 1,\sum_{j=0}^{L}\left(\|\mathbf{W}_{j}\|_{0}+\|\boldsymbol{\nu}_{j}\|_{0}\right)\leq S\bigg{\}},$ (5)
-where $\|\cdot\|_{\infty}$ is the element-wise maximum norm.
+where $\|\cdot\|_{\infty}$ is the entrywise max-norm.
 Assumption 1
-Assume that the MLP class $\mathcal{F}_{\mathrm{id}}(L,W,S)$ is $L_{1}$-Lipschitz with respect to the $\ell_{\infty}$ norm, i.e.,
+The MLP class $\mathcal{F}_{\mathrm{id}}(L,W,S)$ is assumed to be $L_{1}$-Lipschitz with respect to the entrywise max-norm $\|\cdot\|_{\infty}$, that is,
 $\|\mathbf{f}(\mathbf{y}_{1})-\mathbf{f}(\mathbf{y}_{2})\|_{\infty}\leq L_{1}\|\mathbf{y}_{1}-\mathbf{y}_{2}\|_{\infty},\quad\forall\,\mathbf{f}\in\mathcal{F}_{\mathrm{id}}(L,W,S),\ \forall\,\mathbf{y}_{1},\mathbf{y}_{2}\in\mathbb{R}^{K}.$
-This assumption simplifies our analysis.
+This assumption simplifies our theoretical analysis.
 Although the Lipschitz constant $L_{1}$ of an MLP can be bounded in terms of its depth, width, and weight norms, computing its exact value is NP-hard (Virmaux and Scaman, 2018).
@@
 Let $\alpha\geq 0$ and $\mathcal{H}$ be a function class.
-A function $\mathbf{p}=(p_{1},\cdots,p_{K})\in\mathcal{H}$ is said to be $\alpha$-small value bounded (or $\alpha$-SVB) if there exists some constant $C>0$ such that for any $k\in[K]$ and $t\in(0,1]$, the following holds:
+A function $\mathbf{p}=(p_{1},\dots,p_{K})\in\mathcal{H}$ is said to be $\alpha$-small value bounded (or $\alpha$-SVB) if there exists some constant $C>0$ such that for any $k\in[K]$ and $t\in(0,1]$, the following holds:
 $\mathbb{P}\left\{p_{k}(\mathbf{Z})\leq t\right\}\leq Ct^{\alpha}.$
@@
 For any $\beta>0$ and $Q>0$, the $\beta$-Hölder-smooth class of the conditional probability functions is defined as
-$\mathcal{G}(\beta,Q):=\left\{\mathbf{p}_{0}=(p_{1}^{0},\dots,p_{K}^{0}):[0,1]^{d}\to\mathcal{S}^{K},\;p_{k}^{0}\in C^{\beta}([0,1]^{d},Q)\text{ for all }k\in[K]\right\},$ (7)
+$\mathcal{G}(\beta,Q):=\left\{\mathbf{p}_{0}=(p_{1}^{0},\dots,p_{K}^{0})^{\top}:[0,1]^{d}\to\mathcal{S}^{K},\;p_{k}^{0}\in C^{\beta}([0,1]^{d},Q)\text{ for all }k\in[K]\right\},$ (7)
 where $\mathcal{S}^{K}\subset\mathbb{R}^{K}$ denotes the probability simplex. Furthermore, we write $\mathcal{G}_{\alpha}(\beta,Q)$ for all functions in $\mathcal{G}(\beta,Q)$ that are $\alpha$-SVB.
 For a class of conditional class probabilities $\mathcal{F}$, we define its coordinate-wise logarithmic transform as
-$\log(\mathcal{F}):=\left\{\log\mathbf{p}=(\log p_{1},\dots,\log p_{K}):\mathbf{p}\in\mathcal{F}\right\}.$
-Definition 4 ($\delta$-Covering Number and $\delta$-Entropy)
+$\log(\mathcal{F}):=\left\{\log\mathbf{p}=(\log p_{1},\dots,\log p_{K})^{\top}:\mathbf{p}\in\mathcal{F}\right\}.$
+Definition 4 ($\delta$-Covering Number and Metric Entropy)
 Let $(\mathcal{V},d)$ be a pseudometric space with metric $d(\cdot,\cdot)$ and $\mathcal{F}$ be a subset of $\mathcal{V}$.
@@
 Moreover, the $\delta$-covering number of $\mathcal{F}$ with respect to the metric $d$ is the smallest cardinality of such a $\delta$-cover $\mathcal{F}(\delta)$, denoted by $\mathcal{N}(\delta,\mathcal{F},d)$.
-The quantity $\log\mathcal{N}(\delta,\mathcal{F},d)$ is referred to as the $\delta$-entropy of $\mathcal{F}$ with respect to the metric $d$.
-Definition 5 (Empirical and Uniform Supremum Covering Number)
-Given $\mathcal{Z}_{T}=\{\mathbf{Z}^{(1)},\ldots,\mathbf{Z}^{(T)}\}\subset\mathbb{R}^{d}$, we define the pseudometric $d=\|\cdot\|_{\infty,\mathcal{Z}_{T}}$ as
-$d(\mathbf{f},\mathbf{g}):=\max_{1\leq t\leq T}\max_{1\leq k\leq K}\left|\mathbf{f}_{k}(\mathbf{Z}^{(t)})-\mathbf{g}_{k}(\mathbf{Z}^{(t)})\right|.$
-The corresponding metric covering number $\mathcal{N}(\delta,\mathcal{F},\|\cdot\|_{\infty,\mathcal{Z}_{T}})$ is referred as the empirical sup-norm covering number.
-For a fixed $T$, the largest sup-norm covering number over all $\mathcal{Z}_{T}\subset\mathbb{R}^{d}$ is referred to as the uniform empirical sup-norm covering number:
+The quantity $\log\mathcal{N}(\delta,\mathcal{F},d)$ is referred to as the metric entropy of $\mathcal{F}$ at scale $\delta$ with respect to the metric $d$.
+Definition 5 (Uniform Empirical Covering Number and Entropy)
+For any given $\mathcal{Z}_{T}=\{\mathbf{Z}^{(1)},\ldots,\mathbf{Z}^{(T)}\}\subset\mathbb{R}^{d}$, we define the pseudometric $d(\mathbf{f},\mathbf{g})=\|\mathbf{f}-\mathbf{g}\|_{\infty,\mathcal{Z}_{T}}$ as
+$d(\mathbf{f},\mathbf{g})=\|\mathbf{f}-\mathbf{g}\|_{\infty,\mathcal{Z}_{T}}:=\max_{1\leq t\leq T}\max_{1\leq k\leq K}\left|\mathbf{f}_{k}(\mathbf{Z}^{(t)})-\mathbf{g}_{k}(\mathbf{Z}^{(t)})\right|.$
+The corresponding metric covering number $\mathcal{N}(\delta,\mathcal{F},\|\cdot\|_{\infty,\mathcal{Z}_{T}})$ is referred as the empirical covering number with respect to $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$.
+For a fixed $T$, the largest covering number over all $\mathcal{Z}_{T}\subset\mathbb{R}^{d}$ is referred to as the uniform empirical covering number:
 $\mathcal{N}_{\infty}(\delta,\mathcal{F},T):=\sup_{\mathcal{Z}_{T}}\mathcal{N}(\delta,\mathcal{F},\|\cdot\|_{\infty,\mathcal{Z}_{T}}).$
+The uniform empirical entropy is defined as $\mathcal{V}_{\infty}(\delta,\mathcal{F},T):=\log\mathcal{N}_{\infty}(\delta,\mathcal{F},T)$.
 Assumption 2
-Let $\mathcal{F}$ be a class of conditional class probability functions $\mathbf{p}=(p_{1},\dots,p_{K})$ used to estimate conditional class probabilities.
+Let $\mathcal{F}$ be a class of conditional class probability functions $\mathbf{p}=(p_{1},\dots,p_{K})^{\top}$ used to estimate conditional class probabilities.
 Suppose that for any $\mathbf{p}\in\mathcal{F}$, there exist two constants $C_{1}>0$ and $0<C_{2}<1$ such that $p_{k}(\mathbf{Z})\geq C_{1}/T^{C_{2}}$ for all $k\in[K]$ and $\mathbf{Z}\in[0,1]^{d}$.
 Assumption 2 requires all estimated probabilities to be bounded away from zero.
-This avoids the singularity of the KL divergence and enables the use of empirical covering numbers to control the complexity of $\log\mathcal{F}$ in the oracle inequality.
+This avoids the singularity of the KL divergence and enables the use of empirical covering numbers to control the complexity of $\log(\mathcal{F})$ in the oracle inequality.
+For the neural network class $\mathcal{F}(M,H,D,W_{1},B_{1},S_{1},\gamma,L,W,S)$ defined in (3.4), this can be achieved by adding additional layers to the MLP architecture of the network.
+More details on this issue are provided in Remarks 11 and 30.
 Remark 6
@@
 Theorem 7 (Oracle Inequality)
-Let $\mathcal{F}$ be any class of conditional class probability functions that satisfies Assumption 2, and let $\widehat{\mathbf{p}}\in\mathcal{F}$ be any estimator of $\mathbf{p}_{0}\in\mathcal{S}^{K}$.
+Let $\mathcal{F}$ be any class of conditional class probability functions that satisfies Assumption 2, and let $\widehat{\mathbf{p}}\in\mathcal{F}$ be an estimator of $\mathbf{p}_{0}\in\mathcal{S}^{K}$.
 For any truncation level $B\geq 2$ and task number $T$, the expected excess risk satisfies
-$R_{B}(\mathbf{p}_{0},\widehat{\mathbf{p}})\leq C\left(\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})+\Delta_{T}(\mathbf{p}_{0},\widehat{\mathbf{p}})+\frac{B^{2}\log(\mathcal{N}_{T})\,\log^{2}T}{T}\right),$ (8)
+$R_{B}(\mathbf{p}_{0},\widehat{\mathbf{p}})\lesssim\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})+\Delta_{T}(\mathbf{p}_{0},\widehat{\mathbf{p}})+\frac{B^{2}\log(\mathcal{N}_{T})\,\log^{2}T}{T},$ (8)
 […]
-Theorem 7 provides a general oracle inequality for any function class $\mathcal{F}$ satisfying Assumption 2.
+The proof of Theorem 7 is provided in Appendix A.
+It provides a general oracle inequality for any function class $\mathcal{F}$ satisfying Assumption 2.
 This result is independent of the specific architecture used to model the conditional class probabilities.
-In the subsequent analysis, we instantiate $\mathcal{F}$ as a class of neural networks that estimate conditional class probabilities from a given prompt, as defined in (3.4).
 Error Decomposition.
@@
 […]
-To derive the ICL risk, we focus on two key components: the approximation error $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$ and the uniform empirical covering entropy $\mathcal{N}_{T}$ in the oracle inequality (8).
-We first establish the approximation capacity of transformer networks for functions in the Hölder class $\mathcal{G}(\beta,Q)$ in the following lemma.
 Lemma 8
-Let $0<\varepsilon<1$, $\beta>0$, and $Q>0$.
-For every function $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$, there exists a transformer network $\boldsymbol{\phi}\in\mathcal{T}(M,H,D,W_{1},B_{1},S_{1},\gamma)$ with
+Let $0<\varepsilon<1$, $\beta>0$, and $Q>0$. For every function $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$, there exists a transformer network $\boldsymbol{\phi}\in\mathcal{T}$ with
 $M\lesssim\log\left(\frac{Q}{\varepsilon}\right),\quad H\lesssim\left(\frac{Q}{\varepsilon}\right)^{d/\beta},\quad D\lesssim 1,$
@@
 Its proof builds on the observation by Gurevych et al.
-(2022) that transformer architectures, such as deep ReLU networks, can represent multiplication operations ans thus approximate polynomials.
-Since polynomials can can approximate continuous functions arbitrarily well, transformer networks can effectively approximate Hölder-smooth targets.
+(2022) that transformer architectures, similar to deep ReLU networks, are capable of representing multiplication operations and therefore can approximate polynomials.
+Since polynomials can approximate continuous functions with an arbitrarily accuracy, transformer networks can effectively approximate Hölder-smooth targets.
 Without loss of generality, we assume the output of the approximating transformer lies in $[0,1]$; otherwise, a projection layer can be applied without increasing the error, as the target functions $p_{k}^{0}$ are already supported on this interval.
 Next, we apply Lemma 8 to approximate the conditional class probabilities in multiclass classification.
-Specifically, we consider the neural network class $\mathcal{F}(M,H,D,W_{1},B_{1},S_{1},\gamma,L,\\ W,S)$ defined in (3.4), which is constructed to approximate the ground-truth conditional class probabilities $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$ in the $\|\cdot\|_{\infty}$-norm, while ensuring that the output probability vectors are uniformly bounded away from zero.
-This is achieved by composing a transformer encoder with an MLP followed by a softmax transformation.
-The MLP processes the representations produced by the transformer encoder and computes a vector of logits, which are then mapped by the softmax function to class probability vectors.
-The resulting outputs are uniformly bounded away from zero and thus satisfy Assumption 2.
-A similar construction is used in Bos and Schmidt-Hieber (2022).
+Specifically, we consider the neural network class $\mathcal{F}$ defined in (3.4) to approximate the true conditional class probabilities $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$ in the $\|\cdot\|_{\infty}$-norm, while ensuring that the output probabilities are uniformly bounded away from zero.
+This is achieved by composing a transformer encoder with a MLP, followed by a softmax transformation.
+The detailed construction of the approximating network in $\mathcal{F}$ is provided in the proof of Theorem 9 in Appendix B.
 This transformer-MLP hybrid architecture is also motivated by its practical effectiveness.
@@
 Such modular designs are widely used in contemporary deep learning models and have demonstrated strong generalization performance across diverse applications.
-For simplicity, we henceforth denote the neural network classes $\mathcal{T}(M,H,D,W_{1},B_{1},S_{1},\gamma)$ defined in (3.3), $\mathcal{F}_{\mathrm{id}}(L,W,S)$ defined in (5), and $\mathcal{F}(M,H,D,W_{1},B_{1},S_{1},\gamma,L,W,S)$ defined in (3.4) as $\mathcal{T}$, $\mathcal{F}_{\mathrm{id}}$, and $\mathcal{F}$, respectively.
 Theorem 9
 Fix $\beta>0$ and $Q>0$.
-For any $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$ and any $\varepsilon>0$, there exists a neural network $\mathbf{\tilde{q}}=(\tilde{q}_{1},\cdots,\tilde{q}_{K})\in\mathcal{F}$ with
+For any $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$ and $\varepsilon\in(0,1)$, there exists a neural network $\mathbf{\tilde{q}}=(\tilde{q}_{1},\cdots,\tilde{q}_{K})\in\mathcal{F}$ with
 $M\lesssim\log\left(\frac{Q}{\varepsilon}\right),\quad H\lesssim\left(\frac{Q}{\varepsilon}\right)^{d/\beta},\quad D\lesssim 1,$
@@
 such that
-$\|{\tilde{q}}_{k}-{p}^{0}_{k}\|_{\infty}<2\varepsilon(K+1),\quad{\tilde{q}}_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2(K+1)}$
-hold for all $k\in\{1,2,\dots,K\}$ and $\mathbf{Z}\in[0,1]^{d}$.
-Theorem 9 shows that the conditional probability $\mathbf{p}_{0}$ can be uniformly approximated, with arbitrarily small error, by a neural network from $\mathcal{F}$.
+$\|{\tilde{q}}_{k}-{p}^{0}_{k}\|_{\infty}\leq 2\varepsilon(K+1),\quad{\tilde{q}}_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2K}$
+hold for all $k\in[K]$ and $\mathbf{Z}\in[0,1]^{d}$.
+Theorem 9 shows that the conditional probability $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$ can be uniformly approximated, with arbitrarily small error, by a neural network of $\mathcal{F}$.
 Its proof is provided in Appendix B.
-Based on this result, we derive a non-asymptotic upper bound for the minimal risk $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$, under uniform approximation and margin conditions.
-Remark 10
-The neural network $\mathbf{\tilde{q}}$ constructed in Theorem 9 inherently satisfies a natural lower bound on its output for every function $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$.
-Specifically, for any $\varepsilon>0$, the network’s outputs are bounded below by $\frac{\varepsilon}{2(K+1)}$, which is an intrinsic property of the architecture.
-This lower bound aligns with Assumption 2.
-In particular, by selecting an appropriate $\varepsilon$, the neural network $\mathbf{\tilde{q}}$ can be guaranteed to meet the lower bound condition on the probability functions, as specified by Assumption 2.
+Based on this result, we derive a non-asymptotic upper bound for the approximation error $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$.
 Theorem 10
-Assume $\mathbf{p}_{0}\in\mathcal{G}_{\alpha}(\beta,Q)$ for some $\alpha\in[0,1]$. Let $\mathbf{p}=(p_{1},\ldots,p_{K}):[0,1]^{d}\to\mathcal{S}^{K}$ be an approximation of $\mathbf{p}_{0}$ such that
-$\|\mathbf{p}-\mathbf{p}_{0}\|_{\infty}\leq 2(K+1)\varepsilon,\quad\min_{1\leq k\leq K}\inf_{\mathbf{Z}\in[0,1]^{d}}p_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2(K+1)}.$
+Assume $\mathbf{p}_{0}\in\mathcal{G}_{\alpha}(\beta,Q)$ for some $\alpha\in[0,1]$.
+Let $\mathbf{p}=(p_{1},\ldots,p_{K})^{\top}:[0,1]^{d}\to\mathcal{S}^{K}$ be an approximation of $\mathbf{p}_{0}$ such that
+$\|\mathbf{p}-\mathbf{p}_{0}\|_{\infty}\leq 2(K+1)\varepsilon,\quad\min_{1\leq k\leq K}\inf_{\mathbf{Z}\in[0,1]^{d}}p_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2K}.$
 Then we have
 $R(\mathbf{p}_{0},\mathbf{p})\leq C\varepsilon^{1+\alpha}K^{4+\alpha}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right).$
-If we choose $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$ with $C>0$ being a sufficiently small constant, then the risk $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$ is bounded as
-$\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})\lesssim T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}K^{4+\alpha}\log(KT).$
+If we choose $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$ for some constant $C>0$, then
+$\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})\lesssim K^{4+\alpha}T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}\log(KT).$
 Theorem 10 establishes a near-optimal convergence rate for in-context classification, recovering the classical nonparametric minimax bound when $\alpha=1$.
 The proof leverage $f$-divergence theory, bounding the KL divergence via the $\chi^{2}$-divergence (Nguyen et al., 2009; Cai and Lim, 2022).
-More details can be found in Appendix C.
-Next, we deal with the uniform empirical covering entropy $\log(\mathcal{N}_{T})$ in the Oracle Inequality. Recall that
-$\log(\mathcal{N}_{T})=\mathcal{V}_{\infty}\Big{(}\frac{1}{T},\log\mathcal{F},T\Big{)}=\sup_{\mathcal{Z}_{T}}\log\mathcal{N}\Big{(}\frac{1}{T},\log\mathcal{F},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\Big{)}.$
-The following lemma provides an upper bound on the uniform empirical covering entropy $\log(\mathcal{N}_{T})$ of the class $\log(\mathcal{F})$ under the empirical sup-norm.
+More details about the proof can be found in Appendix C.
+Remark 11
+According to the proof of Theorem 9 and Remark 30 in Appendix B, we can add additional ReLU layers to the MLP architecture of the network in $\mathcal{F}$ to ensure that $p_{k}(\mathbf{Z})\geq\varepsilon/(2K)$ for any network $\mathbf{p}=(p_{1},\dots,p_{K})^{\top}\in\mathcal{F}$ and $\varepsilon\in(0,1)$.
+Consequently, without loss of generality, we require that every network in $\mathcal{F}$ has a uniform, strictly positive lower bound $\varepsilon/(2K)$ across all its coordinates.
+Furthermore, if we choose $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$, it is readily seen that the function class $\mathcal{F}$ satisfies Assumption 2.
+This allows us to apply the oracle inequality (Theorem 7) to derive the expected excess risk for transformer-based classifiers.
+Next, we deal with the uniform empirical entropy $\log(\mathcal{N}_{T})$ in the oracle inequality. Recall that
+$\log(\mathcal{N}_{T})=\mathcal{V}_{\infty}\Big{(}\frac{1}{T},\log(\mathcal{F}),T\Big{)}=\sup_{\mathcal{Z}_{T}}\log\mathcal{N}\Big{(}\frac{1}{T},\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\Big{)},$
+where $\mathcal{Z}_{T}=\{\mathbf{Z}^{(1)},\ldots,\mathbf{Z}^{(T)}\}\subset\mathbb{R}^{d}$.
+The following lemma provides an upper bound on the uniform empirical entropy $\log(\mathcal{N}_{T})$ of the function class $\log(\mathcal{F})$.
 Lemma 12
-For any $\delta>0$, and for sufficiently small $\delta$, the following inequality holds:
-$\mathcal{V}_{\infty}\left(\delta,\log\mathcal{F},T\right)\leq\mathcal{V}_{\infty}\left(\tfrac{\delta}{4KL_{1}},\,\mathcal{T},\,T\right)+\mathcal{V}\left(\tfrac{\delta}{4K},\,\mathcal{F}_{\mathrm{id}},\,\|\cdot\|_{\infty}\right),$
+For any $\delta>0$, the following inequality holds:
+$\mathcal{V}_{\infty}\left(\delta,\log(\mathcal{F}),T\right)\leq\mathcal{V}_{\infty}\left(\tfrac{\delta}{4KL_{1}},\,\mathcal{T},\,T\right)+\mathcal{V}\left(\tfrac{\delta}{4K},\,\mathcal{F}_{\mathrm{id}},\,\|\cdot\|_{\infty}\right),$
 where $\mathcal{V}_{\infty}\left(\delta,\,\mathcal{T},\,T\right)=\sup_{\mathcal{Z}_{T}}\log\mathcal{N}(\delta,\mathcal{T},\|\cdot\|_{\infty,\mathcal{Z}_{T}})$, $\mathcal{V}(\delta,\mathcal{F}_{\rm id},\|\cdot\|_{\infty})=\log\mathcal{N}(\delta,\mathcal{F}_{\rm id},\|\cdot\|_{\infty})$, and $L_{1}$ is the Lipschitz constant specified in Assumption 1.
 The proof of Lemma 12 is provided in Appendix D.
-The argument decomposes $\log\mathcal{F}$ into a composition of the transformer class $\mathcal{T}$ and the MLP class $\mathcal{F}_{\mathrm{id}}$, and then applies standard covering number bounds for Lipschitz compositions to control the resulting entropy.
-The next result provides an upper bound on the uniform empirical covering entropy of the transformer class under the empirical sup-norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$, where $\mathcal{Z}_{T}=\{\mathbf{Z}^{(1)},\dots,\mathbf{Z}^{(T)}\}$.
+The argument decomposes $\log(\mathcal{F})$ into a composition of the transformer class $\mathcal{T}$ and the MLP class $\mathcal{F}_{\mathrm{id}}$, and then applies standard covering number bounds for Lipschitz compositions to control the resulting entropy.
+The next result provides an upper bound on the uniform empirical entropy of the transformer class.
 Lemma 13
-For any $\delta_{1}>0$, the uniform empirical covering entropy of the transformer class $\mathcal{T}$ under the empirical sup-norm satisfies
+For any $\delta_{1}>0$, the uniform empirical entropy of the transformer class $\mathcal{T}$ satisfies
 $\mathcal{V}_{\infty}\left(\delta_{1},\mathcal{T},T\right)\lesssim KM^{2}S_{1}\log\left(\max\{M,H,W_{1}\}\right)\log\left(\frac{TB_{1}}{\delta_{1}}\right),$
-where $\mathcal{T}=\mathcal{T}(M,H,D,W_{1},B_{1},S_{1},\gamma)$.
+where $\mathcal{T}=\mathcal{T}(M,H,D,W_{1},B_{1},S_{1},\gamma)$ is given in (3.3).
 The proof of Lemma 13 builds on classical techniques from the neural network literature (Anthony and Bartlett, 2009; Bartlett et al., 2019), and it is closely related to recent analyses of transformer architectures (Gurevych et al., 2022; Jiao et al., 2024).
@@
 Lemma 14
-For any $\delta_{2}>0$, the covering entropy of the function class $\mathcal{F}_{\mathrm{id}}$ under the sup-norm $\|\cdot\|_{\infty}$ satisfies the following upper bound
+For any $\delta_{2}>0$, the metric entropy of the function class $\mathcal{F}_{\mathrm{id}}$ satisfies the following upper bound
 $\mathcal{V}\left(\delta_{2},\,\mathcal{F}_{\mathrm{id}},\,\|\cdot\|_{\infty}\right)\leq 2SL\log\left(\frac{L(W+1)}{3\delta_{2}}\right),$
 where $\mathcal{F}_{\mathrm{id}}=\mathcal{F}_{\mathrm{id}}(L,W,S)$ is defined in (5).
-Lemma 14 gives an upper bound on the uniform covering entropy of the class $\mathcal{F}_{\mathrm{id}}(L,W,S)$, which consists of $S$-sparse ReLU networks with depth $L$, width $W$, and output dimension $K$, measured under the sup-norm over the input domain.
-Its proof is provided in Appendix F.
+[…]
 Theorem 15
 Assume $\mathbf{p}_{0}\in\mathcal{G}_{\alpha}(\beta,Q)$ for some $\alpha\in[0,1]$.
-Let $\hat{\mathbf{p}}$ be an estimator of $\mathbf{p}_{0}$, taking values in the function class $\mathcal{F}(M,H,D,W_{1},B_{1},S_{1},\gamma,L,W,S)$, where the parameters of the class scale with the precision parameter $\varepsilon>0$ as follows:
+Let $\hat{\mathbf{p}}$ be an estimator of $\mathbf{p}_{0}$, taking values in the function class $\mathcal{F}(M,H,D,W_{1},B_{1},S_{1},\gamma,L,W,S)$, where the parameters of this class scale with the precision parameter $\varepsilon>0$ as follows:
 $M\lesssim\log\left(\frac{Q}{\varepsilon}\right),\ H\lesssim\left(\frac{Q}{\varepsilon}\right)^{d/\beta},\ D\lesssim 1,$
@@
 and we choose $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$ for some constant $C>0$, then the ICL risk is bounded by
-$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}\right)\lesssim B^{2}T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}K^{4+\alpha}\log^{5}\left(KT\right).$
+$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}\right)\lesssim B^{2}K^{4+\alpha}T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}\log^{7}\left(KT\right).$
 In particular, for the special case where $\alpha=1$, the ICL risk achieves the bound
-$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}\right)\lesssim B^{2}T^{-\frac{2\beta}{2\beta+d}}K^{5}\log^{5}\left(KT\right).$
-Theorem 15 establishes that the ICL risk attains the minimax-optimal convergence rate (up to logarithmic factors) when $\alpha=1$.
+$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}\right)\lesssim B^{2}K^{5}T^{-\frac{2\beta}{2\beta+d}}\log^{7}\left(KT\right).$
+Theorem 15 establishes that the upper bound of the ICL risk attains the minimax optimal rate $T^{-2\beta/(2\beta+d)}$ (up to logarithmic factors) when $\alpha=1$.
 This result highlights the importance of pretraining task diversity, as observed by Raventós et al.
@@
 Our setup corresponds to the task-scaling regime formalized by Abedsoltan et al.
-(2024), where the risk decreases as the number of pretraining tasks $T$ increases, while the context length remains fixed.
+(2024), where the risk decreases as the number of pretraining tasks $T$ increases, while the context length $N$ remains fixed.
 In our multiclass setting, the achievable risk is fundamentally determined by the number of pretraining tasks $T$.
-Since large language models are trained on virtually unlimited tasks, our bound provides a theoretical explanation for their strong in-context learning performance—even with only a small number of in-context examples at inference time.
-Proof of Theorem 15.
-The proof is based on the oracle inequality (8), which decomposes the error into an approximation error $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$, controlled by Theorem 11, and a complexity term $(\log(\mathcal{N}_{T})\log^{2}T)/T$, bounded using Lemmas 12, 13, and 14, with $\delta=1/T$.
-The approximation scale $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$ is chosen to balance the approximation error and the entropy term, reflecting the trade-off between network expressiveness and statistical complexity.
-These bounds together yields our result.
-$\blacksquare$
+Since large language models are trained on virtually unlimited tasks, our bound provides a theoretical explanation for their strong in-context learning performance, even with only a small number of in-context examples at inference time.
+[…]
 5 Lower Bounds
 To establish the optimality of the convergence rate in Theorem 15, we derive a matching minimax lower bound for the ICL risk.
-Our argument leverages the classical minimax theory for nonparametric estimation under the squared Hellinger distance translate these results to the KL divergence setting.
-The squared Hellinger distance, defined as $H^{2}(P,Q)=\frac{1}{2}\int(\sqrt{dP}-\sqrt{dQ})^{2}$, is a fundamental tool for deriving statistical lower bounds due to its favorable metric properties (van der Vaart and Wellner, 1996; Geer, 2000).
-The rate $T^{-\frac{2\beta}{2\beta+d}}$ is known to be minimax-optimal for estimation under the squared Hellinger loss (Tsybakov, 2008, Chapter 2).
-The key to translating this result is the classical inequality $2H^{2}(P,Q)\leq\text{KL}(P,Q)$, which holds under bounded likelihood ratio conditions, see (7.5) in Birgé and Massart (1998) and Lemma 4 in Haussler and Opper (1997).
-Conversely, under some integrability assumptions, the KL divergence can be bounded by the Hellinger distance up to logarithmic factors (Wong and Shen, 1995, Theorem 5).
+Our argument leverages the classical minimax theory for nonparametric estimation under the squared Hellinger distance and translates these results to the KL divergence setting.
+The squared Hellinger distance between two probability measures $P$ and $Q$ is defined as $H^{2}(P,Q)=\frac{1}{2}\int(\sqrt{dP}-\sqrt{dQ})^{2}$, which is a fundamental tool for deriving statistical lower bounds due to its favorable metric properties (van der Vaart and Wellner, 1996; van de Geer, 2000).
+The rate $T^{-2\beta/(2\beta+d)}$ is known to be minimax optimal for nonparametric estimation under the squared Hellinger loss (Tsybakov, 2008, Chapter 2).
+The key to translating this result is the classical inequality $2H^{2}(P,Q)\leq\text{KL}(P\|Q)$, see Lemma 4 in Haussler and Opper (1997) and (7.5) in Birgé and Massart (1998).
+Conversely, under certain integrability assumptions, the KL divergence can be bounded by the Hellinger distance up to logarithmic factors (Wong and Shen, 1995, Theorem 5).
 This two-sided relationship enables the direct translation of minimax lower bounds from the Hellinger distance to the KL divergence.
@@
 Lemma 16
-Suppose $B\geq 2$. Then the squared Hellinger distance and the truncated KL divergence satisfy
-$H^{2}(P,Q)\leq\tfrac{1}{2}\mathrm{KL}_{2}(P,Q)\leq\tfrac{1}{2}\mathrm{KL}_{B}(P,Q)\leq 5e^{B/2}H^{2}(P,Q).$
+Let $P$ and $Q$ denote two probability measures defined on the same measurable space. Then for any $B\geq 2$,
+$H^{2}(P,Q)\leq\tfrac{1}{2}\mathrm{KL}_{2}(P\|Q)\leq\tfrac{1}{2}\mathrm{KL}_{B}(P\|Q)\leq 5e^{B/2}H^{2}(P,Q).$
 […]
 Theorem 17
-Consider the same setting as in Theorem 15. Then, the minimax lower bound of the ICL risk is given by
-$\inf_{\hat{\mathbf{p}}}\sup_{\mathbf{p}_{0}\in\mathcal{G}_{1}(\beta,Q)}R_{B}(\mathbf{p}_{0},\hat{\mathbf{p}})\gtrsim T^{-\frac{2\beta}{2\beta+d}}.$
-This result aligns with the upper bound in Theorem 15 up to constant factors, thereby establishing the minimax-optimality of the rate $T^{-\frac{2\beta}{2\beta+d}}$ for ICL under the truncated KL loss when $\alpha=1$.
-The case of general $\alpha\in(0,1)$ remains open and will be studied in future work.
+Consider the same setting as in Theorem 15. Then the minimax lower bound of the ICL risk with $\alpha=1$ is given by
+$\inf_{\hat{\mathbf{p}}}\sup_{\mathbf{p}_{0}\in\mathcal{G}_{1}(\beta,Q)}R_{B}(\mathbf{p}_{0},\hat{\mathbf{p}})\gtrsim T^{-\frac{2\beta}{2\beta+d}},$
+where $\mathcal{G}_{\alpha}(\beta,Q)=\{\mathbf{p}_{0}:\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)\ {\rm and}\ \mathbf{p}_{0}\ is\ \alpha-SVB\}$.
+This result aligns with the upper bound in Theorem 15 up to logarithmic factors, thereby establishing the minimax optimality of the convergence rate for the ICL risk under the truncated Kullback-Leibler loss when $\alpha=1$.
+For general $\alpha\in(0,1)$, the minimax optimal rate of the expected excess risk under the truncated KL divergence remains unknown.
+We conjecture that the optimal rate is given by $T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}$, up to logarithmic factors.
+However, a rigorous proof is beyond the scope of the current paper and warrants further study.
 6 In-Context Learning Beyond Transformers
@@
 However, recent literature suggests that ICL is not exclusive to transformer architectures.
-Tong and Pehlevan (2025) empirically demonstrated that MLPs can also learn in-context on standard synthetic ICL taks.
+Tong and Pehlevan (2025) empirically demonstrated that MLPs can also learn in-context on standard synthetic ICL tasks.
 Similarly, Abedsoltan et al.
@@
 In this section, we support these insights by analyzing the expected excess risk of MLP-based classifiers in the setting of multiclass ICL.
-Under appropriate assumptions, we demonstrate that MLPs can achieve minimax-optimal convergence rates for ICL.
-Our analysis of convergence rates provides theoretical support for recent empirical evidence on the effectiveness of MLPs.
+Under appropriate assumptions, we demonstrate that MLPs can achieve the minimax optimal convergence rate for ICL.
+Our analysis provides theoretical support for recent empirical evidence on the effectiveness of MLPs.
 This finding suggests that, while attention mechanisms are powerful, they are not strictly necessary for enabling ICL.
@@
 $\mathcal{F}_{\rm mlp}(L,W,S):=\left\{\tilde{\mathbf{p}}(\mathbf{Z})=\boldsymbol{\Phi}\circ\mathbf{f}(\operatorname{vec}(\mathbf{Z}))\;\middle|\;\mathbf{f}\in\mathcal{F}_{\mathrm{id}}(L,W,S)\right\},$ (9)
-where $\boldsymbol{\Phi}$ denotes the softmax function.
+where $\boldsymbol{\Phi}$ denotes the softmax function and $\operatorname{vec}(\mathbf{Z})$ is the vectorization of the matrix $\mathbf{Z}$.
 Lemma 18
-For every $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$ and $0<\varepsilon<\min\{(K(4+C_{Q,\beta,d}))^{-1},(\beta+1)^{-\beta},(Q+1)^{-\beta/d}e^{-\beta}\}$ with $C_{Q,\beta,d}=(2Q+1)(1+d^{2}+\beta^{2})6^{d}+Q3^{\beta}$, there exist a neural network $\tilde{\mathbf{p}}\in\mathcal{F}_{\rm MLP}(L,W,S)$ with
+For any $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$ and $\varepsilon\in(0,\min\{(K(4+C_{Q,\beta,d}))^{-1},(\beta+1)^{-\beta},(Q+1)^{-\beta/d}e^{-\beta}\})$ with $C_{Q,\beta,d}=(2Q+1)(1+d^{2}+\beta^{2})6^{d}+Q3^{\beta}$, there exist a neural network $\tilde{\mathbf{p}}\in\mathcal{F}_{\rm mlp}(L,W,S)$ with
 $L\lesssim\log\left(\frac{1}{\varepsilon}\right),\quad W\lesssim K\varepsilon^{-d/\beta},\quad S\lesssim K\varepsilon^{-d/\beta}\log\left(\frac{1}{\varepsilon}\right),$
 such that
-$\tilde{p}_{k}\geq\frac{\varepsilon}{4},\quad\|\tilde{p}_{k}-p_{k}^{0}\|_{\infty}\leq\frac{1}{2}K(C_{Q,\beta,d}+4)\varepsilon,$
-for all $k=1,\dots,K$.
-This result slightly modifies Lemma 4.3 in Bos and Schmidt-Hieber (2022), which builds upon a line of work showing that ReLU MLPs can efficiently approximate $\beta$-Hölder smooth functions, with $L\lesssim\log(1/\varepsilon)$, $W\lesssim\varepsilon^{-d/\beta}$, $s\lesssim\varepsilon^{-d/\beta}\log(1/\varepsilon)$ (Yarotsky, 2017; Petersen and Voigtlaender, 2018; Schmidt-Hieber, 2020; Jiao et al., 2023; Chakraborty and Bartlett, 2024b).
+$\|\tilde{p}_{k}-p_{k}^{0}\|_{\infty}\leq\frac{1}{2}K(C_{Q,\beta,d}+4)\varepsilon,\quad\tilde{p}_{k}(\mathbf{Z})\geq\frac{\varepsilon}{4}$
+for all $k\in[K]$ and $\mathbf{Z}\in[0,1]^{d}$.
+This result is a slight modification of Lemma 4.3 in Bos and Schmidt-Hieber (2022), which builds upon a line of work (Yarotsky, 2017; Schmidt-Hieber, 2020; Jiao et al., 2023; Chakraborty and Bartlett, 2024) showing that ReLU MLPs with depth $L\lesssim\log(1/\varepsilon)$, width $W\lesssim\varepsilon^{-d/\beta}$, and size $S\lesssim\varepsilon^{-d/\beta}\log(1/\varepsilon)$ can efficiently approximate $\beta$-Hölder smooth functions.
+Building on Lemma 18, we then derive a non-asymptotic upper bound for the approximation error $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$.
+Theorem 19
+Assume $\mathbf{p}_{0}\in\mathcal{G}_{\alpha}(\beta,Q)$ for some $\alpha\in[0,1]$ and $\mathbf{p}=(p_{1},\ldots,p_{K})^{\top}:[0,1]^{d}\to\mathcal{S}^{K}$ is an approximation of $\mathbf{p}_{0}$ such that $\|\mathbf{p}-\mathbf{p}_{0}\|_{\infty}\leq C_{0}\varepsilon/4$ and $\min_{k}\inf_{\mathbf{Z}\in[0,1]^{d}}p_{k}(\mathbf{Z})\geq\varepsilon/4$ with $C_{0}$ being a positive constant.
+Then we have
+$R(\mathbf{p}_{0},\mathbf{p})\lesssim K(C_{0}+1)^{2+\alpha}\varepsilon^{1+\alpha}\left(1+\frac{\mathbf{1}(\alpha<1)}{1-\alpha}+\log\left(\frac{4}{\varepsilon}\right)\right).$
+Furthermore, if we choose $C_{0}=2K(C_{Q,\beta,d}+4)$ and $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$ for some constant $C>0$, then
+$\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})\lesssim K^{3+\alpha}T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}\log(KT).$
+The proof of this theorem follows the same arguments as in Theorem 10. A similar result can be found in Bos and Schmidt-Hieber (2022, Theorem 3.2).
+Remark 20
+Similar to the arguments in Remark 30, by adding additional ReLU layers to the network in $\mathcal{F}_{\rm id}(L,W,S)$, we can assume that every network $\mathbf{G}=(G_{1},\dots,G_{p})^{\top}\in\mathcal{F}_{\rm id}(L,W,S)$ satisfies $\log(\varepsilon)\leq G_{i}\leq\log(2)$ for all $1\leq i\leq p$.
+Since $\mathcal{F}_{\rm mlp}(L,W,S)=\mathbf{\Phi}(\mathcal{F}_{\rm id}(L,W,S))$, it follows that every network in $\mathcal{F}_{\rm mlp}(L,W,S)$ has a uniform, strictly positive lower bound of $\varepsilon/2K$ for all its coordinates.
+Given that the neural network $\tilde{p}_{k}\geq\varepsilon/4\geq\varepsilon/2K$, this implies that $\tilde{\mathbf{p}}\in\mathcal{F}_{\rm mlp}(L,W,S)$.
+Therefore, with a suitable chosen precision parameter $\varepsilon$, the function class $\mathcal{F}_{\rm mlp}(L,W,S)$ satisfies Assumption 2.
+This allows us to apply the oracle inequality to derive the expected excess risk for MLP-based classifiers in the ICL setting.
+The next lemma establishes an explicit upper bound on the uniform empirical entropy of $\mathcal{F}_{\rm mlp}(L,W,S)$, a key quantity in deriving the generalization error bound in the oracle inequality.
+We write $\mathcal{F}_{\rm mlp}(L,W,S)=\mathbf{\Phi}(\mathcal{F}_{\mathrm{id}}(L,W,S))$.
 Lemma 21
-Consider the function class $\mathcal{F}_{\mathrm{MLP}}(L,W,S)$ defined in (9).
-For any $T\in\mathbb{N}$, the empirical covering entropy of the function class $\log(\mathcal{F}^{\prime}_{\Phi})$ under the empirical sup-norm satisfies
-$\mathcal{V}_{\infty}\left(\tfrac{1}{T},\log(\mathcal{F}_{\rm MLP}(L,W,S)),T\right)\leq C\cdot SL\log S\cdot\log T,$
-for some universal constant $C>0$.
-The proof of Lemma 19 is provided in Appendix H. Building on Lemmas 18 and 19, we establish an upper bound on the expected excess risk for MLP-based classifiers in the ICL setting, stated as follows.
+Suppose $\|\mathbf{G}\|_{\infty}\leq B_{2}<\infty$ for any $\mathbf{G}\in\mathcal{F}_{\mathrm{id}}(L,W,S)$. Then
+$\mathcal{V}_{\infty}\left(\frac{1}{T},\log(\mathcal{F}_{\rm mlp}(L,W,S)),T\right)\leq CKSL\log(S)\log(B_{2}KT),$
+where $\mathcal{V}_{\infty}\left(1/T,\log(\mathcal{F}_{\rm mlp}(L,W,S)),T\right)=\sup_{\mathcal{Z}_{T}}\log\mathcal{N}(1/T,\log(\mathcal{F}_{\rm mlp}(L,W,S)),\|\cdot\|_{\infty,\mathcal{Z}_{T}})$, and $C>0$ is an universal constant.
+The proof of Lemma 21 is provided in Appendix H.
+Building on the oracle inequality (Theorem 7), Theorem 19, and Lemma 21, we derive an upper bound on the expected excess risk for MLP-based classifiers in the ICL setting, as stated below.
 Theorem 22
 Assume $\mathbf{p}_{0}\in\mathcal{G}_{\alpha}(\beta,Q)$ for some $\alpha\in[0,1]$.
-Let $\hat{\mathbf{p}}$ be an estimator of $\mathbf{p}_{0}$, taking values in the function class $\mathcal{F}_{\rm MLP}(L,W,S)$, where the parameters of the class scale with the precision parameter $\varepsilon>0$ as follows:
+Let $\hat{\mathbf{p}}_{\rm mlp}$ be an estimator of $\mathbf{p}_{0}$, taking values in the function class $\mathcal{F}_{\rm mlp}(L,W,S)$, where the parameters of this class scale with the precision parameter $\varepsilon>0$ as follows:
 $L\lesssim\log\left(\frac{1}{\varepsilon}\right),\ W\lesssim K\varepsilon^{-d/\beta},\ S\lesssim K\varepsilon^{-d/\beta}\log\left(\frac{1}{\varepsilon}\right).$
-[…]
-$\Delta_{T}(\mathbf{p}_{0},\widehat{\mathbf{p}})\lesssim T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}K^{4+\alpha},$
-and we choose the precision parameter $\varepsilon$ as
-$\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}},$
-for some constant $C>0$, then the ICL risk is bounded by
-$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}\right)\lesssim B^{2}T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}K^{4+\alpha}\log^{4}\left(KT\right).$
+If we choose the precision parameter $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$ for some positive constant $C$, and the expected empirical risk gap satisfies
+$\Delta_{T}(\mathbf{p}_{0},\widehat{\mathbf{p}}_{\rm mlp})\lesssim K^{5+\alpha}T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}},$
+then the ICL risk is bounded by
+$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}_{\rm mlp}\right)\lesssim B^{2}K^{5+\alpha}T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}\log^{6}\left(KT\right).$
 In particular, for the special case where $\alpha=1$, the ICL risk achieves the bound
-$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}\right)\lesssim B^{2}T^{-\frac{2\beta}{2\beta+d}}K^{5}\log^{4}\left(KT\right).$
+$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}_{\rm mlp}\right)\lesssim B^{2}K^{6}T^{-\frac{2\beta}{2\beta+d}}\log^{6}(KT).$
 Remark 23
-We first show that the risk bounds in Theorems 20 and 15 both achieve the optimal rate $T^{-\frac{2\beta}{2\beta+d}}$, up to logarithmic factors.
+We have shown that the ICL risk bounds in Theorems 15 and 22 both achieve the optimal rate $T^{-2\beta/(2\beta+d)}$, up to logarithmic factors.
 Recent works have empirically shown that MLPs can also exhibit ICL behavior (Tong and Pehlevan, 2025; Kratsios and Furuya, 2025; Abedsoltan et al., 2024).
@@
 Proof of Theorem 22.
-The approximation error is controlled as shown in Lemma 18.
-Using these bounds, we construct an estimator $\hat{\mathbf{p}}\in\mathcal{F}_{MLP}(L,W,S)$ such that $\|\hat{\mathbf{p}}-\mathbf{p}_{0}\|_{\infty}\leq\varepsilon$.
-Applying Theorems 711, and Lemma 19, and expressing the network parameters as functions of $\varepsilon$, we obtain the claimed in-context risk bound.
+Recall that $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$ and $\|\mathbf{G}\|_{\infty}\leq|\log(\varepsilon)|$ for any $\mathbf{G}\in\mathcal{F}_{\mathrm{id}}(L,W,S)$, it follows that $\|\mathbf{G}\|_{\infty}\leq C\log(KT)$ for some positive constant $C$.
+Applying Theorems 719, Lemmas 18 and 21, and expressing the network parameters as functions of $\varepsilon$, we obtain the claimed in-context risk bound.
 $\blacksquare$
 7 Discussions
-The upper bound established in Theorem 15 reflects the curse of dimensionality, as the input prompt dimension $d$ explicitly appears in the convergence rate exponent.
+The upper bounds established in Theorem 15 and  22 reflect the curse of dimensionality, as the input prompt dimension $d$ explicitly appears in the convergence rate exponent.
 This phenomenon is well-known in nonparametric estimation under smoothness constraints, where the optimal minimax rate for estimating a $\beta$-Hölder function scales as $T^{-2\beta/(2\beta+d)}$ (Donoho and Johnstone, 1998).
 Our result demonstrates that this dependence persists in transformer-based ICL under general smoothness assumptions.
-Nevertheless, the practical impact of dimensionality may be mitigated, as the number of tasks $T$ is typically large in realistic ICL scenarios.
+Nevertheless, the practical impact of dimensionality may be mitigated, as the number of tasks $T$ is typically extremely large in realistic ICL scenarios.
 Furthermore, transformer architectures are believed to possess an inherent ability to adapt to latent low-dimensional structure within the input, potentially reducing their effective dependence on the prompt dimension.
@@
 The manifold hypothesis (Pope et al., 2021) suggests that although data appear in high-dimensional spaces, their intrinsic structure is governed by a low-dimensional manifold (Tenenbaum et al., 2000; Roweis and Saul, 2000).
-Under this assumption, deep neural networks have been shown to achieve significantly faster convergence rates by effectively adapting to the intrinsic dimensionality of data (Nakada and Imaizumi, 2020; Chen et al., 2022, 2019; Schmidt-Hieber, 2019; Cloninger and Klock, 2021).
-Recent studies further characterize the statistical properties of deep learning and generative models under low intrinsic dimensionality, providing refined generalization guarantees in these settings (Chakraborty and Bartlett, 2024a, b).
+Under this assumption, deep neural networks have been shown to achieve significantly faster convergence rates by effectively adapting to the intrinsic dimensionality of data (Schmidt-Hieber, 2019; Chen et al., 2019, 2022; Nakada and Imaizumi, 2020; Cloninger and Klock, 2021).
+Recent studies further characterize the statistical properties of deep learning and generative models under low intrinsic dimensionality, providing refined generalization guarantees in these settings (Chakraborty and Bartlett, 2024, 2025).
 Despite these advances, the theoretical understanding of transformers in low-dimensional geometric settings remains limited.
 A recent contribution by Havrilla and Liao (2024) analyzes transformer approximation for a regression problem on low-dimensional manifolds, whose setting is quite different from ours.
-Extending Lemma 8 to manifold-supported functions deserve further study, as this would allow risk bounds to scale with the intrinsic dimension rather than the ambient dimension.
+Extending Lemma 8 to manifold-supported functions warrants further study, as this would allow risk bounds to scale with the intrinsic dimension rather than the ambient dimension.
+Acknowledgments and Disclosure of Funding
+We gratefully acknowledge the support of the National Natural Science Foundation of China (NSFC12471276, NSFC12131006, NSFC12201048, NSFC12301383, NSFC12471276) and the Scientific and Technological Innovation Project of China Academy of Chinese Medical Science (CI2023C063YLL).
 Appendix A Proof of Theorem 7
 Proof  We begin with some notations that will be used in the proof.
-  • The expected KL risk defined as:
+  • The expected excess risk defined as:
     $R_{B}(\mathbf{p}_{0},\hat{\mathbf{p}})=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[\boldsymbol{p}_{0}(\mathbf{Z})^{\top}\left(B\wedge\log\left(\frac{\boldsymbol{p}_{0}(\mathbf{Z})}{\hat{\boldsymbol{p}}(\mathbf{Z})}\right)\right)\right].$
@@
     Here, we omit the subscript $N+1$ from $\mathbf{y}_{N+1}^{(t)}$ for simplicity.
-  • The empirical risk gap is defined as:
+  • The expected empirical risk gap is defined as:
     $\Delta_{T}(\boldsymbol{p}_{0},\hat{\boldsymbol{p}}):=\mathbb{E}_{\mathcal{D}_{T}}\left[-\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}^{(t)\top}\log(\hat{\boldsymbol{p}}(\mathbf{Z}^{(t)}))-\min_{\boldsymbol{p}\in\mathcal{F}}\left(-\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}^{(t)\top}\log(\boldsymbol{p}(\mathbf{Z}^{(t)}))\right)\right].$
@@
 $\bigg{|}\mathbb{E}_{\mathcal{D}_{T},(\mathbf{Z},\mathbf{y})}y_{k}\bigg{[}B\wedge\log\frac{p_{k}^{0}(\mathbf{Z})}{\widehat{p}_{k}(\mathbf{Z})}\bigg{]}\mathbb{I}\left(p_{k}^{0}(\mathbf{Z})<\frac{C_{T}}{T}\right)\bigg{|}\leq\frac{C_{T}\{\log(T/C_{T})+B\}}{T}$
-$\bigg{|}\mathbb{E}_{\mathcal{D}_{T}}\left[y_{k}^{(t)}\left(B\wedge\log\frac{p_{k}^{0}(\mathbf{Z}^{(t)})}{\hat{p}_{k}(\mathbf{Z}^{(t)})}\right)\mathbb{I}\left(p_{k}^{0}(\mathbf{Z}^{(t)})<\frac{C_{T}}{T}\right)\right]\bigg{|}\leq\frac{C_{T}\{\log(T/C_{T})+B\}}{T},$
+$\bigg{|}\mathbb{E}_{\mathcal{D}_{T}}\left[y_{k}^{(t)}\left(B\wedge\log\frac{p_{k}^{0}(\mathbf{Z}^{(t)})}{\widehat{p}_{k}(\mathbf{Z}^{(t)})}\right)\mathbb{I}\left(p_{k}^{0}(\mathbf{Z}^{(t)})<\frac{C_{T}}{T}\right)\right]\bigg{|}\leq\frac{C_{T}\{\log(T/C_{T})+B\}}{T},$
 for any $t\in\{1\dots T\}$ and $k\in\{1,\dots,K\}$.
@@
 Similarly,
-$\bigg{|}\mathbb{E}_{\mathcal{D}_{T}}\left[y_{k}^{(t)}\left(B\wedge\log\frac{p_{k}^{0}(\mathbf{Z}^{(t)})}{\hat{p}_{k}(\mathbf{Z}^{(t)})}\right)\mathbb{I}\left(p_{k}^{0}(\mathbf{Z}^{(t)})<\frac{C_{T}}{T}\right)\right]\bigg{|}$
+$\bigg{|}\mathbb{E}_{\mathcal{D}_{T}}\left[y_{k}^{(t)}\left(B\wedge\log\frac{p_{k}^{0}(\mathbf{Z}^{(t)})}{\widehat{p}_{k}(\mathbf{Z}^{(t)})}\right)\mathbb{I}\left(p_{k}^{0}(\mathbf{Z}^{(t)})<\frac{C_{T}}{T}\right)\right]\bigg{|}$
 $\leq$ $\mathbb{E}_{\mathcal{D}_{T}}\left[y_{k}^{(t)}\{|\log(p_{k}^{0}(\mathbf{Z}^{(t)}))|+B\}\mathbb{I}\left(p_{k}^{0}(\mathbf{Z}^{(t)})<\frac{C_{T}}{T}\right)\right]$
@@
 $\leq$ $\frac{C_{T}\{\log(T/C_{T})+B\}}{T}.$
-This complete the proof of this lemma.
+This completes the proof of Lemma 24.
 Lemma 25
@@
 Adding $\mathbb{E}_{\mathbf{Z}}[\mathbf{p}_{0}(\mathbf{Z})^{\top}\log(\mathbf{p}_{0}(\mathbf{Z}))]$ to both sides yields the desired inequality.
-Recall that $\mathbf{p}(\mathbf{Z})=(p_{1}(\mathbf{Z}),\dots,p_{K}(\mathbf{Z}))$ and $g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})=\mathbf{y}^{\top}\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\mathbf{p}(\mathbf{Z})}\right)_{\geq C_{T}/T}$, where $B\geq 2$.
+Recall that $\mathbf{p}(\mathbf{Z})=(p_{1}(\mathbf{Z}),\dots,p_{K}(\mathbf{Z}))^{\top}$and $g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})=\mathbf{y}^{\top}\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\mathbf{p}(\mathbf{Z})}\right)_{\geq C_{T}/T}$, where $B\geq 2$.
 We then have the following result, which is used to derive the oracle inequality.
@@
 $\geq\ \sum_{k=1}^{K}y_{k}\left[B\wedge\log\left(\frac{C_{T}}{T}\right)\right]=\log\left(\frac{C_{T}}{T}\right).$
-Write $B_{T}=B\vee\left|\log(C_{T}/T)\right|$, it follows that $|g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})|\leq B_{T}$. The rest of the proof will be divided into six steps to enhance the readability.
+Write $B_{T}=B\vee\left|\log(C_{T}/T)\right|$, and it follows that $|g_{\mathbf{p}}(\mathbf{Z},\mathbf{y})|\leq B_{T}$. The rest of the proof will be divided into six steps to enhance the readability.
 Step 1. Symmetrization.
 Let $\tilde{\mathcal{D}}_{T}=\{(\mathbf{\tilde{Z}}^{(t)},\mathbf{\tilde{y}}^{(t)})\}_{t=1}^{T}$ be i.i.d.
-random variables with same distribution as $(\mathbf{Z},\mathbf{y})$ and independent of $\mathcal{D}_{T}$.
+random variables with the same distribution as $(\mathbf{Z},\mathbf{y})$ and independent of $\mathcal{D}_{T}$.
 Consider the event
@@
 $\leq$ $B\mathbb{E}\left(\sum_{k=1}^{K}p_{k}^{0}(\mathbf{Z})\left[B\wedge\log\left(\frac{p_{k}^{0}(\mathbf{Z})}{p_{k}(\mathbf{Z})}\right)\right]\right),$
-where the second equation is due to the fact that $\mathbf{y}$ is one-hot encoded, and the second inequality follows from Lemma 25 in the following.
+where the second equation is due to the fact that $\mathbf{y}$ is one-hot label vector, and the second inequality follows from Lemma 27 in the following.
 Since for any $\mathbf{p}\in\mathcal{F}$, we have $p_{k}(\mathbf{Z})\geq C_{1}/T^{C_{2}}$ for some constants $C_{1}>0$ and $0<C_{2}<1$, and $C_{T}\leq C_{1}T^{1-C_{2}}$, it follows that
@@
 $+\frac{\epsilon(1-\epsilon)}{4(1+\epsilon)B_{T}}\frac{1}{T}\sum_{t=1}^{T}g^{2}_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-\frac{\epsilon^{2}(a+b)}{4(1+\epsilon)B_{T}}\Bigg{\}}.$
-Now we investigate the last probability in (21) condition on $\mathcal{D}_{T}$.
-Fixing $\mathcal{D}_{T}=\{(\mathbf{Z}^{(t)},\mathbf{y}^{(t)}):t=1,\dots,T\}$, let $\delta>0$ and let $\mathcal{G}_{\delta}=\{g_{i}:i=1,\dots,M\}$ be a $\delta$-cover of $\mathcal{F}_{1}=\{g_{\mathbf{p}}:\mathbf{p}\in\mathcal{F}\}$ with respect to the norm $\|\cdot\|_{L_{1},\mathcal{D}_{T}}$ on the sample $\mathcal{D}_{T}$, where $M=N(\delta,\mathcal{F}_{1},\|\cdot\|_{L_{1},\mathcal{D}_{T}})$.
-Without loss of generalization, we assume that $\mathcal{G}_{\delta}\subset\mathcal{F}_{1}$.
-For a given $\mathbf{p}\in\mathcal{F}$, there exist $g_{i}\in\mathcal{F}_{1}$ such that
+Now we study the last probability in (21), conditioned on $\mathcal{D}_{T}$.
+Fix $\mathcal{D}_{T}=\{(\mathbf{Z}^{(t)},\mathbf{y}^{(t)}):t=1,\dots,T\}$, let $\delta>0$, and define $\mathcal{G}_{\delta}=\{g_{i}:i=1,\dots,M\}$ as a $\delta$-cover of $\mathcal{F}_{1}=\{g_{\mathbf{p}}:\mathbf{p}\in\mathcal{F}\}$ with respect to $\|\cdot\|_{L_{1},\mathcal{D}_{T}}$, where $M=N(\delta,\mathcal{F}_{1},\|\cdot\|_{L_{1},\mathcal{D}_{T}})$.
+Without loss of generalization, we assume that $\mathcal{G}_{\delta}\subset\mathcal{F}_{1}$. For a given $\mathbf{p}\in\mathcal{F}$, there exist $g_{i}\in\mathcal{F}_{1}$ such that
 $\frac{1}{T}\sum_{t=1}^{T}\big{|}g_{i}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})-g_{\mathbf{p}}(\mathbf{Z}^{(t)},\mathbf{y}^{(t)})\big{|}\leq\delta.$
@@
 $\leq$ $10\sup_{\mathcal{Z}_{T}}\mathcal{N}(\frac{\epsilon b}{8},\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}})\exp\Bigg{(}-\frac{\epsilon^{2}(1-\epsilon)aT}{15(1+\epsilon)B_{T}^{2}}\Bigg{)}.$
-This completes the proof of this Lemma.
-The following result is the Lemma 3.7 in Bos and Schmidt-Hieber (2022), which is used in Step 1 of the proof of Lemma 24 with $m=2$.
+This completes the proof of Lemma 26.
+The following lemma, adapted from Lemma 3.7 in Bos and Schmidt-Hieber (2022), is used in Step 1 of the proof of Lemma 26 with $m=2$.
 We call $(p_{1},\dots,p_{K})\in\mathbb{R}^{K}$ a probability vector if $\sum_{i=1}^{K}p_{i}=1$ and $p_{i}\geq 0$ for all $1\leq i\leq K$.
@@
 Appendix B Proof of Theorem 9
-Proof  By composing the neural networks established in Lemma 8 and Lemma 27, we construct a neural network $\mathbf{G}(\boldsymbol{\phi})=\big{(}G(\phi_{1}),\dots,G(\phi_{K})\big{)}$ such that, for any $\varepsilon\in(0,1)$,
+The following lemma is a key component in the proof of Theorem 9.
+It provides an efficient approximation of the logarithm function by a scalar-valued neural network with controlled complexity.
+This result is adapted from Theorem 4.1 of Bos and Schmidt-Hieber (2022), with some modifications to the notations.
+Lemma 28 (Theorem 4.1 of Bos and Schmidt-Hieber (2022))
+For any $\varepsilon\in(0,1)$ and $\beta>0$, there exists a scalar-valued ReLU neural network $G:[0,1]\to\mathbb{R}$ with the depth $L$, width $W$, and size $S$ (the total number of nonzero parameters) satisfying
+$L\lesssim\log\left(\frac{1}{\varepsilon}\right),\quad W\lesssim\left(\frac{1}{\varepsilon}\right)^{\frac{1}{\beta}},\quad S\lesssim\left(\frac{1}{\varepsilon}\right)^{\frac{1}{\beta}}\log\left(\frac{1}{\varepsilon}\right),$
+such that
+$\left|e^{G(x)}-x\right|\leq\varepsilon,\quad\text{and}\quad G(x)\geq\log(\varepsilon),\quad\text{for all }x\in[0,1].$
+Remark 29
+Note that the neural network $G$ constructed in Lemma 28 also satisfies $G(x)\leq\log(x+\varepsilon)\leq\log(2)$. This upper bound of $G(x)$ will be used in the proof of Theorem 9.
+Proof [Proof of Theorem 9] By composing the neural networks established in Lemma 8 and Lemma 28, we construct a neural network $\mathbf{G}(\boldsymbol{\phi})=\big{(}G(\phi_{1}),\dots,G(\phi_{K})\big{)}$ such that, for any $\varepsilon\in(0,1)$,
 $\|\boldsymbol{\phi}-\mathbf{p}_{0}\|_{\infty}\leq\varepsilon\quad{\rm and}\quad\|e^{\mathbf{G}(\boldsymbol{\phi})}-\boldsymbol{\phi}\|_{\infty}\leq\varepsilon,$
-where $\boldsymbol{\phi}=(\phi_{1},\phi_{2},\dots,\phi_{K})^{\top}\in\mathcal{T}(M,D,H,W_{1},B_{1},S_{1},\gamma)$ and $\mathbf{G}\in\mathcal{F}(L,KW,KS)$. Consequently, for any $k=1,\dots,K$, we have
-$\left\|e^{G(\phi_{k})}-p_{k}^{0}\right\|_{\infty}\leq\left\|e^{G(\phi_{k})}-\phi_{k}\right\|_{\infty}+\left\|\phi_{k}-p_{k}^{0}\right\|_{\infty}\leq\varepsilon+\varepsilon=2\varepsilon.$
+where $\boldsymbol{\phi}=(\phi_{1},\phi_{2},\dots,\phi_{K})^{\top}\in\mathcal{T}(M,D,H,W_{1},B_{1},S_{1},\gamma)$ and $G\in\mathcal{F}_{\mathrm{id}}(L,W,S)$ are provided in Lemma 8 and Lemma 28, respectively.
+Here we temporarily allow the notation $\mathcal{F}_{\mathrm{id}}(L,W,S)$ to refer to scalar-valued networks with depth $L$, width $W$, and size $S$.
+It is readily seen that $\mathbf{G}\in\mathcal{F}_{\mathrm{id}}(L,KW,KS)$.
+Consequently, for any $k=1,\dots,K$, we have
+$\left\|e^{G(\phi_{k})}-p_{k}^{0}\right\|_{\infty}\leq\left\|e^{G(\phi_{k})}-\phi_{k}\right\|_{\infty}+\left\|\phi_{k}-p_{k}^{0}\right\|_{\infty}\leq 2\varepsilon.$
 Now, define the vector-valued function $\widetilde{\mathbf{q}}=\big{(}\widetilde{q}_{1},\cdots,\widetilde{q}_{K}\big{)}^{\top}$ component-wisely as
@@
 $\leq$ $\sum_{\ell=1}^{K}\left\|p_{\ell}^{0}-e^{G(\phi_{\ell})}\right\|_{\infty}\left\|\frac{e^{G(\phi_{k})}}{\sum_{j=1}^{K}e^{G(\phi_{j})}}\right\|_{\infty}+\left\|e^{G(\phi_{k})}-p_{k}^{0}\right\|_{\infty}$
-$\leq$ $2K\varepsilon+2\varepsilon=2\varepsilon(K+1).$
-Next, we investigate the low bound of $\widetilde{q}_{k}(\mathbf{Z})$. According to Lemma  27, we have $e^{G(\phi_{k})}\geq\varepsilon$. Also note that
-$\left\|\sum_{j=1}^{K}e^{G(\phi_{j}(\mathbf{Z}))}-1\right\|_{\infty}=\left\|\sum_{j=1}^{K}\{e^{G(\phi_{j}(\mathbf{Z}))}-p_{j}^{0}\}\right\|_{\infty}\leq 2K\varepsilon.$
-Consequently,
-$\widetilde{q}_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2K\varepsilon+1}\geq\frac{\varepsilon}{2K+1}\geq\frac{\varepsilon}{2(K+1)}.$
+$\leq$ $2K\varepsilon+2\varepsilon$
+$=$ $2\varepsilon(K+1).$
+Next, we investigate the low bound of $\widetilde{q}_{k}(\mathbf{Z})$. According to Lemma  28 and Remark 29, we have $\log(\varepsilon)\leq G(\phi_{k})\leq\log(2)$. Consequently,
+$\widetilde{q}_{k}(\mathbf{Z})=\frac{e^{G(\phi_{k}(\mathbf{Z}))}}{\sum_{j=1}^{K}e^{G(\phi_{j}(\mathbf{Z}))}}\geq\frac{\varepsilon}{2K}.$
 This completes the proof of Theorem 9.
-The approximation capabilities of deep ReLU networks have been extensively studied, establishing sharp rates for various function classes (Yarotsky, 2017; DeVore et al., 2021; Elbrächter et al., 2021; Berner et al., 2021; Siegel, 2023; Schmidt-Hieber, 2020; Shen et al., 2019; Lu et al., 2021; Shen et al., 2022; Jiao et al., 2023).
-A key component in our subsequent analysis is the efficient approximation of the logarithm function by a scalar-valued neural network with controlled complexity.
-To this end, we present the following lemma, which directly adapts the construction from Theorem 4.1 of Bos and Schmidt-Hieber (2022).
-Lemma 26 (Scalar approximation of $\log x$)
-For any $\varepsilon\in(0,1)$ and $\beta>0$, there exists a scalar-valued ReLU neural network $G:[0,1]\to\mathbb{R}$ such that
-$\left|e^{G(x)}-x\right|\leq\varepsilon,\quad\text{and}\quad G(x)\geq\log(\varepsilon),\quad\text{for all }x\in[0,1],$
-with network depth $L$, width $W$, and size $S$ (total number of non-zero parameters) satisfying
-$L\lesssim\log\left(\frac{1}{\varepsilon}\right),\quad W\lesssim\left(\frac{1}{\varepsilon}\right)^{\frac{1}{\beta}},\quad S\lesssim\left(\frac{1}{\varepsilon}\right)^{\frac{1}{\beta}}\log\left(\frac{1}{\varepsilon}\right).$
-Lemma 27 (Vectorized approximation via coordinate-wise application)
-Let $K\geq 1$, and let $G$ be a scalar-valued neural network with depth $L$, width $W$, and at most $S$ nonzero parameters. Define the vector-valued network
-$\mathbf{G}(\boldsymbol{\phi}(\mathbf{Z})):=\bigl{(}G(\phi_{1}(\mathbf{Z})),\dots,G(\phi_{K}(\mathbf{Z}))\bigr{)}^{\top},$
-where $\boldsymbol{\phi}(\mathbf{Z})=(\phi_{1}(\mathbf{Z}),\dots,\phi_{K}(\mathbf{Z}))^{\top}\in[0,1]^{K}$.
-Then $\mathbf{G}\in\mathcal{F}_{\mathrm{id}}(L,KW,KS)$, and for all $k\in[K]$ and $\mathbf{Z}\in[0,1]^{d}$,
-$\left|e^{G(\phi_{k}(\mathbf{Z}))}-\phi_{k}(\mathbf{Z})\right|\leq\varepsilon,\quad\text{and}\quad G(\phi_{k}(\mathbf{Z}))\geq\log(\varepsilon).$
-Proof  Although $\mathcal{F}_{\mathrm{id}}(L,W,S)$ as defined in (5) refers to vector-valued neural networks mapping $0,1]^{K}\to\mathbb{R}^{K}$, we temporarily allow the notation to refer to scalar-valued networks when the context is clear.
-To prove the lemma, it suffices to construct a scalar-valued ReLU neural network $G:[0,1]\to\mathbb{R}$ such that
-$\left|e^{G(x)}-x\right|\leq\varepsilon\quad\text{and}\quad G(x)\geq\log(\varepsilon)\quad\text{for all }x\in[0,1].$
-Given such a function $G$, we define the vector-valued network $\mathbf{G}:[0,1]^{K}\to\mathbb{R}^{K}$ by
-$\mathbf{G}(\mathbf{x}):=\bigl{(}G(x_{1}),G(x_{2}),\dots,G(x_{K})\bigr{)}^{\top}.$
-By construction, for each $k\in[K]$, the coordinate function satisfies
-$\left|e^{G(x_{k})}-x_{k}\right|\leq\varepsilon,\quad\text{and}\quad G(x_{k})\geq\log(\varepsilon),\quad\text{for all }x_{k}\in[0,1],$
-so $\mathbf{G}$ satisfies the desired approximation and positivity conditions coordinate-wise.
-Since the same network $G$ is applied independently to each coordinate, the resulting network $\mathbf{G}$ lies in the class $\mathcal{F}_{\mathrm{id}}(L,KW,KS)$, completing the proof.
+Remark 30
+According to the proof of Theorem 4.1 of Bos and Schmidt-Hieber (2022), for any ReLU neural network $G_{0}$ and $\varepsilon\in(0,1)$, we can always construct a network $G(x)$ by adding additional ReLU layers to $G_{0}$ such that $G(x)=\{G_{0}(x)\vee\log(\varepsilon)\}\wedge\log(2)$, where $a\vee b=\max\{a,b\}$ and $a\wedge b=\min\{a,b\}$.
+If $|e^{G_{0}(x)}-x|\leq\varepsilon$ for all $x\in[0,1]$, then $|e^{G(x)}-x|\leq\varepsilon$ for all $x\in[0,1]$.
+This construction also ensures $\log(\varepsilon)\leq G(x)\leq\log(2)$.
+Therefore, by applying this construction to the MLP architecture in the network of $\mathcal{F}$ and adapting the arguments in proof of Theorem 9, we can ensure that for any network $\mathbf{p}=(p_{1},\dots,p_{K})\in\mathcal{F}$,
+$p_{k}\geq\frac{\varepsilon}{2K},\quad k=1,\dots,K,$
+where $\mathcal{F}=\mathcal{F}(M,H,D,W_{1},B_{1},S_{1},\gamma,L,W,S)$ is defined in (3.4).
 Appendix C Proof of Theorem 10
-To establish an upper bound for the risk $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$ in terms of the Kullback–Leibler (KL) divergence, we first provide two Lemmas that will be used in the proof of Theorem 11.
+To establish an upper bound for the approximation error $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$ in terms of the Kullback-Leibler (KL) divergence, we first provide two Lemmas that will be used in the proof of Theorem 10.
 Lemma 31
@@
 $\int g(\mathbf{z})\,d\mathbb{P}_{\mathbf{Z}}(\mathbf{z})=\int_{0}^{\infty}\mathbb{P}_{\mathbf{Z}}\bigl{(}g(\mathbf{Z})\geq u\bigr{)}\,du.$
-Set
-$g(\mathbf{z})=\frac{1}{p_{k}(\mathbf{z})}\mathbf{1}_{\{p_{k}(\mathbf{z})\geq h\}},$
-thus
-$\int_{\{p_{k}(\mathbf{z})\geq h\}}\frac{1}{p_{k}(\mathbf{z})}\,d\mathbb{P}_{\mathbf{Z}}(\mathbf{z})=\int_{0}^{\infty}\mathbb{P}_{\mathbf{Z}}\left(\frac{1}{p_{k}(\mathbf{Z})}\mathbb{I}_{\{p_{k}(\mathbf{Z})\geq h\}}\geq u\right)\,du.$
+Write $g(\mathbf{z})=\frac{1}{p_{k}(\mathbf{z})}\mathbb{I}(p_{k}(\mathbf{z})\geq h)$, it follows that
+$\int_{\{p_{k}(\mathbf{z})\geq h\}}\frac{1}{p_{k}(\mathbf{z})}\,d\mathbb{P}_{\mathbf{Z}}(\mathbf{z})=\int_{0}^{\infty}\mathbb{P}_{\mathbf{Z}}\left(\frac{1}{p_{k}(\mathbf{Z})}\mathbb{I}(p_{k}(\mathbf{z})\geq h)\geq u\right)\,du.$
 Since
-$\frac{1}{p_{k}(\mathbf{Z})}\mathbf{1}_{\{p_{k}(\mathbf{Z})\geq h\}}\geq u>0\quad\Longrightarrow\quad h\leq p_{k}(\mathbf{Z})\leq\frac{1}{u},$
+$\frac{1}{p_{k}(\mathbf{Z})}\mathbb{I}(p_{k}(\mathbf{z})\geq h)\geq u>0\quad\Longrightarrow\quad h\leq p_{k}(\mathbf{Z})\leq\frac{1}{u},$
 it follows that $u\in(0,1/h]$ and
@@
 Lemma 32 (KL-Chi-square Divergence Inequality)
-Let $P$ and $Q$ be two probability measures satisfying $P\ll Q$ with the Radon-Nikodym derivative
-$g(z)=\frac{dP}{dQ}(z).$
-Then, the following inequality holds:
+Let $P$ and $Q$ be two probability measures satisfying $P\ll Q$ with the Radon-Nikodym derivative $g(z)=\frac{dP}{dQ}(z)$. Then, the following inequality holds:
 $\mathrm{KL}(P\|Q)\leq\chi^{2}(P\|Q),$
@@
 This completes the proof of Lemma 32.
-Proof  Recall that $R(\mathbf{p}_{0},\mathbf{p})=\mathbb{E}_{\mathbf{Z}}\left[\mathbf{p}_{0}(\mathbf{Z})^{\top}\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\mathbf{p}(\mathbf{Z})}\right].$ Applying Lemma 29 to our setting yields
+Proof [Proof of Theorem 10] Recall that $R(\mathbf{p}_{0},\mathbf{p})=\mathbb{E}_{\mathbf{Z}}\left[\mathbf{p}_{0}(\mathbf{Z})^{\top}\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\mathbf{p}(\mathbf{Z})}\right].$ Applying Lemma 32 to our setting yields
 $R(\mathbf{p}_{0},\mathbf{p})=\mathbb{E}_{\mathbf{Z}}\left[KL(\mathbf{p}_{0}(\mathbf{Z}),\mathbf{p}(\mathbf{Z}))\right]\leq\mathbb{E}_{\mathbf{Z}}\left[\sum_{k=1}^{K}\frac{(p_{k}^{0}(\mathbf{Z})-p_{k}(\mathbf{Z}))^{2}}{p_{k}(\mathbf{Z})}\right].$
 Recall that
-$\|\mathbf{p}-\mathbf{p}_{0}\|_{\infty}\leq 2(K+1)\varepsilon,\quad\text{and}\quad\min_{1\leq k\leq K}\inf_{\mathbf{Z}\in[0,1]^{d}}p_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2(K+1)}.$
+$\|\mathbf{p}-\mathbf{p}_{0}\|_{\infty}\leq 2(K+1)\varepsilon,\quad\text{and}\quad\min_{1\leq k\leq K}\inf_{\mathbf{Z}\in[0,1]^{d}}p_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2K}.$
 Therefore, for any $k\in\{1,\dots,K\}$ and any $\mathbf{Z}\in[0,1]^{d}$, it holds that
-$p_{k}(\mathbf{Z})\geq\max\left\{p_{k}^{0}(\mathbf{Z})-2(K+1)\varepsilon,\;\frac{\varepsilon}{2(K+1)}\right\}.$
+$p_{k}(\mathbf{Z})\geq\max\left\{p_{k}^{0}(\mathbf{Z})-2(K+1)\varepsilon,\;\frac{\varepsilon}{2K}\right\}.$
 Furthermore, if
-$p_{k}^{0}(\mathbf{Z})-2(K+1)\varepsilon\geq\frac{\varepsilon}{2(K+1)},$
+$p_{k}^{0}(\mathbf{Z})-2(K+1)\varepsilon\geq\frac{\varepsilon}{2K},$
 then we have
-$p_{k}^{0}(\mathbf{Z})-2(K+1)\varepsilon\geq p_{k}^{0}(\mathbf{Z})\left(1-\frac{2(K+1)}{2(K+1)+\frac{1}{2(K+1)}}\right)=\frac{p_{k}^{0}(\mathbf{Z})}{1+4(K+1)^{2}}.$
+$p_{k}^{0}(\mathbf{Z})-2(K+1)\varepsilon\geq p_{k}^{0}(\mathbf{Z})\left(1-\frac{2(K+1)}{2(K+1)+\frac{1}{2K}}\right)=\frac{p_{k}^{0}(\mathbf{Z})}{1+4K(K+1)}.$
 This gives rise to the following upper bound:
-$\begin{split}\frac{\left(p_{k}^{0}(\mathbf{Z})-p_{k}(\mathbf{Z})\right)^{2}}{p_{k}(\mathbf{Z})}&\leq 8(K+1)^{3}\varepsilon\cdot\mathbb{I}\left\{p_{k}^{0}(\mathbf{Z})\leq 2(K+1)\varepsilon+\frac{\varepsilon}{2(K+1)}\right\}\\ &+\frac{4(K+1)^{2}\left[1+4(K+1)^{2}\right]\varepsilon^{2}}{p_{k}^{0}(\mathbf{Z})}\cdot\mathbb{I}\left\{p_{k}^{0}(\mathbf{Z})>2(K+1)\varepsilon+\frac{\varepsilon}{2(K+1)}\right\}.\end{split}$
+$\begin{split}\frac{\left(p_{k}^{0}(\mathbf{Z})-p_{k}(\mathbf{Z})\right)^{2}}{p_{k}(\mathbf{Z})}&\leq 8K(K+1)^{2}\varepsilon\cdot\mathbb{I}\left\{p_{k}^{0}(\mathbf{Z})\leq 2(K+1)\varepsilon+\frac{\varepsilon}{2K}\right\}\\ &\quad+\frac{4(K+1)^{2}\left[1+4K(K+1)\right]\varepsilon^{2}}{p_{k}^{0}(\mathbf{Z})}\cdot\mathbb{I}\left\{p_{k}^{0}(\mathbf{Z})>2(K+1)\varepsilon+\frac{\varepsilon}{2K}\right\}.\end{split}$
 Taking the expectation on both sides with respect to $\mathbf{Z}$, we obtain
-$\mathbb{E}_{\mathbf{Z}}\left[\frac{\left(p_{k}^{0}(\mathbf{Z})-p_{k}(\mathbf{Z})\right)^{2}}{p_{k}(\mathbf{Z})}\right]\leq 8(K+1)^{3}\varepsilon\,\mathbb{P}\left\{p_{k}^{0}(\mathbf{Z})\leq 2(K+1)\varepsilon+\frac{\varepsilon}{2(K+1)}\right\}$
-$+4(K+1)^{2}\left[1+4(K+1)^{2}\right]\varepsilon^{2}\,\mathbb{E}_{\mathbf{Z}}\left[\frac{\mathbb{I}\left\{p_{k}^{0}(\mathbf{Z})>2(K+1)\varepsilon+\frac{\varepsilon}{2(K+1)}\right\}}{p_{k}^{0}(\mathbf{Z})}\right].$
+$\mathbb{E}_{\mathbf{Z}}\left[\frac{\left(p_{k}^{0}(\mathbf{Z})-p_{k}(\mathbf{Z})\right)^{2}}{p_{k}(\mathbf{Z})}\right]\leq 8K(K+1)^{2}\varepsilon\,\mathbb{P}\left\{p_{k}^{0}(\mathbf{Z})\leq 2(K+1)\varepsilon+\frac{\varepsilon}{2K}\right\}$
+$\quad+4(K+1)^{2}\left[1+4K(K+1)\right]\varepsilon^{2}\,\mathbb{E}_{\mathbf{Z}}\left[\frac{\mathbb{I}\left\{p_{k}^{0}(\mathbf{Z})>2(K+1)\varepsilon+\frac{\varepsilon}{2K}\right\}}{p_{k}^{0}(\mathbf{Z})}\right].$
 By applying the $\alpha$-SVB condition to bound the first term and using Lemma 31 to bound the second term, we obtain
 $\mathbb{E}_{\mathbf{Z}}\left[\frac{\left(p_{k}^{0}(\mathbf{Z})-p_{k}(\mathbf{Z})\right)^{2}}{p_{k}(\mathbf{Z})}\right]$
-$\leq$ $C\,\varepsilon^{1+(\alpha\wedge 1)}\,8(K+1)^{3}\left(\frac{4(K+1)^{2}+1}{2(K+1)}\right)^{\alpha\wedge 1}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right),$
-where $\alpha\wedge\beta=\min\{\alpha,\beta\}$. Also note that $0\leq\alpha\leq 1$ and $\frac{4(K+1)^{2}+1}{2(K+1)}\leq 3(K+1)$. Consequently,
+$\leq$ $C\,\varepsilon^{1+(\alpha\wedge 1)}\,8K(K+1)^{2}\left(\frac{4K(K+1)+1}{2K}\right)^{\alpha\wedge 1}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right),$
+where $\alpha\wedge\beta=\min\{\alpha,\beta\}$. Also note that $0\leq\alpha\leq 1$ and $\frac{4K(K+1)+1}{2K}\leq 3(K+1)$. Consequently,
 $R(\mathbf{p}_{0},\mathbf{p})$ $=\mathbb{E}_{\mathbf{Z}}\left[\sum_{k=1}^{K}\frac{\left(p_{k}^{0}(\mathbf{Z})-p_{k}(\mathbf{Z})\right)^{2}}{p_{k}(\mathbf{Z})}\right]$
-$\leq C\varepsilon^{1+(\alpha\wedge 1)}K(K+1)^{3}\left(\frac{4(K+1)^{2}+1}{2(K+1)}\right)^{\alpha\wedge 1}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\!\left(\frac{1}{\varepsilon}\right)\right)$
-$\leq C\varepsilon^{1+(\alpha\wedge 1)}K(K+1)^{3+(\alpha\wedge 1)}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right)$
-$=C\varepsilon^{1+\alpha}K(K+1)^{3+\alpha}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right)$
+$\leq C\varepsilon^{1+(\alpha\wedge 1)}K^{2}(K+1)^{2}\left(\frac{4K(K+1)+1}{2K}\right)^{\alpha\wedge 1}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\!\left(\frac{1}{\varepsilon}\right)\right)$
+$\leq C\varepsilon^{1+(\alpha\wedge 1)}K^{2}(K+1)^{2+(\alpha\wedge 1)}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right)$
+$=C\varepsilon^{1+\alpha}K^{2}(K+1)^{2+\alpha}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right)$
 $\leq C\varepsilon^{1+\alpha}K^{4+\alpha}\left(2+\frac{I(\alpha<1)}{1-\alpha}+\log\left(\frac{1}{\varepsilon}\right)\right).$
@@
 $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})\lesssim T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}K^{4+\alpha}\log(KT).$
-This concludes the proof of Theorem  11.
+This concludes the proof of Theorem 10.
 Appendix D Proof of Lemma 12
-Proof  We first establish a useful Lipschitz continuity property of the log-softmax function. This intermediate result is critical for bounding covering numbers in Lemma 12.
+Proof [Proof of Lemma 12] We first establish a useful Lipschitz continuity property of the log-softmax function. This intermediate result is critical for bounding covering numbers in Lemma 12.
 Lemma 33 (Lipschitz Continuity of Log-Softmax)
-Let $\mathbf{y}=(y_{1},\dots,y_{K})\in\mathbb{R}^{K}$, and define the softmax function $\mathbf{\Phi}:\mathbb{R}^{K}\to\Delta^{K-1}$ by
+Let $\mathbf{y}=(y_{1},\dots,y_{K})^{\top}\in\mathbb{R}^{K}$, and define the softmax function $\mathbf{\Phi}:\mathbb{R}^{K}\to\mathcal{S}^{K}$ by
 $\mathbf{\Phi}(\mathbf{y}):=\frac{1}{\sum_{j=1}^{K}\exp(y_{j})}\begin{bmatrix}\exp(y_{1})\\ \vdots\\ \exp(y_{K})\end{bmatrix}$,
-where $\Delta^{K-1}$ is the probability simplex in $\mathbb{R}^{K}$. Then $\log\circ\mathbf{\Phi}$ is $K$-Lipschitz with respect to the $\|\cdot\|_{\infty}$-norm, i.e.,
+where $\mathcal{S}^{K}$ is the probability simplex in $\mathbb{R}^{K}$. Then $\log\circ\mathbf{\Phi}$ is $K$-Lipschitz with respect to the $\|\cdot\|_{\infty}$-norm, i.e.,
 $\|\log(\mathbf{\Phi}(\mathbf{y}_{1}))-\log(\mathbf{\Phi}(\mathbf{y}_{2}))\|_{\infty}\leq K\|\mathbf{y}_{1}-\mathbf{y}_{2}\|_{\infty},\quad\forall\,\mathbf{y}_{1},\mathbf{y}_{2}\in\mathbb{R}^{K}.$
-Proof  Consider the log-softmax function
-$\log(\mathbf{\Phi}(\mathbf{y}))_{i}=\mathbf{y}_{i}-\log\left(\sum_{j=1}^{K}e^{\mathbf{y}_{j}}\right).$
-Taking derivatives, for $i,k\in\{1,\dots,K\}$, we have
-$\frac{\partial}{\partial\mathbf{y}_{k}}\log(\mathbf{\Phi}(\mathbf{y}))_{i}=\delta_{ik}-\mathbf{\Phi}(\mathbf{y})_{k},$
+Proof  Consider the log-softmax function $\log(\mathbf{\Phi}(\mathbf{y}))_{i}=y_{i}-\log\left(\sum_{j=1}^{K}e^{y_{j}}\right)$. Taking derivatives, we have for $i,k\in\{1,\dots,K\}$,
+$\frac{\partial}{\partial y_{k}}\log(\mathbf{\Phi}(\mathbf{y}))_{i}=\delta_{ik}-\mathbf{\Phi}(\mathbf{y})_{k},$
 where $\delta_{ik}$ is the Kronecker delta.
 As $0\leq\mathbf{\Phi}(\mathbf{y})_{k}\leq 1$, the Jacobian’s entries $\delta_{ik}-\mathbf{\Phi}(\mathbf{y})_{k}$ are bounded by 1 in absolute value.
-Thus, the Lipschitz constant under the $\ell_{\infty}$-norm is at most $K$.
+Thus, the Lipschitz constant is at most $K$.
 We now define an auxiliary function class excluding the softmax layer:
@@
 This allows us to represent the original function class conveniently as
-$\log\mathcal{F}=\log(\mathbf{\Phi}(\mathcal{F}^{\prime})).$
-The next lemma establish the covering number bound of $log\mathcal{F}$ with respect to the norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$.
+$\log(\mathcal{F})=\log(\mathbf{\Phi}(\mathcal{F}^{\prime})).$
+The next lemma establish the covering number bound of $\log\mathcal{F}$ with respect to the norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$.
 Lemma 34
 Given a sample $\mathcal{Z}_{T}=\{\mathbf{Z}^{(1)},\dots,\mathbf{Z}^{(T)}\}\subset[0,1]^{d}$, then for any $\delta>0$, we have
-$\mathcal{N}\left(\delta,\log\mathcal{F},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\mathcal{N}\left(\tfrac{\delta}{2K},\mathcal{F}^{\prime},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right).$
+$\mathcal{N}\left(\delta,\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\mathcal{N}\left(\tfrac{\delta}{2K},\mathcal{F}^{\prime},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right).$
 Proof  Fix $\delta>0$.
-Let $\{\mathbf{f}_{j}\}_{j=1}^{J}$ be the centers of a minimal $\delta/(2K)$-cover of $\mathcal{F}^{\prime}$ under the empirical sup-norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$.
+Let $\{\mathbf{f}_{j}\}_{j=1}^{J}$ be the centers of a minimal $\delta/(2K)$-cover of $\mathcal{F}^{\prime}$ with respect to $\|\cdot\|_{\mathcal{Z}_{T}}$.
 That is, for every $\mathbf{f}\in\mathcal{F}^{\prime}$, there exists $j\in\{1,\dots,J\}$ such that
@@
 Lemma 35 (Empirical Covering Bound for Composed Class)
-Let $\mathcal{T}$ be a class of functions from $0,1]^{d}$ to $\mathbb{R}^{K}$, and let $\mathcal{F}_{\mathrm{id}}$ be a class of functions from $\mathbb{R}^{K}$ to $\mathbb{R}^{K}$, such that every $\mathbf{f}\in\mathcal{F}_{\mathrm{id}}$ is $L_{1}$-Lipschitz with respect to the sup-norm $\|\cdot\|_{\infty}$.
+Let $\mathcal{T}$ be a class of functions from $0,1]^{d}$ to $\mathbb{R}^{K}$, and let $\mathcal{F}_{\mathrm{id}}$ be a class of functions from $\mathbb{R}^{K}$ to $\mathbb{R}^{K}$, such that every $\mathbf{f}\in\mathcal{F}_{\mathrm{id}}$ is $L_{1}$-Lipschitz with respect to $\|\cdot\|_{\infty}$.
 Define the composed function class
@@
 Proof  Fix $\delta>0$.
-Let $\widehat{\mathcal{T}}$ be a $\tfrac{\delta}{4KL_{1}}$-cover of $\mathcal{T}$ with respect to the empirical sup-norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$, and let $\widehat{\mathcal{F}}_{\mathrm{id}}$ be a $\tfrac{\delta}{4K}$-cover of $\mathcal{F}_{\mathrm{id}}$ with respect to the sup-norm $\|\cdot\|_{\infty}$.
+Let $\widehat{\mathcal{T}}$ be a $\tfrac{\delta}{4KL_{1}}$-cover of $\mathcal{T}$ with respect to $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$, and let $\widehat{\mathcal{F}}_{\mathrm{id}}$ be a $\tfrac{\delta}{4K}$-cover of $\mathcal{F}_{\mathrm{id}}$ with respect to $\|\cdot\|_{\infty}$.
 Define the composite approximation set
 $\widehat{\mathcal{F}^{\prime}}:=\left\{\hat{\mathbf{f}}\circ\hat{\boldsymbol{\phi}}:\hat{\mathbf{f}}\in\widehat{\mathcal{F}}_{\mathrm{id}},\;\hat{\boldsymbol{\phi}}\in\widehat{\mathcal{T}}\right\}$.
-Next we show that $\widehat{\mathcal{F}^{\prime}}$ is a $\tfrac{\delta}{2K}$-cover of $\mathcal{F}^{\prime}$ with respect to the empirical sup-norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$.
-Indeed, for any $\mathbf{f}\in\mathcal{F}_{\mathrm{id}}$ and $\boldsymbol{\phi}\in\mathcal{T}$, there exist $\hat{\mathbf{f}}\in\widehat{\mathcal{F}}_{\mathrm{id}}$ and $\hat{\boldsymbol{\phi}}\in\widehat{\mathcal{T}}$ such that
-$\|\boldsymbol{\phi}-\hat{\boldsymbol{\phi}}\|_{\infty,\mathcal{Z}_{T}}\leq\tfrac{\delta}{4KL_{1}}\quad\text{and}\quad\|\mathbf{f}-\hat{\mathbf{f}}\|_{\infty}\leq\tfrac{\delta}{4K}.$
-Then for each $t\in\{1,\ldots,T\}$, we have
+Next, we show that $\widehat{\mathcal{F}^{\prime}}$ is a $\tfrac{\delta}{2K}$-cover of $\mathcal{F}^{\prime}$ with respect to $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$.
+Indeed, for any $\mathbf{f}\in\mathcal{F}_{\mathrm{id}}$ and $\boldsymbol{\phi}\in\mathcal{T}$, there exist $\hat{\mathbf{f}}\in\widehat{\mathcal{F}}_{\mathrm{id}}$ and $\hat{\boldsymbol{\phi}}\in\widehat{\mathcal{T}}$ such that Then for each $t\in\{1,\ldots,T\}$, we have
 $\|\mathbf{f}(\boldsymbol{\phi}(\mathbf{Z}^{(t)}))-\hat{\mathbf{f}}(\hat{\boldsymbol{\phi}}(\mathbf{Z}^{(t)}))\|_{\infty}$ $\leq\|\mathbf{f}(\boldsymbol{\phi}(\mathbf{Z}^{(t)}))-\mathbf{f}(\hat{\boldsymbol{\phi}}(\mathbf{Z}^{(t)}))\|_{\infty}+\|\mathbf{f}(\hat{\boldsymbol{\phi}}(\mathbf{Z}^{(t)}))-\hat{\mathbf{f}}(\hat{\boldsymbol{\phi}}(\mathbf{Z}^{(t)}))\|_{\infty}$
 $\leq L_{1}\|\boldsymbol{\phi}(\mathbf{Z}^{(t)})-\hat{\boldsymbol{\phi}}(\mathbf{Z}^{(t)})\|_{\infty}+\tfrac{\delta}{4K}$
-$\leq L_{1}\tfrac{\delta}{4KL_{1}}+\tfrac{\delta}{4K}=\tfrac{\delta}{2K}.$
+$\leq L_{1}\tfrac{\delta}{4KL_{1}}+\tfrac{\delta}{4K}$
+$=\tfrac{\delta}{2K}.$
 Taking the maximum over $t=1,\ldots,T$ yields
@@
 Finally, by Lemma 34 and Lemma 35, we obtain
-$\mathcal{N}\left(\delta,\log\mathcal{F},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)$ $\leq$ $\mathcal{N}\left(\tfrac{\delta}{2K},\mathcal{F}^{\prime},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)$
+$\mathcal{N}\left(\delta,\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)$ $\leq$ $\mathcal{N}\left(\tfrac{\delta}{2K},\mathcal{F}^{\prime},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)$
 $\leq$ $\mathcal{N}\left(\tfrac{\delta}{4KL_{1}},\mathcal{T},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\cdot\mathcal{N}\left(\tfrac{\delta}{4K},\mathcal{F}_{\mathrm{id}},\|\cdot\|_{\infty}\right).$
@@
 Appendix E Proof of Lemma 13
-Before presenting the detailed proof of Lemma 13, we first introduce the notion of pseudo-dimension which will be used in the proof of Lemma  13.
+Before presenting the detailed proof of Lemma 13, we first introduce the notion of pseudo-dimension which will be used in the proof of Lemma 13.
 Definition 36
@@
 (2019).
-The main idea is to bound the pseudo-dimension of subsets of $\mathcal{T}$ and then utilize well-known results developed by Anthony and Bartlett (2009) to bound the covering entropy by pseudo-dimensions.
+The main idea is to bound the pseudo-dimension of subsets of $\mathcal{T}$ and then utilize well-known results developed by Anthony and Bartlett (2009) to bound the entropy by pseudo-dimensions.
 Here $\mathcal{T}=\mathcal{T}(M,H,D,W_{1},B_{1},S_{1},\gamma)$ defined in (3.3) in the main text.
@@
 Since $\sup_{\mathbf{Z}}|g(\mathbf{Z},\mathbf{\theta})|\leq B_{1}$, it follows that
-$\mathcal{N}(\delta_{1},\mathcal{G},\|\cdot\|_{\infty,\mathcal{Z}_{T}})\leq\mathcal{N}(\delta_{1},\mathcal{B}_{B_{1}},\|\cdot\|_{\infty})\leq\left(\frac{3B_{1}}{\delta_{1}}\right)^{T}\leq\left(\frac{3B_{1}}{\delta_{1}}\right)^{{\rm Pdim}(\mathcal{G})},\ \text{for}\ T<{\rm Pdim}(\mathcal{G}).$
+$\mathcal{N}(\delta_{1},\mathcal{G},\|\cdot\|_{\infty,\mathcal{Z}_{T}})\leq\mathcal{N}(\delta_{1},\mathcal{B}_{B_{1}},\|\cdot\|_{\infty})\leq\left(\tfrac{3B_{1}}{\delta_{1}}\right)^{T}\leq\left(\tfrac{3B_{1}}{\delta_{1}}\right)^{\mathrm{Pdim}(\mathcal{G})},\qquad\,T<\mathrm{Pdim}(\mathcal{G})\,.$
 Therefore, in all cases, we have
@@
 This complete the proof of Lemma 13.
-In the end of this subsection, we present some lemmas that are used in the proof of Lemma  13.
+In the end of this subsection, we present some lemmas that are used in the proof of Lemma 13.
 Lemma 37 (Theorem 8.3 of Anthony and Bartlett (2009))
@@
 Lemma 39 (Theorem 12.2 of Anthony and Bartlett (2009))
-Let $\mathcal{F}$ be a class of real-valued functions defined on a domain $\mathcal{X}$ with range bounded in the interval $[0,B]$.
-Let $\epsilon>0$, and suppose that the pseudo-dimension of $\mathcal{F}$ is $d$.
-Then the $\ell_{\infty}$ covering number satisfies
+Let $\mathcal{F}$ denote a class of real-valued functions on a domain $\mathcal{X}$, with range in $[0,B]$. Let $\epsilon>0$, and assume $\mathcal{F}$ has pseudo-dimension $d$. Then
 $\mathcal{N}_{\infty}(\epsilon,\mathcal{F},m)\leq\sum_{i=1}^{d}\binom{m}{i}\left(\frac{B}{\epsilon}\right)^{i},$
-which is further bounded by
-$\left(\frac{emB}{\epsilon d}\right)^{d}\quad\text{for }m\geq d.$
+which is further bounded by $\left(\frac{emB}{\epsilon d}\right)^{d}$ for $m\geq d$.
 Appendix F Proof of Lemma 14
@@
 $=$ $\left\|\sum_{k=1}^{L}\mathcal{B}_{k+1}(\mathbf{g})\circ(W^{(k)}\cdot+b^{(k)})\circ\mathcal{A}_{k}(\mathbf{f})(\mathbf{x})-\mathcal{B}_{k+1}(\mathbf{g})\circ({W^{{}^{\prime}(k)}}\cdot+{b^{{}^{\prime}(k)}})\circ\mathcal{A}_{k}(\mathbf{f})(\mathbf{x})\right\|_{\infty}$
-$\leq$ $\sum_{k=1}^{L}\|\mathcal{B}_{k+1}(\mathbf{g})\|_{Lip}\|(W^{(k)}\cdot+b^{(k)})\circ\mathcal{A}_{k}(\mathbf{f})(\mathbf{x})-({W^{{}^{\prime}(k)}}\cdot+{b^{{}^{\prime}(k)}})\circ\mathcal{A}_{k}(\mathbf{f})(\mathbf{x})\|_{\infty}.$
-Next, we explicitly derive the following uniform bound for each intermediate mapping. According the definition of $\mathcal{F}_{\rm id}(L,W,S)$ given in (5), we have
+$\leq$ $\sum_{k=1}^{L}\|\mathcal{B}_{k+1}(\mathbf{g})\|_{Lip}\cdot\|(W^{(k)}\cdot+b^{(k)})\circ\mathcal{A}_{k}(\mathbf{f})(\mathbf{x})-({W^{{}^{\prime}(k)}}\cdot+{b^{{}^{\prime}(k)}})\circ\mathcal{A}_{k}(\mathbf{f})(\mathbf{x})\|_{\infty},$
+where $\|\mathcal{B}_{k+1}(\mathbf{g})\|_{Lip}$ denotes the corresponding Lipschitz constant of $\mathcal{B}_{k+1}(\mathbf{g})$.
+Next, we explicitly derive the following uniform bound for each of the intermediate mappings.
+According to the definition of $\mathcal{F}_{\rm id}(L,W,S)$ provided in (5), we obtain
 $\|\mathcal{A}_{k}(\mathbf{f})\|_{\infty}$ $\leq$ $\max_{j}\|W^{(k-1)}_{j,:}\|_{1}\|\mathcal{A}_{k-1}(\mathbf{f})\|_{\infty}+\|b^{(k-1)}\|_{\infty}$
@@
 $\|\mathbf{f}(\mathbf{y})-\mathbf{g}(\mathbf{y})\|_{\infty}\leq\sum_{k=1}^{L}W^{L-k}[W(W+1)^{k-1}+1]\delta_{2}\leq L(W+1)^{L}\delta_{2}.$
-Similar to the arguments for Lemma  13, we fix the positions of these nonzero parameters of the network $F_{id}(L,W,S)$ and then its covering number with respect to the $\|\cdot\|_{\infty}$-norm is bounded by
+Similar to the arguments for Lemma 13 in Appendix E, we fix the positions of these nonzero parameters of the network $F_{id}(L,W,S)$ and then its covering number with respect to the $\|\cdot\|_{\infty}$-norm is bounded by
 $\left(\frac{3\delta_{2}}{L(W+1)^{L}}\right)^{-S}.$
@@
 Appendix G Proof of Lemma 16
-The Kullback–Leibler divergence and the Hellinger distance are two classical measures of discrepancy between probability distributions; see Tsybakov (2008, Section 2.4) and Birgé and Massart (1998, Lemma 5).
+The Kullback-Leibler divergence and Hellinger distance are two classical measures of discrepancy between probability distributions; see Tsybakov (2008, Section 2.4) and Birgé and Massart (1998, Lemma 5).
 These two measures are closely related and satisfy several useful inequalities, which are fundamental in nonparametric estimation theory.
-Recall that the squared Hellinger distance between $P$ and $Q$ is defined as
-$H^{2}(P,Q)=\left(1-\int\sqrt{dP\,dQ}\right),$
-as given in Birgé and Massart (1998, Section 7.3).
+Recall that the squared Hellinger distance between $P$ and $Q$ is defined as $H^{2}(P,Q)=1-\int\sqrt{dP\,dQ}$ (Birgé and Massart, 1998, Section 7.3).
 We proceed to establish the inequality stated in Lemma 16.
-Proof [Proof of Lemma 16] We first show that $H^{2}(P,Q)\leq\frac{1}{2}\mathrm{KL}_{2}(P,Q)\leq\frac{1}{2}\mathrm{KL}_{B}(P,Q)$.
-As the second inequality follows from the definition of the truncated KL-divergence, we only show $H^{2}(P,Q)\leq\frac{1}{2}\mathrm{KL}_{2}(P,Q)$.
+Proof [Proof of Lemma 16] We first show that $H^{2}(P,Q)\leq\frac{1}{2}\mathrm{KL}_{2}(P\|Q)\leq\frac{1}{2}\mathrm{KL}_{B}(P\|Q)$.
+As the second inequality follows from the definition of the truncated KL-divergence, we only show $H^{2}(P,Q)\leq\frac{1}{2}\mathrm{KL}_{2}(P\|Q)$.
 Let $P=P^{a}+P^{s}$ be the Lebesgue decomposition of $P$ with respect to $Q$, where $P^{a}\ll Q$ and $P^{s}\perp Q$.
@@
 Since $x\leq-\log(1+x)$ for x>1x>-1, it follows that
-$\mathrm{KL}_{2}(P,Q)$ $=$ $\int\left(2\wedge\log\frac{dP}{dQ}\right)dP$
+$\mathrm{KL}_{2}(P\|Q)$ $=$ $\int\left(2\wedge\log\frac{dP}{dQ}\right)dP$
 $=$ $\int\left(2\wedge 2\log\sqrt{p/q}\right)pd\mu$
@@
 $=$ $2H^{2}(P,Q).$
-Thus, we obtain $H^{2}(P,Q)\leq\frac{1}{2}KL_{2}(P,Q)$.
-The proof of the upper bound on $\mathrm{KL}_{B}(P,Q)$ in terms of $H^{2}(P,Q)$ follows a similar approach to that of Theorem 5 of Wong and Shen (1995) and Lemma 3.4 of Bos and Schmidt-Hieber (2022).
+Therefore, we obtain $H^{2}(P,Q)\leq\frac{1}{2}KL_{2}(P\|Q)$.
+The proof of the upper bound on $\mathrm{KL}_{B}(P\|Q)$ in terms of $H^{2}(P,Q)$ follows a similar approach to that of Theorem 5 of Wong and Shen (1995) and Lemma 3.4 of Bos and Schmidt-Hieber (2022).
 For completeness, we also provide a detailed derivation.
@@
 This implies
-$\mathrm{KL}_{B}(P,Q)$ $=$ $\int_{p^{a}/q\leq e^{B}}\frac{p^{a}}{q}\log\left(\frac{p^{a}}{q}\right)qd\mu+B\int_{\frac{dP}{dQ}>e^{B}}dP$
+$\mathrm{KL}_{B}(P\|Q)$ $=$ $\int_{p^{a}/q\leq e^{B}}\frac{p^{a}}{q}\log\left(\frac{p^{a}}{q}\right)qd\mu+B\int_{\frac{dP}{dQ}>e^{B}}dP$
 $\leq$ $2\int_{p^{a}/q\leq e^{B}}(p^{a}-\sqrt{p^{a}q})d\mu+e^{B/2}\int_{p^{a}/q\leq e^{B}}(\sqrt{p^{a}}-\sqrt{q})^{2}d\mu+B\int_{\frac{dP}{dQ}>e^{B}}dP$
@@
 Consequently,
-$\mathrm{KL}_{B}(P,Q)\leq 2\left(e^{B/2}+B(1-e^{-B/2})^{-2}\right)H^{2}(P,Q).$
+$\mathrm{KL}_{B}(P\|Q)\leq 2\left(e^{B/2}+B(1-e^{-B/2})^{-2}\right)H^{2}(P,Q).$
 Case 2: $\int_{p^{a}/q\leq e^{B}}(p^{a}-\sqrt{p^{a}q})d\mu>0$. Some elementary calculations show that
@@
 It follows that
-$\mathrm{KL}_{B}(P,Q)\leq$ $\ 2\int_{p^{a}/q\leq e^{B}}(p^{a}-\sqrt{p^{a}q})d\mu+2e^{B/2}H^{2}(P,Q)+B\int_{\frac{dP}{dQ}>e^{B}}dP$
+$\mathrm{KL}_{B}(P\|Q)\leq$ $\ 2\int_{p^{a}/q\leq e^{B}}(p^{a}-\sqrt{p^{a}q})d\mu+2e^{B/2}H^{2}(P,Q)+B\int_{\frac{dP}{dQ}>e^{B}}dP$
 $\leq$ $\ 2e^{B/2}H^{2}(P,Q)+B(1-e^{-B/2})^{-1}\int_{p^{a}/q\leq e^{B}}(p-\sqrt{pq})d\mu$
@@
 $B(1-e^{-B/2})^{-1}\leq 2e^{B/2},\quad B(1-e^{-B/2})^{-2}\leq 4e^{B/2}.$
-Thus, we obtain $\mathrm{KL}_{B}(P,Q)\leq 10e^{B/2}H^{2}(P,Q)$. Consequently,
-$\frac{1}{2}\mathrm{KL}_{B}(P,Q)\leq 5e^{B/2}H^{2}(P,Q).$
+Therefore, we obtain $\mathrm{KL}_{B}(P\|Q)\leq 10e^{B/2}H^{2}(P,Q)$. Consequently,
+$\frac{1}{2}\mathrm{KL}_{B}(P\|Q)\leq 5e^{B/2}H^{2}(P,Q).$
 This completes the proof of Lemma 16.
 Appendix H Proof of Lemma 21
-Proof [Proof of Lemma 19] Let $\mathcal{F}_{\mathrm{MLP}}(L,W,S)$ be defined as in (9), where $\boldsymbol{\Phi}$ denotes the softmax mapping.
-For any empirical input set $\mathcal{Z}_{T}\subset[0,1]^{d}$, we aim to bound the covering number of the log-transformed class $\log\big{(}\mathcal{F}_{\mathrm{MLP}}(L,W,S)\big{)}$ under the empirical sup-norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$.
-By Lemma 31, for any $\delta>0$, we have
-$\mathcal{N}\left(\delta,\log\left(\mathcal{F}_{\mathrm{MLP}}(L,W,S)\right),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\mathcal{N}\left(\tfrac{\delta}{2K},\mathcal{F}_{\mathrm{id}}(L,W,S),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right).$
-Setting $\delta=\tfrac{1}{T}$, it follows that
-$\mathcal{N}\left(\tfrac{1}{T},\log\left(\mathcal{F}_{\mathrm{MLP}}(L,W,S)\right),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\mathcal{N}\left(\tfrac{1}{2KT},\mathcal{F}_{\mathrm{id}}(L,W,S),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right).$
-Applying the standard pseudo-dimension-based covering bound (see Lemma 36, Theorem 12.2 of Anthony and Bartlett (2009)), we obtain
-$\mathcal{N}\left(\tfrac{1}{2KT},\mathcal{F}_{\mathrm{id}}(L,W,S),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\left(\frac{2eKC\log T\cdot T^{2}}{\operatorname{Pdim}(\mathcal{F}_{\mathrm{id}})}\right)^{\operatorname{Pdim}(\mathcal{F}_{\mathrm{id}})},$
-where the output range of $\mathcal{F}_{\mathrm{id}}$ is bounded by $C\log T$ for some universal constant $C>0$.
+Proof [Proof of Lemma 21] Recall that $\mathcal{F}_{\rm mlp}(L,W,S)=\mathbf{\Phi}(\mathcal{F}_{\mathrm{id}}(L,W,S))$, it follows from Lemma 34 that
+$\mathcal{N}\left(\frac{1}{T},\log\left(\mathcal{F}_{\mathrm{mlp}}(L,W,S)\right),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\mathcal{N}\left(\frac{1}{2KT},\mathcal{F}_{\mathrm{id}}(L,W,S),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right).$
+Consider the function classes
+${\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)}^{+}$ $=$ $\{\sigma(\mathbf{e}_{k}^{\top}\mathbf{G}):\mathbf{G}=(G_{1},\dots,G_{K})^{\top}\in\mathcal{F}_{\mathrm{id}}(L,W,S)\},$
+${\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)}^{-}$ $=$ $\{\sigma(-\mathbf{e}_{k}^{\top}\mathbf{G}):\mathbf{G}=(G_{1},\dots,G_{K})^{\top}\in\mathcal{F}_{\mathrm{id}}(L,W,S)\},$
+where $\sigma(x)=\max(x,0)$ is the ReLU activation function and $\mathbf{e}_{k}=(0,\dots,0,1,0,\dots,0)^{\top}\in\mathbb{R}^{K}$ with $1$ lying in the $k$-th component of $\mathbf{e}_{k}$.
+It is readily seen that both $\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)$$^{+}$ and $\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)$$^{-}$ consist of ReLU networks with depth $L+1$, width $W$, and size $S+1$.
+Since $\|\mathbf{G}\|_{\infty}\leq B_{2}$ for any $\mathbf{G}\in\mathcal{F}_{\mathrm{id}}(L,W,S)$, it follows that $|\sigma(\mathbf{e}_{k}^{\top}\mathbf{G})|\leq B_{2}$ and $|\sigma(-\mathbf{e}_{k}^{\top}\mathbf{G})|\leq B_{2}$.
+Therefore, by Lemma 39 (Theorem 12.2 of Anthony and Bartlett (2009)), we obtain
+$\mathcal{N}\left(\frac{1}{4KT},{\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)}^{+},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)$ $\leq$ $\left(\frac{4eKB_{2}T^{2}}{\mathrm{Pdim}({\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)}^{+})}\right)^{\mathrm{Pdim}({\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)}^{+})},$
 Furthermore, Theorems 3 and 6 of Bartlett et al. (2019) imply that there exist universal constants $\underline{c},\overline{c}>0$ such that
-$\underline{c}\cdot SL\log(S/L)\leq\operatorname{Pdim}(\mathcal{F}_{\mathrm{id}}(L,W,S))\leq\overline{c}\cdot SL\log S.$
+$\underline{c}(S+1)(L+1)\log\frac{S+1}{L+1}\leq\mathrm{Pdim}({\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)}^{+})\leq\overline{c}(S+1)(L+1)\log(S+1).$
+Consequently,
+$\mathcal{N}\left(\frac{1}{4KT},{\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)}^{+},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\left(\frac{4eKB_{2}T^{2}}{\underline{c}(S+1)(L+1)\log\frac{S+1}{L+1}}\right)^{\overline{c}(S+1)(L+1)\log(S+1)},$
+Similarly, we have
+$\mathcal{N}\left(\frac{1}{4KT},{\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)}^{-},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\left(\frac{4eKB_{2}T^{2}}{\underline{c}(S+1)(L+1)\log\frac{S+1}{L+1}}\right)^{\overline{c}(S+1)(L+1)\log(S+1)},$
+Also note that $G_{k}=\sigma(\mathbf{e}_{k}^{\top}\mathbf{G})-\sigma(-\mathbf{e}_{k}^{\top}\mathbf{G})$, it follows that
+$\mathcal{N}\left(\frac{1}{2KT},{\mathcal{F}_{\mathrm{id}}(L,W,S)},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\left(\frac{4eKB_{2}T^{2}}{\underline{c}(S+1)(L+1)\log\frac{S+1}{L+1}}\right)^{2\overline{c}K(S+1)(L+1)\log(S+1)}.$
 Taking the logarithm and the supremum over all $\mathcal{Z}_{T}\subset[0,1]^{d}$, we conclude that
-$\mathcal{V}_{\infty}\left(\tfrac{1}{T},\log\left(\mathcal{F}_{\mathrm{MLP}}(L,W,S)\right),T\right)\leq C\cdot SL\log S\cdot\log T,$
+$\mathcal{V}_{\infty}\left(\frac{1}{T},\log\left(\mathcal{F}_{\mathrm{mlp}}(L,W,S)\right),T\right)\leq CKSL\log(S)\log(B_{2}KT),$
 for some universal constant $C>0$.